A Complete Bibliography of Publications in the Journal of Software: Evolution and Process

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

15 March 2023 Version 1.08

Title word cross-reference

2 [BMP13]. _{4} [GGGBEM18]. _{NRFixer} [GS23]. p [IAA18a]. v [BSB16].

-algorithm [BSB16].

1 [WCS14]. 12th [KDK13]. 13th [FKW13]. 14971 [FMC^{+}15]. 15939 [ADCG12]. 18th [PP14]. 19th [OP14].

20000 [WCS14]. 20000-1 [WCS14]. 2010 [BMP13, CDF13]. 2011 [CT13, PP14, RPZ14]. 2012 [CLC14, MC14, OP14]. 2015 [AÖAG16, De 16, DHH^{+}16, KTHS16, LSBV17, LSCP16, MPS^{+}16, MMJB16, PKBT16, Pot16, RFM16, SA17]. 2016 [KOPR17]. 2019 [HAS21]. 20th [BMCP14c]. 26262 [IK14]. 29110 [LSCP16, MPS^{+}16, MM14, MPM^{+}20, SGSGYO19, TKS^{+}14]. 29110-based [MPS^{+}16].

3 [vdGK12]. 3.0 [MKMW18]. 31000 [BMM19]. 31000-based [BMM19]. 33014 [PHJ15].

4-year [MMJB16]. 4th [PF18].

5 [GCMS14, RMOP20].

60880 [VNM14]. 61508 [VNM14]. 62304 [RFM16].

9 [LSBV17]. 9126 [ADCG12].

A/B [FDM^{+}18]. abandonment [KC22]. abilities [SRSEGA14]. abnormality [AAYK20]. abstract [NSDD17, YZT22, BGM20]. academia [RBE^{+}21]. access [DLM23]. accessibility [SGSGYO19]. accidents [CKMN15]. according [LLLL21, MM14]. Accounting [Al 15]. accuracy [IAA18a]. Accurate [WL16, KTYK16, TUF20, WWQ^{+}21]. Achieving [BJF12, GCMS14]. Acknowledgements [Ano17m]. across [BDS^{+}22, JPvdW^{+}17, LSFG20]. action [KMP22, ST23]. actionable [MGN18]. activities [WWQ^{+}21]. actual [LHEG18]. ad [OB17]. adapted [HGGASVD14]. adaptive [GNIS18, LAF^{+}18, RM22]. ADAS [MKMW18]. Additional [JN12]. Addressing [BCPM20a, SL20]. Adopting [MMC14, NMA^{+}17]. adoption [ASA^{+}20, CV21, eLAA^{+}23, TST17]. Advanced [QHY^{+}22]. Advances [BCPM19, BMCP15]. affecting [KC22, WKB^{+}15]. affects [GWA15, RHD^{+}16]. after [EKV15, SJG^{+}14]. again [TPB^{+}17]. Agency [PBDJ16]. aggregating [MV14]. aggregation [MAL^{+}13]. Agile [GGJ20, HGGASVD14, PL14, RK18a, SVK^{+}14, TWFF18, ASA^{+}20, AAM^{+}21, ALK^{+}17, Ano21b, DT18, Dow14, El 18, ECRJVA^{+}18, JW12, KHR18, KMB17, KTP22, Kos21, KMP22, Kru13, MPB16, MMC14, NGAEE22, OMTR12, PA14, Pot16, PSMM19, RK18b, ST23, SKK^{+}18, TGE17, TST17, MGW^{+}17, RBG15]. Agile MANTEMA [PRGP12]. AHP [Yag18]. aid [El 18, MBD^{+}22]. Akari [HOD^{+}14]. algorithm [BSB16, KTYK16, KRG^{+}13, LSRT16]. algorithms [GAA^{+}22, KRR^{+}16]. Alignment [MGW^{+}17]. alleviate [GS23]. allocation [EKG18, FPA17, II17, RMCLGAM14]. allocation-based [EKG18]. Allura [TP21]. alternatives [LMdVT^{+}12]. AmaLgam [WL16]. among [AAM^{+}21, FM23, PBSA17, SS18]. amongst [AHKL20]. amplification [SAD22]. AmPyfier [SAD22]. analogy [AI21, HIA19, IAA18b]. analogy-based [IAA18b]. Analysing [LGNEAO20, KLWA20, SVK^{+}14]. Analysis [ECRJVA^{+}18, FPA17, KLG17, MPM^{+}20, AVC20, AL17, AAYK20, Ano21c, Ano21t, BSEA14, BHLO16, CPR15, CLC14, EPG18, FM23, JMCC18, Jun12, KKK^{+}21, Kos13, KFYI13, LSBV16, LSBV17, LHR18, LHA^{+}16, NSDD17, PMT20, QB21, RAM^{+}22, SJG^{+}14, SZ12, SP22b, Stu12, UY13, XLWZ15, YAMM22, ZWW^{+}20, CS13]. Analytic [KHZ21, SKK^{+}18]. analytical [AKKM20, ASK^{+}21, AKM^{+}21, KS20]. Analyzing [AN17, LLLW18, ABB14, BMZP14, RF16]. Android [MDV^{+}18, QB21]. animations [Sir18]. Anniversary [BMCP14c]. Ansible [QT22]. Antecedents [CCUP21]. Anti [MPSM22]. Anti-bloater [MPSM22]. Apache [TP21, TFAL21]. API [CdSdAM16, GRK17, HZ23, ND20, PJS^{+}17]. App [GLL^{+}21, DGVP21, GLL^{+}20, LLLL21, TRC^{+}17, LLLW18]. Application [BMCP14a, LLLL21, OCCSBDBP20, AHH17, AAYK20, BCC^{+}15, BSEA14, BW22, BMCP14b, GdSC21, HGGASVD14, LC23, Ost14]. Applications [WLL^{+}23, AMB18, ADFT16, BCD13, CYK20, CLAD17, DMT13, JSI^{+}22, LGS^{+}17, LB13, OFAT^{+}18, RRDPD^{+}13, UIK20, WRC^{+}13]. applied [PCCP14]. Applying [RMAB15, BM12a, De 15]. Appraisal [dH13]. appraisals [De 15, De 16, SV21]. Approach [ARSH16, SDG21, AKKM20, AHKL20, Ano21p, AAA13, BM13, BHZK12, BCD14, BCD16, CS19, CLW22, DCVM20, DP15, EIA19, FLGB22, GP16, GEK14, GKHE18, GS23, GDAG13, HF12, HH12, IK14, JMR18, JKIU22, KLC^{+}14, KKK^{+}21, Kos21, KMP22, Kuh15, KKW18, MAD^{+}14, NLNL20, NHLTL17, OKÓ^{+}17, ÖTM18, PGM21, RAM^{+}22, SLHS18, SKWD15, SSL15, SNP22, SLMW21, SBPS12, SAC21, UWKRS21, UTZ13, UIK20, WSB14, YF22, ZKK13]. Approaches [BDB^{+}18, BJF12, BMCP15, DRW14, KJS^{+}13, LM18, PUWS21, SH23]. Approaching [GCPM19]. ArchiMate [PCCRP21]. ArchiRev [PCCRP21]. ArchiRev-Reverse [PCCRP21]. architects [VKPvV13]. architectural [CNPR13, MB17, SAPF22]. architecture [AJP14, Bab12, LGNEAO20, LLSA22, NSDP15, OEC17, PJW17, WRC^{+}13, YYF^{+}22]. architectures [MKMW18, PJW17, PDP18]. Archive [MCSGBSA20]. Arg [KLWA20]. arguments [KLWA20]. art [GS18b]. artifacts [CDO^{+}13]. artificial [KTYK16]. aspect [AAJE13, BJF12, JMR18]. aspect-oriented [AAJE13, JMR18]. aspects [AY20, ABG18, KGW13, KJS^{+}13, MSB^{+}14, Pot16]. assessed [GWA15, YFL^{+}22]. Assessing [CRP14, LSCP16, RBFM15, TST17, UFF18, SG17]. assessment [Abu16, BMM19, BMP13, BMCP14a, CC13, CS13, CYK20, CG14, CWBC17, CPSGAA20, JN12, Jun12, KKN^{+}19, KMH^{+}15, LS14, LCM^{+}15, LMC16, LdONB22, dSLBF15, MMK16, MV14, ÖTM18, PPdAPP21, SKTW14, SN14, SBPS12, TKS^{+}14]. assessments [CRP14, VMCN20]. assessor [ISK15]. assets [Ano21v]. Assigning [KGPH12]. assignment [KRR^{+}16, LMdVT^{+}12, SBS20, ZKN^{+}21]. Assisting [PGAP17]. association [AAM^{+}21, AHKL20, LLJ^{+}14]. assumptions [EMRF14]. assurance [BW18, HMMR18, LML^{+}18, Pot16, PBDJ16]. asynchrony [JGHA14]. athletic [ZDV22]. Atlas [SNP22]. attack [QHY^{+}22]. attackers [IZJG18]. attribute [Jun12]. attributes [NBF18]. automate [GP16]. Automated [AMB18, HBG21, JRCMWE23, MBD^{+}22, MLM16, PPPdG20, VMSH13, WFB^{+}18, dGNFAT17, DLM23, JKIU22, NBD17, UIK20]. Automatic [FHS13, WPVS14, XBS^{+}21, NBCM17, MTG22]. Automatically [ND20]. Automating [SGD^{+}13]. automation [AEC23, JRCMWE23]. Automotive [Kra13, MKMW18, NRB13, NGAEE22, SSB^{+}20, VKC^{+}21]. autonomic [YF22]. autonomous [HMB^{+}18, ST22]. autonomy [GdSC21]. autotuning [DMZ^{+}22]. availability [BM13]. Avispa [ABB14]. aware [FAT22, MSK18, PKY14].

B [BAAO^{+}20, FDM^{+}18]. back [TPB^{+}17]. Bad [MRJ20]. bagging [ZHC^{+}18]. balance [HD17]. Balanced [dH13]. banking [Ano21b]. barriers [AHKL20, ALK^{+}20, BDB^{+}18]. Base [KL18]. based [AMH15, AAM^{+}20, AIA20, AKM^{+}21, ALR^{+}22, ASACG13, ABQO14, ADFT16, AL15, AI21, Ano21v, APL^{+}14, AYB^{+}18, BMM19, BH16, BSEA14, BCD16, BDS^{+}22, CDO^{+}13, CS19, CZY^{+}22, DRW14, DMT13, EKG18, EIA19, EMRF14, EKP18, FAT22, FTW16, GLL^{+}21, GGGBEM18, GP16, GAA^{+}22, GA20, GEK14, GKHE18, dH13, IAA18b, JbAPT18, JKIU22, KMB17, KKT12, KÓ16, KHZ21, KKK^{+}21, KST14, KFYI13, LMdVT^{+}12, LGNEAO20, aRL21, MLCW22, MKA20, MPS^{+}16, MA12, MCL^{+}18, NHLTL17, NSDP15, OCCSBDBP20, PMT20, PCCP14, RLMC19, RCL19, SLHS18, SDWA20, SKWD15, SKTW14, SG20, SGD^{+}13, TKS^{+}14, UIK20, WLL^{+}23, WSB14, YZT22, ZKK13, ZWY^{+}20, ZLM^{+}21, ZLW22, dSPF^{+}12, dGNA^{+}16, dGNFAT17]. basic [CMCG^{+}12, MPM^{+}20]. Bayesian [PGAP17]. becoming [IZJG18]. BEFTIGRE [CLAD17]. behavior [JMCC18]. behavioral [AHA17, GD21]. Behaviour [CLAD17]. Behaviour-driven [CLAD17]. behaviours [RGSR20]. being [DT18, RGSR20]. Bellwether [MK22]. benchmarking [MA20]. benchmarks [BAAO^{+}20]. benefits [JbAPT18, LMC16]. best [RAM^{+}22, TUF20]. better [KL18, NXC13]. between [ASG18, AN17, DRM21, KJ14, KCH^{+}17, KMH^{+}15, LSBV16, LSBV17, MGW^{+}17, SP22a, SAPF22, WFB^{+}18]. beyond [KPG14]. bi [SKWD15]. bi-level [SKWD15]. Big [SKR15, KKK^{+}21]. binary [QSM15]. Blended [ÇD19]. Blending [NGAEE22]. bloater [MPSM22]. blockchain [ALR^{+}22, CPSGAA20]. Blocks [WPVS14]. blueprint [SSB^{+}20]. boolean [PCL21]. boosts [ST23]. born [CDR12]. both [GLL^{+}21]. bottlenecks [PRNR14]. bounds [BGV18]. BPaaS [HMB22]. BPM [CV21, GCPM19]. BPMN [AYB^{+}18]. BPMN-based [AYB^{+}18]. Brazil [GdOK18]. BS [GGGBEM18]. BSCBAS [dH13]. Bug [WL16, ATR^{+}17, DRM21, HZ23, NHLTL17, RHD^{+}16, SBS20, SAK17, WWQ^{+}21, WSB^{+}21, ZTGG16]. bug-assignment [SBS20]. bugs [DRW14, GS23, KRR^{+}16, XLWZ15]. Build [DBP22, NH14]. Building [SP22a, CKP15]. bulk [MA20]. Business [Kos21, ARD^{+}12, CPMS^{+}14, DMS12, FAT22, GNIS18, HBG21, JMR18, LWY^{+}16, NSBR12, PCCP14, PPdAPP21, RLR21, ST23, Sch12, SZ12, Stu12]. business-to-business [FAT22, ST23]. bytecode [PB17]. ByteDance [WCDY22].

C [HD17, BGM20, GJ13, HZ23, LSBV16, LSBV17]. Can [COGFCA22, KMH^{+}15, TPB^{+}17, Kra13, PPdAPP21]. capabilities [FDM^{+}18]. capability [Cao12, GdOK18, SG20, UC12, WCS14, SKTW14]. care [EKP18, PDP18, WFB^{+}18]. Case [TGE17, AIA20, AY20, Ano21b, AN17, BLGZ18, BABW12, BCSS18, EKV15, GAA^{+}22, GP13, HD17, HZ14, KFYI13, KO13, LSW13, MKA20, MGUA14, MPB16, MMPN20, MBMR19, NH14, OCCSBDBP20, PMT20, PCCRP21, PCMCAC14, QHY^{+}22, RGS^{+}15, RCL19, RMOP20, RBE^{+}21, SDG21, TFAL21, TG20, TMDE21, VvV15]. case-based-reasoning-based [AIA20]. cases [GRK17, KMB17, LdONB22, PPPdG20]. catalog [PGCMA15]. Categorization [WLL^{+}23]. Category [KL18]. causes [RF16]. CC [LSBV16, LSBV17]. centered [GD21, SH23]. centric [NBD17, WPXZ13]. certification [CPMS^{+}14, KBM^{+}12]. Challenges [AM20, CKR18, CdMSNdCM^{+}14, IK17, SK18, AKKM20, AEC23, Cap13, GFK^{+}20, KHR18, KJ14, KHZ21, KKK^{+}21, KMP22, PUWS21, SKK^{+}18, UWGP15, MVRP20]. challenging [KS20]. Change [PJW17, AMA^{+}20, AKKM20, AMH^{+}20, ASK^{+}21, Ano21q, CMCG^{+}12, CdMSNdCM^{+}14, CS19, ETT^{+}21, EAK13, Ghe17, HLP15, JGHA14, KGPH12, KA20, LHEG18, PHJ15, SSB^{+}20, YZT22, ZHC^{+}18, ZLW22]. change-level [ETT^{+}21]. change-prone [EAK13]. change-proneness [ZHC^{+}18, ZLW22]. changes [Bol13, CKS18, EAK13, ECRJVA^{+}18, HCL^{+}17, JRSS13, JPW^{+}17, SAPF22, THP^{+}16]. changing [HD17, RM22, Rya16]. characterisation [GCPM19]. characteristics [SVK^{+}14, WSB^{+}21]. Characterizing [SAEH16, WCDY22, TBP^{+}17]. check [LP17]. checking [CGU22, FHS13, LHA^{+}16, NHLTL17, NL18]. Choice [KL18, NPLP12]. Choosing [PHJ15]. choreographies [FAT22]. class [Al 15, GS23, MPSM22]. class-imbalance [GS23]. Classes [SVB^{+}17, EAK13, FM23, SAC21]. classical [HD17, SKR15]. Classification [AJP14, GJC^{+}18, MK22, NLNL20, RLMC19, ZTGG16]. click [IZJG18]. click-fraud [IZJG18]. client [AMA^{+}20]. client-vendor [AMA^{+}20]. Clone [CKS18, PTK13, BKZ13, GK13, Kos14, SSL15, SKR15]. Cloned [HG13]. clones [KR15]. cloning [PCP18]. Cloud [EKP18, PJW17, AKM^{+}21, AMB22, CLAD17, FHS13, KKK^{+}21, WCDY22]. Cloud-based [EKP18, AKM^{+}21]. Cluster [DMT13, KMB17, JMCC18]. Cluster-based [DMT13, KMB17]. clustering [AI21, AFK22, CKP15, LS19, PCCP14]. clustering-based [AI21]. clusters [SJG^{+}14]. CMMI [GP13, GCMS14, RMOP20]. CMMI-DEV [RMOP20]. Co [dGNA^{+}16]. Co-located [dGNA^{+}16]. coaches [KTP22]. coaching [PL14]. Coadapting [KBHG17]. COBOL [HF12]. Code [ARSH16, KLG17, SVB^{+}17, WPVS14, AVC20, AM16, BHLO16, BGM20, CLC14, DRGP13, GDJV18, HG13, HCL^{+}17, JPW^{+}17, KMH^{+}15, LHEG18, MLM16, MdGC^{+}17, NBCM17, OKÓ^{+}17, PCP18, PB17, QSM15, QT22, QBO^{+}13, RA15, SLHS18, SWT^{+}16, San14, SP22a, SAPF22, SLVOP18, SLL^{+}21, SAK17, XBS^{+}21, YYF^{+}22, ZBGH22]. coding [HD17]. coengineering [TMDE21]. coherence [CNS14]. cohesion [Al 15]. Cohesive [KO13]. collaboration [RDD^{+}12]. Collaborative [MDSSMS15, DFGP18, JPW^{+}17, KLC^{+}14]. collection [KGW13]. combination [BJF12, ZHC^{+}18]. combinatorial [PPPdG20]. combined [HD17]. Combining [CS13, ZTGG16]. come [RZL^{+}15]. comments [MBD^{+}22]. commercial [Dow14]. commit [BKS^{+}17]. commits [GFKH18, TBP^{+}17, WG18]. communicated [SJA^{+}14]. communication [ECRJVA^{+}18, NPLP12, PK14, RBE^{+}21]. communicators [UWKRS21]. communities [RCO20]. community [EKP18]. companies [ALK^{+}17, BO21, FDM^{+}18, GP13, LCMCPA15, MPB16, ÖTM18, SH23]. company [GCMS14, MFA12]. Comparative [FZW^{+}18, Nia15]. compare [MKPB23]. Comparing [BKS^{+}17, DRW14]. comparison [AJP14, AG21, Ano21t]. competence [SRSEGA14]. competences [HLP15]. Compilation [WLL^{+}23]. compile [TPB^{+}17]. complete [TG20]. Completion [ARSH16, YYF^{+}22]. Complex [BBD^{+}22, BM13, FLGB22, LSW13, WYW^{+}13]. Compliance [CGU22, DMS12, NGAEE22, VMCN20]. compliant [Gal20, KK14, RFM16]. component [MA12, SP22b]. component-based [MA12]. components [CMCG^{+}12, LMOLG23, NEU^{+}19]. Composing [WL16]. composite [Ano21v, WRC^{+}13]. composition [GJC^{+}18, LM18, UTZ13]. compositions [AYB^{+}18]. comprehending [ADFT16]. Comprehension [RBM16, BMC18, GAA18, PA16]. comprehensive [Ano21b, LLJ^{+}14, MCL^{+}18, SLVOP18, TNLJ^{+}17, UWKRS21, WSB^{+}21, MCSGBSA20]. computations [WZ17]. computer [MDSSMS15, RGSR20]. computing [AMB22, HMB22, KKK^{+}21, LAF^{+}18, NMA^{+}17, Sir18]. concentration [BPKZ16]. concept [IK14]. conceptual [UWGP15, UGI^{+}14, ZPFV14]. Conceptualising [KLWA20]. concerns [BCD16, CGN17, MBMCP18]. conclusion [MK22]. Concrete [BLGZ18]. Conference [CDF13, FKW13, OP14, PF18, PP14, BMCP14c]. Confidence [ST22]. Confidence-driven [ST22]. configuration [BDGR12, DMZ^{+}22, NBD17]. conflict [WFB^{+}18]. conflicts [DBP22]. conformance [FHS13]. connected [OFAT^{+}18, CKR18]. connections [ASG18]. consequences [CCUP21]. consistency [HLWL15, LHA^{+}16]. consistent [TMDE21]. constraints [KL18]. construct [AAA13]. construction [Ano21a, GDJV18, NBCM17, TKK^{+}21]. consultancy [GCMS14]. consultants [GdOK18]. consumer [HB16]. consumption [AVC20, FAA^{+}17]. contagious [MB17]. content [BCC^{+}15, Cap13]. Context [ARSH16, HMB^{+}18, SLHS18, EMRF14, FAT22, LSFG20, PCP18]. context-aware [FAT22]. Context-based [SLHS18]. Context-dependent [HMB^{+}18]. Context-sensitive [ARSH16]. context-specific [EMRF14]. Contextual [CV21]. Contextualizing [Kru13]. continuity [PJW17]. Continuous [OEC17, DLM23, PGAP17, RM22, BDG^{+}23]. contract [AVC20]. contributions [LML^{+}18]. control [DLM23, ECRJFZ13, OCCSBDBP20, QSM15]. controlled [QBO^{+}13]. conventions [CNSR23]. conversation [BAAO^{+}20]. coordinated [SR19]. coordination [SSM21]. COQUALMO [HBH12]. corpus [LSBV16, LSBV17]. Correct [GDJV18]. Correct-by-construction [GDJV18]. correctness [KKW18]. correlation [GWA15]. Correlations [LS19, CKP15]. Corrigendum [LSBV17]. Cost [SV20, KTYK16, NSDP15, PBSA17]. Cost-driven [SV20]. cost-effective [NSDP15]. costs [BGV18]. coupling [APL^{+}14, CN12, KCH^{+}17]. Cover [GS18a, Raj18a]. coverage [CS19]. coverage-based [CS19]. Crafting [Kuh15]. CRAN [MCSGBSA20]. crash [NHLTL17]. create [KTP22]. Creating [Sir18]. Creation [RFM16]. crisis [ZCL^{+}18]. criteria [HB16, IAA18a, KKT12]. criterion [KL18]. critical [BZ21, CPR15, CPSGAA20, JN12, eLAA^{+}23, LYL^{+}12, MSBE20, SLMW21, ST22, Wal18]. criticality [DGVP21]. cross [CMQ^{+}20, CLW22, HH12]. cross-project [CMQ^{+}20, CLW22]. cross-section [HH12]. crosscutting [BCD16]. crossroads [ZRB^{+}14]. crowd [CdSdAM16]. CrowdRE [KLWA20]. CrowdRE-Arg [KLWA20]. Crowdsourcing [ZKN^{+}21]. CSCW [TNLJ^{+}17]. CSMR [Ano13, FKW13, MC14]. CSP [CZY^{+}22]. Cultural [MVP12, BABW12]. culture [ST23]. Current [BCPM20b, KKT12, LM18, SN14]. currently [DT18]. customer [KJ14, KPG14, RMAB15]. customized [LMdVT^{+}12]. cyber [BZ21, BGV18, BMS21, CYK20, GAA^{+}22, GdSC21, MSBE20, PMT20, PDP18, QHY^{+}22, RMP^{+}20, TMDE21, ZIDM20, ZIDM21]. cyber-attack [QHY^{+}22]. cyber-physical [BZ21, BGV18, BMS21, GAA^{+}22, GdSC21, MSBE20, PMT20, PDP18, TMDE21]. Cybersecurity [Wal18, MBMCP18]. cycle [AÖAG16, Sch12, ZPW12]. cycles [FM23, Hou14, MPM^{+}20].

Data [MA20, VBMH22, AIA20, Al 15, Ano21t, CYK20, CSGR17, GA20, JSI^{+}22, KKK^{+}21, KMH^{+}15, LHEG18, RL14, Stu12, SKR15, ZPFV14, ZTGG16, CS13]. Data-driven [MA20]. data-intensive [ZPFV14]. Database [MAzGB18]. databases [DMZ^{+}22, PCdGCP13]. datasets [AHH17, QT22]. De-motivators [ASA^{+}20, AAM^{+}21]. Dealing [KMP22]. debt [CWBC17, LSFG20, MB17, TFAL21, ZBGH22]. debugging [Ano21q, CS19, VHBK22]. December [LSBV17]. decentralised [FLGB22]. decision [Ano21v, El 18, FPA17, GP16, HB16, JMCC18, KS20, aRL21, dSLBF15, WKB^{+}15]. decision-making [Ano21v, JMCC18, KS20]. decision-support [GP16]. declarative [FLGB22, ZPW12]. decomposition [SL20, VKPvV13]. deep [ZLM^{+}21, ZLW22]. deep-learning-based [ZLM^{+}21]. Defect [CPR15, Ano21u, CMQ^{+}20, CLW22, ETT^{+}21, QT22, RLMC19, SLL^{+}21]. defective [CMQ^{+}20]. Defectivity [HBH12]. defects [HLWL15, JWP^{+}16, KCH^{+}17]. define [MA20]. definition [KPG14]. delivery [RGS^{+}15]. Delta [AMH15, VHBK22]. Delta-based [AMH15]. deltas [DCVM20]. demand [HF12]. demand-driven [HF12]. demonstration [SGD^{+}13]. demonstration-based [SGD^{+}13]. dependability [ZWY^{+}20]. dependence [DRW14, SJG^{+}14, UY13]. dependence-based [DRW14]. dependences [APL^{+}14, yCP12]. dependencies [BPKZ16, KMH^{+}15, LHEG18, SS18]. dependency [Ano21q, BPKZ16, FM23, JbAPT18]. dependent [HMB^{+}18]. dephase [JGHA14]. deployed [SPC16]. Deployment [ETT^{+}21, NSBR12, PGM21, RGSR20]. deployments [LC23]. derivation [OMTR12]. Deriving [JSI^{+}22, Sha15]. descriptions [DS17, LLLW18]. Design [BCD14, BGM20, CsZZ^{+}13, FMR13, WPVS14, BM13, BAAO^{+}20, BCD13, FAA^{+}17, NSBR12, OKÓ^{+}17, PVA^{+}21, PPdAPP21, RTCL16, SGSGYO19, SP22a, SV21]. Designing [LdONB22]. desirability [LdONB22]. detailed [CPR15]. detect [DP15]. Detecting [ATR^{+}17, BMZP14, JGHA14, CKS18]. detection [AAG21, BCD14, CS19, GK13, Kos14, MRJ20, MAMN19, PCL21, SSL15, SKR15, Wan21, WFB^{+}18, ZPFV14]. detectors [SKR15]. determining [DRW14]. DEV [GP13, RMOP20]. develop [GJC^{+}18]. developer [CSGR17, DRM21, KRR^{+}16, KC22, TBP^{+}17, WWQ^{+}21]. developer-related [TBP^{+}17]. developers [BCSS18, COGFCA22, KGPH12, KTHS16, MGN18, WPXZ13, XLWZ15]. Developing [SV21, MMC14, NBF18, PKY14]. Development [LCM^{+}15, LMC16, ASA^{+}20, AY20, AHA17, AMH^{+}20, ASK^{+}21, AKM^{+}21, AÖAG16, ASG18, AI21, Ano21b, Bab12, BM12a, BM13, BCC^{+}15, BH16, BJBM21, Cao12, CN12, ÇD19, DLM23, DT18, DHH^{+}16, EIA19, FPA17, GS18b, GKS^{+}12, GD21, HD17, HGGASVD14, HLP15, HIA19, IAA18a, IAA18b, IK17, II17, JWP^{+}16, KLC^{+}14, KKN^{+}19, KA20, KHZ21, KGW13, KTP22, Kru13, LSW13, LMdVT^{+}12, LC23, LLLW18, MMK16, MKA20, MPS^{+}16, MVRP20, MGUA14, MA20, MA12, NMA^{+}17, NPLP12, ND20, ÖTM18, PL14, PBSA17, PRNR14, PK14, PUWS21, RGSR20, RCL19, RMOP20, RK18a, RK18b, RMCLGAM14, RFM16, SSM21, SK18, SRSEGA14, SKK^{+}18, SAEH16, SLMW21, SH23, TGE17, TKK^{+}21, TWFF18, TRC^{+}17, THP13, TST17, VKC^{+}21, WWQ^{+}21, YAMM22, dSPF^{+}12, AL17]. device [CC13, LCM^{+}15, LMC16, ÖTM18]. DevOpRET [BDG^{+}23]. DevOps [ADH20, BEB21, BDG^{+}23, EAD17, JbAPT18, KS20, RYA^{+}21, RAM^{+}22]. diagram [SNP22]. diagrams [GAA18, MMPN20]. dictionary [HBG21]. differences [KJ14]. differencing [DCVM20]. different [CMQ^{+}20, MPM^{+}20]. differential [BH16]. Digital [BO21, OB20]. digitalization [PPdAPP21]. dimension [GSB^{+}20]. direct [San14]. directed [NHLTL17]. directions [AEC23, MGW^{+}17, RPV12]. Dirichlet [EKG18]. disasters [McQ12]. discipline [GdSC21]. Discovering [yCP12, PJS^{+}17, BSB16]. discovery [MLM16]. discriminative [HCL^{+}17]. disjunctive [PCL21]. dispatching [GAA^{+}22]. Disruption [OB20]. distance [Pri12]. distinguished [OP14, PP14]. Distributed [BBD^{+}22, BCSS18, FPA17, FLGB22, LMdVT^{+}12, Pri12, RBG15, SKK^{+}18, THP13, UFF18, YF22, dSPF^{+}12, dGNA^{+}16]. distribution [Ano21u, CN12, PBSA17, SP22b]. distrust [PNL12]. divergence [MGW^{+}17]. diversity [SVSC20]. Do [CMQ^{+}20, GWA15, LHEG18, DS17]. documentation [NSDP15, PJS^{+}17]. does [GKS^{+}12, RZL^{+}15, SWT^{+}16]. domain [ATR^{+}17, APL^{+}14, BSEA14, BO21, JN12, KKN^{+}19, MMC14, RPZ14, RLR21]. domain-based [APL^{+}14]. domains [BMCP14b, DT18]. down [AAA13]. drive [NRB13]. driven [AMH15, Ano21p, BJF12, BCD14, CLAD17, De 16, DP15, GGPT12, HF12, HD17, MA20, PGM21, SZ12, SV20, ST22, WWKC18]. driver [CNPR13]. drivers [SSB^{+}20]. DRIVES [SSB^{+}20]. DRIVES-EU [SSB^{+}20]. driving [MKMW18, ST22]. DSL [BCD14]. DSL-driven [BCD14]. Dthreads [FZW^{+}18]. Dual [XLWZ15]. duplicate [ATR^{+}17]. duration [BEB21, Hou14]. during [CKG15, CKS18, ECRJVA^{+}18, GJ13, HH12, PPdAPP21, SPC16]. Dynamic [Cao12, BCD16, CsZZ^{+}13, GS12, HBH12, QVR^{+}21, RPZ14, SAM13, WRC^{+}13]. Dynamics [II17, dSLBF15].

each [Kra13]. Early [PRNR14, AN18, LHEG18]. economical [Pot16]. Ecosystem [BO18, MCSGBSA20, OB17, TFAL21]. ecosystems [AY20, DRM21, HB16, OB20]. ECQA [ISK15, KBM^{+}12]. editing [ZKZH14]. edition [PF18]. editor [GW12, HAE14, HZR16, KLG17, MTG22, RBM16]. Editorial [Ano13, BMP13, BMCP14a, BMCP14b, BMCP15, BMCP16, CDR17, CDF13, CLC14, CT13, FKW13, GW12, HOJ15, HAE14, HZR16, KÓ16, KLG17, KDK13, LP17, MC14, MAD12, MGW^{+}17, MP16, PP14, RPZ14, RBM16]. Editorials [BM12b, BMCP14c, CDR12, NSBR12, RPV12]. edits [ZKK13]. education [Sir18]. educational [De 16]. effect [FAA^{+}17, VHBK22, ZKZH14]. effective [EPG18, GJC^{+}18, NSDP15, VvV15]. effectiveness [ASG18, AL15, HH12, RMOP20, VMCN20]. Effectivity [Pot16]. efficiency [JPvdW^{+}17, LSW13, RMOP20]. Efficient [ARSH16, ACM^{+}14, SSL15]. Effort [Ano21a, QB21, RCL19, TGE17, AL15, AI21, Ano21r, AN18, BCC^{+}15, BH16, EIA19, GS18b, HOJ15, HIA19, IAA18a, IAA18b, KFYI13, LHA^{+}16, MKA20, MA20, PBSA17, TUF20, vdGK12]. electronic [OCCSBDBP20]. electronics [HB16]. elevators [GAA^{+}22]. elicitation [AL17, DFGP18, SDWA20, SL20, YAMM22]. Embedded [BSEA14, BO21, DT18, MPB16, OB20]. embedding [SLL^{+}21]. Embracing [GJ13]. Emerging [BW18]. Empirical [ABG18, AAJE13, AAYK20, CAKA21, EIA19, HSvG17, LSBV16, LSBV17, PA14, AAM^{+}20, ALK^{+}17, AVC20, ALK^{+}20, AG21, BKZ13, CO13, CWBC17, CV21, CNS14, FQA17, GFK^{+}20, GdOK18, HZ23, JPvdW^{+}17, JRSS13, KRR^{+}16, KA20, LMB12, LGS^{+}17, LML^{+}18, NXC13, NMA^{+}17, PCP18, RHD^{+}16, SP22a, SS18, TBP^{+}17, VMSH13, YAMM22, ZKZH14, BAB^{+}22]. employing [MPB16]. employment [dSVB^{+}22]. empowering [MMJB16]. enabled [MLCW22]. Enabling [PB17]. enactment [THP^{+}16]. encapsulation [Al 15]. encryption [LGNEAO20]. End [LMOLG23]. End-user [LMOLG23]. Energy [JPvdW^{+}17, FAA^{+}17, SWT^{+}16]. engine [OCCSBDBP20]. engineered [GAA18]. Engineering [BBD^{+}22, OP14, PMT20, PP14, RRDPD^{+}13, AIA20, ARD^{+}12, ATR^{+}17, AHC21, BSEA14, BJF12, BCPM20b, Bor16, BDB^{+}18, CKR18, GKHE18, GGPT12, GWA15, HBM^{+}16, JW12, KÓ16, KFT14, LMB12, LP17, LFT19, LHIE18, MAD12, MBMR19, MPM^{+}20, OEC17, OMTR12, PCCRP21, PNL12, Pri12, RPZ14, RPV12, RBE^{+}21, RDD^{+}12, RBG15, SKWD15, SL20, WKB^{+}15, YFL^{+}22, dGNFAT17, dSVB^{+}22]. engineers [MVP12, PPdAPP21]. engines [MAzGB18]. enhanced [ZPW12].

Enhancing [Ano21b]. ensemble [Ano21t]. ensembles [Ano21r, HIA19]. Ensuring [TMDE21]. enterprise [GPM^{+}12]. enterprises [Abu16, ABG18, BAB^{+}22, CG14, ECRJFZ16, LCMCPA15, Pot16, SG20]. entities [HCL^{+}17, MPS^{+}16, MPM^{+}20]. entrepreneurial [TWFF18]. entropy [EIA19]. entropy-based [EIA19]. Envelopment [CS13, Ano21t]. environment [AMB22, AN17, SKK^{+}18]. environmental [JMCC18]. environments [GGGBEM18, PSMM19]. Equation [eLAA^{+}23]. equivalent [NLNL20]. erosion [LLSA22]. Error [LHA^{+}16, WLL^{+}23]. essential [PGM21]. estimate [BH16]. Estimating [AI21]. estimation [AL15, Ano21r, AN18, BCC^{+}15, EIA19, GS18b, HOJ15, HIA19, IAA18a, IAA18b, MKA20, MA20, PBSA17, QB21, RCL19, TGE17, TUF20]. ETGM [KRG^{+}13]. ethical [RGSR20]. ethics [MdGC^{+}17]. EU [SSB^{+}20]. European [FKW13, CDF13]. EuroSPI [BMCP14c, AÖAG16, De 16, DHH^{+}16, KTHS16, LSCP16, MPS^{+}16, MMJB16, Pot16, RFM16, BMP13]. evaluate [FQA17]. Evaluating [HH12, HIA19, IAA18a, QBO^{+}13, ISK15, JSI^{+}22, SBS20]. Evaluation [VBMH22, WPVS14, AN18, Bab12, CLAD17, El 18, EIA19, ECRJFZ13, Gal20, HHW12, LMdVT^{+}12, LB13, MBD^{+}22, NSBR12, RMP^{+}20, UGI^{+}14, VMCN20, VMSH13, WCS14, YT22]. Event [BAAO^{+}20]. Event-B [BAAO^{+}20]. events [ZCL^{+}18]. evidence [Ano21v, PA14, dSPF^{+}12]. evidence-based [Ano21v, dSPF^{+}12]. Evolution [CNPR13, KDK13, LLT20, MAzGB18, MCSGBSA20, PF18, QVR^{+}21, TFAL21, AJP14, BZ21, BH16, BCD13, CDR12, CKG15, CS13, CKS18, CPSGAA20, GLL^{+}20, GK13, GG14, GBRHO14, HH12, HZ23, KRG^{+}13, KKW18, LGS^{+}17, LLLW18, MDV^{+}18, MP16, NXC13, PTK13, RHL15, SJA^{+}14, SR19, TWL^{+}21, WYW^{+}13, ZDV22]. evolution-based [BH16]. evolutionary [BPKZ16, Bol13, JPW^{+}17, KRR^{+}16, KCH^{+}17]. evolvability [Raj18b]. Evolving [FDM^{+}18, TP21, BCPM20a, HSvG17, HAS21, JWP^{+}16, RPZ14]. EXACT [ADFT16]. examination [CO13, RGS^{+}15, SS18]. examining [Pri12]. example [GKHE18]. examples [GRK17, GEK14]. Excel [AHH17, ADFT16]. executables [CsZZ^{+}13]. Execute [SJG^{+}14]. Execution [HAN22, AAA13, AM16, BJF12, CsZZ^{+}13, MAD^{+}14, SJA^{+}14, UTZ13]. exhaustive [KGW13]. Experience [AÖAG16, BMP13, De 15, KK14, BRDL21, BDS^{+}22, HOD^{+}14, LSCP16, PMT20, SV20, TKS^{+}14, TP21, WWKC18]. experiences [BMP13, BMCP14c, LS19, RBFM15]. Experiment [AL15, De 16]. Experimentation [FDM^{+}18]. experiments [BAAO^{+}20, QBO^{+}13]. expert [MKA20, OCCSBDBP20, RHD^{+}16]. expert-based [MKA20]. experts [GSB^{+}20]. explication [HBH12]. Exploiting [SAK17]. exploration [ALK^{+}20]. explorative [HLP15]. exploratory [Ano21c, CNSR23, MPSM22, MDV^{+}18, Nia12, QT22, SP22a, SGAP14, SJA^{+}14, SKR15, WPXZ13, ZKJ12]. explore [RCO20]. Exploring [GFK^{+}20, KC22, WKB^{+}15]. expressions [PPPdG20]. Extended [LLJ^{+}14, QSM15]. Extending [KL18, KPG14, MKMW18]. extent [CO13]. external [LML^{+}18, SSM21]. extracting [KLWA20]. extreme [BW22].

faced [BCPM20a]. factors

[ASK^{+}21, AKM^{+}21, AN17, BABW12, BDB^{+}18, CV21, ECRJFZ13, JMCC18, KHR18, KC22, KS20, Nia15, RK18b, TBP^{+}17, WKB^{+}15, Yag18, YAMM22]. factory [LSCP16]. failure [ZKK13]. failure-inducing [ZKK13]. failures [ST22]. fair [EKV15]. fault [AAYK20, CS19, JSI^{+}22, LAF^{+}18, LLJ^{+}14, PCL21, PA16, SPC16, Wan21, ZLM^{+}21]. fault-prone [AAYK20]. faults [Ano21q, BKZ13, DP15]. FaultTracer [ZKK13]. Fear [PNL12]. feasibility [KFT14]. Feature [DRGP13, Ano21r, BDGR12, EKG18, HZ14, HCL^{+}17, LYL^{+}12, LLLL21, MLM16, WPXZ13]. features [YYF^{+}22, YZT22]. feelings [GWA15]. FermaT [PB17]. fertilised [BMCP15]. few [MA20]. file [RM22, ZKZH14]. filter [Ano21r, HIA19]. financial [KMP22, QHY^{+}22, VvV15]. find [BO18]. Finding [GJC^{+}18]. Findings [SK18, CKP15]. Fine [FM23]. Fine-grained [FM23]. firefly [KTYK16]. Five [Raj18b, CDR17]. fix [Hou14, TBP^{+}17]. fix-inducing [TBP^{+}17]. fixability [GS23]. fixing [OKÓ^{+}17]. flexibility [NBF18]. flexible [GNIS18]. FlexParser [RM22]. FLOSS [GBRHO14]. flow [QSM15]. fluctuations [PCMCAC14]. fluid [QT22]. focus [TRC^{+}17]. focused [KLC^{+}14]. focusing [CMCG^{+}12]. forecasting [Hou14, QHY^{+}22]. Foreword [SA17]. forges [TP21]. Formal [BRDL21, BAAO^{+}20, NL18, AMH15, BGM18, SLMW21, WZ17]. Formalizing [BDGR12, RRL22]. formation [AHKL20, ALK^{+}20]. formulas [AHH17]. Formulation [NEU^{+}19]. forums [KLWA20]. forward [ACM^{+}14]. FOSS [DP15]. four [GS12, PBSA17]. four-loop [GS12]. four-stage [PBSA17]. fragmentation [CSGR17]. Framework [SS18, AMH15, AAM^{+}20, AAM^{+}21, AKM^{+}21, AHKL20, Bab12, BCC^{+}15, BLGZ18, Bor16, DLM23, ECRJFZ13, KLWA20, KHZ21, KMP22, KO13, LS14, LCM^{+}15, LMC16, LB13, MMK16, NMA20, NBF18, OKA14, PKY14, RMCLGAM14, SDWA20, TGE17, TNLJ^{+}17, UWGP15, UGI^{+}14, ZIDM20, TST17]. frameworks [YT22]. Framing [PA16]. fraud [IZJG18]. full [CLAD17]. full-tier [CLAD17]. function [LC23]. functional [GNIS18, LGNEAO20, LLT20, MBMCP18, SL20]. functions [LSBV16, LSBV17, QSM15]. Further [ZBGH22]. Future [MGW^{+}17, EAK13, McQ12, MKMW18]. Fuzzy [AIA20, ASK^{+}21, AI21, El 18, HIA19, KKK^{+}21, LC23, RAM^{+}22, Yag18]. fuzzy-TOPSIS [KKK^{+}21].

game [PVA^{+}21]. gamification [AM20]. genealogies [BKZ13]. generalized [Hou14]. generate [GA20]. generating [LSRT16]. generation [BDS^{+}22, HBG21, JKIU22, MMPN20, NRB13, PPPdG20, UIK20, XBS^{+}21]. generator [BGM20]. geographic [CN12]. German [PBDJ16]. getting [GKS^{+}12]. GitHub [GFKH18, LML^{+}18, ZWW^{+}20]. GLOB [NMA20]. Global [JW12, MVRP20, AMA^{+}20, AMH^{+}20, ASK^{+}21, Bab12, BJBM21, CN12, GKS^{+}12, HLP15, IK17, II17, JWP^{+}16, KKN^{+}19, KA20, KHZ21, LMB12, MAD12, NMA^{+}17, NMA20, NPLP12, PL14, PNL12, Pri12, RK18a, RPV12, RMCLGAM14, SRSEGA14, UWGP15, YAMM22, dGNFAT17, AL17]. globally [BCSS18, Pri12, THP13]. go [dSVB^{+}22]. goal [GLL^{+}20, MKPB23, PGM21, SZ12]. goal-driven [PGM21, SZ12]. goal-question-metric [MKPB23]. Goals [PUWS21]. Going [OB20]. good [NEU^{+}19]. Governance [BBD^{+}22, ISK15, MVRP20, PPdAPP21]. grained [FM23]. grammar [AHH17]. graph [BCD14, CKP15, DMZ^{+}22, YYF^{+}22]. graphs [QSM15]. GreCo [MdGC^{+}17]. Green [MdGC^{+}17, CLAD17, RK18a, RK18b, SK18, CV21, GCPM19]. greenness [MAzGB18]. Group [RMOP20, ASG18, KMP22, MPB16]. Grouping [JMCC18]. groupware [Bab12]. groupware-supported [Bab12]. growth [Ano21q, CS19, FDM^{+}18, HSvG17]. GSD [AKKM20, IK17, MVP12, RK18b]. GSEPIM [KKN^{+}19]. guardian [KKW18]. Guest [CLC14, GW12, HZR16, KLG17, MC14, MTG22, RBM16]. GUI [AMB18, San14]. guide [MPM^{+}20]. guided [LLLW18]. Guidelines [SBS20].

handling [SN14, VvV14]. hands [Bor16]. hands-on [Bor16]. hardware [OCCSBDBP20]. hardware-in-the-loop [OCCSBDBP20]. hashing [Kos14]. HCI [GdOK18]. Health [BNBA21, RGSR20, OFAT^{+}18, PDP18, RLMC19, CKR18]. healthcare [ALR^{+}22]. healthiness [BZ21]. heart [ZDV22]. help [PPdAPP21]. hemodialysis [LHR18]. heuristic [GA20]. Hierarchical [VHBK22]. hierarchy [AKKM20, ASK^{+}21, AKM^{+}21, KS20, KHZ21, SKK^{+}18]. High [GCMS14, BW18, HMMR18]. High-maturity [GCMS14]. higher [BPKZ16]. highly [TUF20]. historical [EAK13]. history [OKA14]. hoc [OB17]. hoisting [VHBK22]. homogeneous [Ano21r, AFK22, HIA19]. Horizontal [HZ14, UC12]. human [ABG18, WPXZ13]. human-centric [WPXZ13]. Hungarian [SH23]. Hybrid [BHZK12, HAS21, ADH20, FPA17, GS23, JMR18, ÖTM18, PUWS21, RAM^{+}22, SGAP14, TKK^{+}21]. hyperheuristic [BLGZ18].

I/O [BMZP14]. ICONIX [LLT20]. ICPC [RBM16]. ICSM [CT13]. ICSSP [HAS21, KOPR17, PKBT16, RPZ14]. idea [NRB13]. Identification [KKK^{+}21, Ano21v, HCL^{+}17, LMdVT^{+}12, PRNR14, RAS^{+}13]. identifier [GDAG13]. identify [CMQ^{+}20]. Identifying [CKP15, GFKH18, KKT12, QSM15, SVB^{+}17, ND20]. IEC [LSCP16, MFA12, MM14, MPM^{+}20, PHJ15, RFM16, SGSGYO19, TKS^{+}14, VNM14, WCS14].

IEEE [PF18]. II [GS12]. IM [GS23]. Image [GS18a, Raj18a, XBS^{+}21]. image2emmet [XBS^{+}21]. imbalance [GS23]. imbalanced [Ano21u]. Impact [EKG18, SJG^{+}14, BL15, BABW12, CN12, HLWL15, LHEG18, PJ16, SWT^{+}16, SVSC20, SS18, TWL^{+}21]. impacts [PCP18]. imperfect [Ano21q, CS19]. implement [ALR^{+}22]. implementation [AMA^{+}20, BAB^{+}22, CsZZ^{+}13, FMC^{+}15, KK14, LCMCPA15, Nia12, Nia15, RYA^{+}21, SGSGYO19, SS18, SDWA20]. implementations [MPM^{+}20]. Implementing [PVA^{+}21, SDG21, AÖAG16]. implications [HSvG17, WCS14]. improve [AMA^{+}20, DCVM20, McQ12, MK22, RMAB15, VKPvV13]. improved [ARD^{+}12, PCMCAC14, SDWA20, SAK17]. Improvement [BM12b, BMCP16, GSB^{+}20, AM20, ABG18, BAB^{+}22, BMP13, BMCP14a, BMCP14b, BMCP15, BCPM19, CKMN15, CMCG^{+}12, CKP15, CG14, CCUP21, CO13, CPMS^{+}14, ECRJFZ13, GPM^{+}12, GP13, HHW12, KK14, KKT12, KKN^{+}19, KJS^{+}13, Kuh15, MBMCP18, MM14, Nia12, Nia15, NEU^{+}19, OKA14, PGAP17, PSMM19, RAS^{+}13, RMP^{+}20, SSA15, SG20, TGE17]. Improving [BM12a, BM13, CDO^{+}13, LYL^{+}12, THP13, VvV14, ZLM^{+}21, dH13, Kos21, MMK16, UFF18]. imputation [AIA20, IAA18b]. incident [VvV14, VvV15]. incomplete [AIA20]. inconsistent [CKS18]. Increasing [LSW13]. Incremental [Kos13, GK13]. independent [CLW22]. indexed [LLLL21]. indexing [CDO^{+}13]. India [ŠW12]. Indian [CCUP21, SG20]. indicators [GKS^{+}12]. individual [JMCC18]. inducing [TBP^{+}17, ZKK13]. Industrial [BMP13, AY20, BMCP14c, ETT^{+}21, GAA^{+}22, KCH^{+}17, LSW13, LS14, PCCRP21, RL14, SK18]. industry [BW22, BCPM20a, BDS^{+}22, CCUP21, CWBC17, GFK^{+}20, MDSSMS15, MAL^{+}13, NGvD13, NRB13, RBE^{+}21, SSA15, SSB^{+}20]. industry-academia [RBE^{+}21]. infer [MBMR19]. inference [WWQ^{+}21]. influence [ECRJFZ13, RTCL16]. influencing [JMCC18]. informal [DMS12]. Information [Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22a]. Information [Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, VvV15, WL16, CG14, LHIE18, PCCRP21, RGSR20, SJA^{+}14, VvV14]. informed [Gal20]. infrastructure [HOD^{+}14, QT22]. infrastructures [FHS13]. inhibitors [SS18]. initiatives [CPSGAA20, KKT12, SS18]. Innovation [BMCP16, BCPM19, MBMCP18, NRB13, RDD^{+}12]. insertion [PCL21]. Insights [KTP22]. inspection [CKS18, EMRF14, MFCP14]. INSPEX [BRDL21]. instability [MK22]. institutions [QHY^{+}22]. instructions [Bor16]. Integrated [HBM^{+}16, BM13, BMM19, ISK15, RDD^{+}12]. Integrating [EMRF14, KTYK16, THP^{+}16, VNM14].

Integration [SKTW14, SGSGYO19, vdGK12, IK17, KLC^{+}14, SG20, WG18]. Integration-based [SKTW14, SG20]. integration-focused [KLC^{+}14]. intelligence [MSK18]. intelligent [YF22]. intensive [AAA13, KGW13, OB20, ZPFV14]. intentional [NBD17]. inter [Kos14, MPB16]. inter-group [MPB16]. inter-system [Kos14]. interaction [CSGR17, MPB16, MMPN20, San14, WWQ^{+}21]. interactions [DRM21, ECRJVA^{+}18]. interactive [BDGR12]. interest [MB17]. interface [LLLL21, XBS^{+}21]. Internal [LML^{+}18]. International [KDK13]. Internet [BNBA21, IZJG18]. interoperability [YF22]. interpretation [NSDD17]. interpretive [AHKL20]. interprocedural [yCP12]. interrupts [KTHS16]. interview [KTP22]. introducing [OKÓ^{+}17]. Introduction [BMC18, FKW13, GGPT12, KÓ16, KOPR17, MP16, PKBT16, GW12, HAE14, HZR16, KLG17, MTG22, RBM16, Wan21]. Investigating [DGVP21, FAA^{+}17, SVSC20, JWP^{+}16]. investigation [BAB^{+}22, CPR15, JRSS13, KRR^{+}16, KA20, LML^{+}18, PBSA17, ZBGH22]. Involving [Lin15]. IoV [CZY^{+}22]. IR [CDO^{+}13]. IR-based [CDO^{+}13]. ISM [RAM^{+}22]. ISO [ADCG12, BMM19, FMC^{+}15, IK14, KO13, LSCP16, MPS^{+}16, MFA12, MM14, MPM^{+}20, PHJ15, SGSGYO19, TKS^{+}14, WCS14]. ISO/IEC [LSCP16, MFA12, MM14, MPM^{+}20, PHJ15, SGSGYO19, TKS^{+}14, WCS14]. Issue [Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22a]. Issue [Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, BMP13, FKW13, KLG17, KOPR17, PKBT16, RBM16, Ano13, BMC18, BBD^{+}22, CKR18, CLC14, HAS21, KÓ16, KDK13, MC14, MAD12, MTG22, MP16, NSBR12, PP14, RPZ14, SA17]. issues [DGVP21, LM18, MVP12, SN14]. IT-supported [NRB13]. items [ZBGH22]. iterative [HBM^{+}16]. ITIL [LCMCPA15]. ITMark [LSCP16]. IWSC’13 [KR15].

Java [AAYK20, LSBV16, LSBV17, SAC21, TFPB14]. JavaScript [NSDD17, SVB^{+}17]. joint [PGM21, RBE^{+}21]. JointConf [DMZ^{+}22]. Jointly [DMZ^{+}22]. journal [CDR12, PF18]. Jsvee [Sir18]. just [HZ14]. just-in-time [HZ14].

Kanban [ALK^{+}17, SDG21]. KASL [GS12]. KASL-II [GS12]. Kelmu [Sir18]. kernel [NH14]. kernels [WYW^{+}13]. key [Yag18]. know [GKS^{+}12]. Knowledge [BABW12, PPdAPP21, ATR^{+}17, AJP14, CdSdAM16, CCUP21, CPMS^{+}14, CNS14, CNPR13, GS12, LLLL21, NEU^{+}19, RCO20, RHD^{+}16, SP22a].

label [MAD^{+}14]. Labelling [ZCL^{+}18]. Language [BZ21, dGNA^{+}16, HAN22, SNP22]. languages [Ano21p]. LAPPI [RAS^{+}13].

Large [Kos14, ASA^{+}20, AHH17, AAA13, DRM21, KHR18, KCH^{+}17, KRG^{+}13, LSBV16, LSBV17, MPB16, PRNR14, Pot16, SSM21, SSL15, SJA^{+}14, TUF20, WKB^{+}15]. Large-scale [Kos14, ASA^{+}20, SSM21, WKB^{+}15]. late [BKZ13]. latent [EKG18]. law [SP22b]. laws [GBRHO14, GG14]. Layer [OB17]. LDA [BHLO16]. lead [WKB^{+}15]. leakage [LHA^{+}16]. Lean [PSMM19, SSA15, ALK^{+}17, RL14]. learn [COGFCA22, Kra13]. learned [BMCP16, CRP14, MFA12, MMJB16, NGvD13, RGS^{+}15, RBFM15, ŠW12, VKPvV13]. Learning [Ost14, GAA^{+}22, GS12, LMB12, NLNL20, SLL^{+}21, SAC21, YZT22, ZLM^{+}21, ZLW22]. learning-based [GAA^{+}22]. Least [eLAA^{+}23]. Legacy [SVB^{+}17, BCSS18, HF12, PCdGCP13]. Lehman [GG14]. Lessons [MFA12, MMJB16, ŠW12, BMCP16, CRP14, NGvD13, RGS^{+}15, RBFM15, VKPvV13]. Level [RMOP20, ETT^{+}21, SKWD15, TST17, UIK20]. levels [GCMS14]. leverage [NRB13]. leveraged [BMCP14b]. LHV [Ano21b]. library [TFPB14]. life [AÖAG16, MPM^{+}20, SDWA20, Sch12, ZPW12]. life-cycle [ZPW12]. lifecycle [NSBR12]. lifetime [Ano21s, BPKZ16]. light [RAS^{+}13]. light-weight [RAS^{+}13]. lightweight [CC13]. limit [CLMZ13]. line [HOD^{+}14, KFT16, LHIE18, NBD17, OMTR12]. linear [AL15]. lines [BSEA14, BSB16, CGN17, Gal20, QVR^{+}21]. linguistic [MVP12]. Linux [NH14, WYW^{+}13, WG18]. literal [PCL21]. Literature [NBF18, AAG21, AHC21, BNBA21, BGM18, BDB^{+}18, CGU22, JbAPT18, KA20, PDP18, RL15, SK18, SSB^{+}20, VKC^{+}21, YFL^{+}22, YT22, ZKN^{+}21]. lived [GBRHO14]. LLP [RGS^{+}15]. load [ABS^{+}23]. local [AAYK20]. Localization [WL16, LLJ^{+}14, PA16, SAK17, ZLM^{+}21]. localizing [ZKK13]. located [dGNA^{+}16]. location [DRGP13, EKG18, FAT22, WPXZ13]. location-based [FAT22]. locators [LSRT16]. Log [VBMH22, RM22, Sha15]. logging [Ano21c]. logistics [KO13]. logs [UTZ13]. long [GBRHO14, HSvG17]. long-lived [GBRHO14]. long-term [HSvG17]. lookup [WSB14]. loop [GS12, OCCSBDBP20].

Machine [GAA^{+}22, LHR18, NLNL20, SAC21]. Machines [BGM20]. maintainability [ASACG13, LSW13]. Maintenance [BBD^{+}22, FKW13, WRC^{+}13, Ano13, CDF13, CLMZ13, KFYI13, MBD^{+}22, MP16, PRGP12, QT22, SGD^{+}13, TWL^{+}21, UWGP15]. Make [SBPS12, LYL^{+}12]. making [Ano21v, JMCC18, KS20, aRL21]. Malaysian [ABG18]. malware [MDV^{+}18]. manage [IK14]. management [AAM^{+}20, ASA^{+}20, ALK^{+}17, AMA^{+}20, AKKM20, AMH^{+}20, ASK^{+}21, Ano21t, BHZK12, BCC^{+}15, BMM19, CYK20, CG14, EKP18, FLGB22, GGPT12, HMB22, ISK15, KA20, KJS^{+}13, Kos21, KST14, Lin15, MMK16, MMJB16, MM14, MPM^{+}20, NRB13, NMA20, OB17, ZCL^{+}18, dSPF^{+}12]. manager [KBM^{+}12]. Managing [AY20, GPM^{+}12, HMMR18, JRSS13, LSFG20, UWGP15]. manifesto [KBM^{+}12, GSB^{+}20]. Manipulation [KLG17, CLC14]. map [KGW13]. mapping [CdMSNdCM^{+}14, CAKA21, KPG14, KFT14, LLSA22, LHIE18, MMPN20, MdlRT19, RRDPD^{+}13, TWL^{+}21, ZIDM21, dSPF^{+}12, dSVB^{+}22]. MapReduce [MdlRT19]. market [GGPT12, HGGASVD14]. market-driven [GGPT12]. Markovian [BEB21]. mass [HGGASVD14]. mass-market [HGGASVD14]. matching [BCD14]. matter [GKS^{+}12, GWA15]. Maturity [CG14, KJ14, SKTW14, ALR^{+}22, ASG18, GdOK18, GGPT12, GGJ20, GCMS14, HHW12, Kos21, PPdAPP21, RMOP20, SVK^{+}14, SG20, TST17]. MDE [ABQO14]. MDE-based [ABQO14]. MDevSPICE(R) [LCM^{+}15, LMC16]. mean [RL14]. Meaningful [WPVS14, BGV18]. means [CS13, LSCP16]. Measurement [ADCG12, Al 15, KO13, RLTTC14]. measures [LLT20, LLJ^{+}14, MV14]. Measuring [SG20]. mechanism [RCO20, RMP^{+}20]. mechanisms [LGNEAO20]. mechatronic [BM12a]. media [CPMS^{+}14, NPLP12]. medical [BGM18, CC13, Kra13, LCM^{+}15, LMC16, MMK16, MMC14, ÖTM18, PK14, TRC^{+}17, WFB^{+}18]. MedITNet [MMK16]. medium [Abu16, ABG18, BAB^{+}22, CG14, LCMCPA15, SG20]. medium-sized [LCMCPA15]. Merge [WG18]. Merge-Tree [WG18]. Meta [Sch12, AHA17, RRDPD^{+}13]. meta-model [AHA17, RRDPD^{+}13]. Meta-service [Sch12]. metamodel [SNP22]. Metamorphic [AEC23]. Method [WPVS14, Abu16, AL17, AAYK20, AFK22, AN17, CC13, DMS12, DHH^{+}16, GA20, HB16, JWP^{+}16, KMH^{+}15, KFT14, MV14, RLR21, SPC16, SZ12, SAEH16, SR19, UWKRS21]. Methodological [RMCLGAM14]. methodologies [ASA^{+}20, NGAEE22, UFF18]. Methodology [CMCG^{+}12, BCC^{+}15, FPA17, KST14, PRGP12]. methods [AG21, AAJE13, BRDL21, BGM18, CMQ^{+}20, El 18, FQA17, HHW12, LSBV16, LSBV17, PJS^{+}17, QBO^{+}13, RGS^{+}15, RK18b, SAC21, TKK^{+}21, ZWW^{+}20, ZKN^{+}21, ZHC^{+}18]. metric [MKPB23]. Metrics [Bol13, VKC^{+}21, AVC20, CS13, EAK13, KFYI13, MRJ20, MAL^{+}13, PGAP17, SVSC20, Sha15, SAEH16, SAC21]. Mexican [ECRJFZ16]. Micro [KFYI13, FMR13]. micro-structures [FMR13]. microservice [MLCW22]. microservices [OEC17, WCDY22]. migrating [FHS13]. migration [NGvD13, RL14, RL15, SV20]. migrations [TFPB14]. milestone [ZWW^{+}20]. minds [RTCL16]. mindset [KTP22]. mined [LLLL21]. Mining [BCD16, GRK17, yCP12, JGHA14, JPW^{+}17, PJS^{+}17, RA15, SKTW14, UTZ13, ZTGG16]. minority [Ano21u]. Misaligned [LFT19]. Mission [eLAA^{+}23, CPR15]. Mission-critical [eLAA^{+}23, CPR15]. mitigation [BJBM21, ZIDM20]. mixed [Ano21t, HMB^{+}18, ZWW^{+}20]. mixed-methods [ZWW^{+}20]. ML [GCMS14]. ML-5 [GCMS14]. Mobile [PKY14, TRC^{+}17, AMB18, CLAD17, LGS^{+}17, LC23, UIK20]. Model [BDS^{+}22, GEK14, GKHE18, OB17, SKWD15, AMH15, AHA17, AMH^{+}20, ALR^{+}22, ASACG13, Ano21p, Ano21q, BDGR12, BMM19, BH16, BSEA14, BJF12, ÇD19, DP15, FTW16, GLL^{+}20, GGGBEM18, GS12, GdSC21, GJC^{+}18, GdOK18, GGPT12, GD21, HD17, Hou14, JKIU22, JSI^{+}22, eLAA^{+}23, LMdVT^{+}12, LC23, dSLBF15, MBMR19, MFCP14, MAMN19, NHLTL17, NL18, OMTR12, PGCMA15, PBSA17, QHY^{+}22, RYA^{+}21, RCL19, RF16, RRL22, SRSEGA14, SLVOP18, SVK^{+}14, SSA15, SG20, SGD^{+}13, SG17, TST17, UIK20, Wan21, WWKC18, dSPF^{+}12, RRDPD^{+}13, SKTW14]. Model-based [BDS^{+}22, BSEA14, GGGBEM18, UIK20]. model-driven [AMH15, Ano21p, BJF12, DP15, HD17]. modeler [PJ16]. Modeling [CZY^{+}22, HAN22, MGUA14, AAM^{+}21, Ano21p, BHZK12, BJF12, COGFCA22, GGGBEM18, JMR18, LMOLG23, MCL^{+}18, PJ16, RAS^{+}13, RMAB15, SS18, TNLJ^{+}17, ZKJ12, ZWK18, ZWY^{+}20, eLAA^{+}23]. Modell [KFT16]. modelling [AHKL20, CS19, FZW^{+}18, PMT20, RLR21]. models [ABB14, AG21, Ano21t, DS17, HAN18, HHW12, KTYK16, KLC^{+}14, LdONB22, NL18, PJW17, PCCRP21, RRDPD^{+}13, VMSH13, YT22, ZKJ12]. modern [KJS^{+}13]. modernization [PCdGCP13]. modification [HAN18]. modularization [DMT13]. modularized [DMZ^{+}22]. modules [AFK22, BPKZ16, CMQ^{+}20]. monitoring [AYB^{+}18, LWY^{+}16, MLCW22]. Motivators [NMA^{+}17, ASA^{+}20, KA20, AAM^{+}21]. moving [AL15, MK22]. MPT [SLL^{+}21]. MPT-embedding [SLL^{+}21]. Multi [Ano21q, AFK22, BCSS18, GKHE18, HB16, KK14, OKÓ^{+}17, UC12, UWKRS21]. multi-case [BCSS18]. multi-criteria [HB16]. multi-method [UWKRS21]. multi-objective [GKHE18, OKÓ^{+}17]. multi-sourcing [UC12]. multi-standard [KK14]. Multi-upgradation [Ano21q]. multicase [BO21]. Multicloud [LM18]. Multicriteria [El 18, KS20, aRL21]. multidimension [KBHG17]. multilevel [VMSH13]. multimodel [KKT12, LS14]. multiobjective [KRR^{+}16]. multiplatform [SR19]. multiple [BM12a, MPB16]. multisourcing [SK18]. multisystem [DRM21]. multivocal [AMA^{+}20]. mutants [NLNL20]. mutation [BW22].

NasBASE [KÓ16]. natural [dGNA^{+}16]. natural-language [dGNA^{+}16]. nature [CN12]. navigability [VMSH13]. NDN [CZY^{+}22]. NDN-based [CZY^{+}22]. need [DS17]. needed [ABS^{+}23]. needs [MDSSMS15, RDD^{+}12]. net [SNP22]. nets [BEB21]. Network [MCSGBSA20, AHC21, JbAPT18, KTYK16, WYW^{+}13]. networks [MMK16, PGAP17, VvV15, dSVB^{+}22]. neural [HMB22, KTYK16, YYF^{+}22, dSVB^{+}22]. next [RPV12, dSVB^{+}22]. no [CKMN15]. non [BEB21, COGFCA22, YFL^{+}22, GS23]. non-developers [COGFCA22]. non-Markovian [BEB21]. non-quality [YFL^{+}22]. Non-Reproducible [GS23]. nonfunctional [SGD^{+}13]. nonlinear [Wan21]. nonradial [Ano21t]. nonterminating [WZ17]. normalization [BL15]. Norway [SDWA20]. Norway-based [SDWA20]. notation [COGFCA22]. noun [CDO^{+}13]. noun-based [CDO^{+}13]. nuclear [VNM14].

O [BMZP14]. obfuscation [SWT^{+}16]. object [Al 15, EAK13, JWP^{+}16, KRG^{+}13]. object-oriented [Al 15, EAK13, JWP^{+}16, KRG^{+}13]. Objective [De 16, AFK22, GKHE18, KRR^{+}16, OKÓ^{+}17]. Objective-driven [De 16]. offshoring [BABW12]. Onboarding [BCSS18]. Online [BGV18, FDM^{+}18]. Ontology [PCCP14, AAM^{+}20]. Ontology-based [PCCP14, AAM^{+}20]. Open [Wan21, Ano21s, Cap13, CNSR23, DRM21, HZ14, KC22, KFYI13, eLAA^{+}23, LGS^{+}17, NXC13, QB21, RCO20, YT22]. open-source [Ano21s, CNSR23, eLAA^{+}23, LGS^{+}17, NXC13]. OpenPPM [MMJB16]. operation [NSBR12]. operations [KFT16]. operator [ZCL^{+}18]. operators [PGM21]. opportunities [Cap13, CdMSNdCM^{+}14, MRJ20]. optimization [NSDP15, WSB14]. optimization-based [NSDP15, WSB14]. optimizing [HMMR18]. oracled [PPPdG20]. oracles [GAA^{+}22]. Orchestration [HAN22]. order [SAK17]. ordinal [MV14]. organization [NBF18, RMOP20, WWKC18]. organizations [BMM19, ÇD19, ECRJFZ13, ECRJFZ16, dH13, KHR18, LFT19, PRGP12, RYA^{+}21]. oriented [ARD^{+}12, Al 15, AAJE13, BHZK12, EAK13, Gal20, JWP^{+}16, JMR18, KRG^{+}13, RLR21, SV21, UC12, WPXZ13, WRC^{+}13, ZWK18]. origins [DRW14]. osmosis [LMB12]. Osmotic [FLGB22]. other [Kra13]. outcome [WKB^{+}15]. outsourced [CRP14]. outsourcing [AKM^{+}21, AHKL20, ALK^{+}20, EKV15, UC12]. over-sampling [Ano21u]. Owners [UWKRS21].

package [MCSGBSA20]. packages [CKP15, RMCLGAM14, SVSC20]. Pakistan [SSA15]. PAM [Abu16]. Papers [BMP13, CDF13, KDK13, OP14, PP14]. paradigm [YAMM22]. parallel [AMB18, LWY^{+}16, SSL15]. parameters [DMZ^{+}22]. parser [RM22]. Parsing [AHH17]. Partial [eLAA^{+}23]. partially [CRP14]. Partition [KL18]. partnership [AHKL20, ALK^{+}20]. passing [LdONB22]. past [McQ12]. pathways [WFB^{+}18]. Pattern [ZPFV14, ZWY^{+}20, BCD14]. Pattern-based [ZWY^{+}20]. patterns [FAA^{+}17, FMR13, HAN18, HZ23, JGHA14, JPW^{+}17, MGUA14, OKÓ^{+}17, PVA^{+}21, RRL22, SGSGYO19, SP22a, UTZ13, ZKZH14]. payment [EKV15]. people [KJS^{+}13]. perceived [Pri12]. Perceptions [VBMH22, PJ16]. perform [WPXZ13]. Performance [AMB22, BMZP14, CV21, EKP18, GAA^{+}22, dH13, LB13, OCCSBDBP20, PMT20, PJ16, RF16, UIK20, VvV14]. performance-based [PMT20]. personal [RF16]. personality [AY20]. perspective [ASA^{+}20, BPKZ16, GGJ20, KPG14, Lin15, OEC17, SVK^{+}14, WYW^{+}13]. perspectives [BCPM20b, MKPB23]. Petri [BEB21, SNP22]. phase [Ano21a, IK14]. phases [LSFG20, LLT20, PBSA17]. phenomenon [MB17]. physical [BZ21, BGV18, BMS21, CYK20, GAA^{+}22, GdSC21, MSBE20, PMT20, PDP18, RMP^{+}20, TMDE21, ZIDM20, ZIDM21]. PI [NEU^{+}19]. picture [UWKRS21]. pilot [ADH20, GSB^{+}20]. placement [MSK18]. plan [RFM16]. planning [ASG18, KO13]. plans [PCMCAC14]. platform [BM12a]. platforms [FHS13, ZKN^{+}21]. PLC [LSFG20]. PLM [GGGBEM18]. PMBOK(R) [MM14]. POEML [HAN22]. point [Ano21q, CS19]. points [AN17, BCC^{+}15, MKA20]. policy [AYB^{+}18]. popularity [SVSC20]. portfolio [ALK^{+}17]. porting [SR19]. possible [KJ14]. post [BKS^{+}17]. post-commit [BKS^{+}17]. power [SP22b]. Practical [BGV18, BW22, RAS^{+}13, RBFM15]. practice [EAD17, GGGBEM18, GdOK18, MKPB23, MGW^{+}17, OKA14, PA14, SKTW14, SZ12, SH23, UWKRS21]. Practices [KHR18, BABW12, CWBC17, HGGASVD14, JW12, MPS^{+}16, MGUA14, MMC14, RAM^{+}22, RK18a, SSA15, TWFF18]. Practitioner [ALK^{+}20]. practitioners [DS17]. Pragmatic [AM16]. pragmatism [LP17]. pre [BKS^{+}17]. pre-commit [BKS^{+}17]. Pred [IAA18a]. predict [EAK13]. Predicting [Ano21s, APL^{+}14, GS23, ST22, SAC21]. prediction [AG21, Ano21a, Ano21u, BEB21, CMQ^{+}20, CLW22, ETT^{+}21, LHEG18, SLL^{+}21, YZT22, ZHC^{+}18, ZLW22]. predictions [KTYK16]. Predictive [HMB22, QT22]. predictors [TUF20]. Preface [KR15]. prefactoring [SLHS18]. preprocessor [GJ13]. presence [SJG^{+}14]. principle [SKTW14]. prioritization [BLGZ18, KMB17, KHZ21, KKK^{+}21, LYL^{+}12]. prioritize [ASK^{+}21, SLHS18]. Prioritizing [SKK^{+}18, Yag18]. proactive [RCO20]. probabilistic [NL18]. probability [CS19]. problem [GS23]. problems [KJ14, KKT12, MVP12, RF16]. Process [BM12b, BMCP15, BMCP16, GSB^{+}20, Jun12, PK14, RDD^{+}12, RRDPD^{+}13, SKTW14, SZ12, VBMH22, Abu16, AMA^{+}20, AKKM20, ASK^{+}21, AKM^{+}21, ABB14, ABQO14, AM20, ADH20, ABG18, Bab12, BHZK12, BAB^{+}22, BMM19, BKS^{+}17, BJF12, BMP13, BMCP14a, BMCP14b, BCPM19, BCPM20b, BSB16, CKMN15, CMCG^{+}12, CKP15, CG14, CCUP21, CO13, CLMZ13, CNS14, De 15, De 16, DMS12, DS17, Dow14, ECRJFZ13, Gal20, GGGBEM18, GPM^{+}12, GP16, GP13, GdOK18, GGPT12, HAN18, HD17, HHW12, HLWL15, JMR18, JRCMWE23, JN12, KK14, KLC^{+}14, KKT12, KKN^{+}19, KS20, KA20, KHZ21, KBHG17, KJS^{+}13, Kos21, KMP22, Kuh15, KFT16, KFYI13, LCM^{+}15, LMC16, LYL^{+}12, LHA^{+}16, Lin15, aRL21, LLT20, MBMCP18, MV14, MA12, Nia12, Nia15, NEU^{+}19, NSBR12, OMTR12, OKA14, PCCP14, PJ16, PSMM19, RAS^{+}13, RMAB15, RF16, RLR21]. process [SDWA20, SGSGYO19, SL20, Sch12, SN14, SV21, SGAP14, SKK^{+}18, SG20, SBPS12, Stu12, THP^{+}16, UC12, VMCN20, WCS14, WPXZ13, WWQ^{+}21, ZKJ12, ZRB^{+}14, ZWK18, ZWY^{+}20, HAN22, PF18, CDR12]. process-based [KHZ21]. process-oriented [SV21, UC12, WPXZ13]. process-related [KMP22]. processes [ARD^{+}12, BSEA14, CGU22, CRP14, DMT13, DHH^{+}16, EMRF14, FAT22, GNIS18, HAS21, LWY^{+}16, Ost14, RPZ14, RMAB15, Rya16, SH23, THP13, ZPW12]. product [BSEA14, BDS^{+}22, CGN17, GGPT12, HOD^{+}14, HBH12, JPvdW^{+}17, LHIE18, NBD17, OMTR12, PBDJ16, QVR^{+}21, RDD^{+}12, UWKRS21]. production [NSDP15, SPC16, TMDE21]. productivity [AN17, AN18, GWA15, HOJ15]. products [BPKZ16, KGW13, ŠW12]. profile [MCL^{+}18, MPM^{+}20]. profiling [HBH12]. Program [GAA18, RBM16, BMC18, KK14, KFYI13, MM14, PA16, San14, Sir18, WZ17, WSB14, ZKK13]. programme [MFA12]. programming [AAJE13, yCP12, LLLL21]. programs [MdlRT19]. progressive [PBSA17]. Project [AN18, AHA17, ADH20, Ano21a, Ano21t, AN17, CMQ^{+}20, CLW22, Dow14, GBRHO14, MGUA14, MMJB16, MM14, NMA20, PCP18, PCMCAC14, Pri12, RGS^{+}15, RBE^{+}21, SDWA20, SSB^{+}20, dSPF^{+}12]. projects [AAM^{+}20, AHA17, Ano21s, BM12a, BCSS18, CKMN15, Cap13, CN12, DFGP18, GKS^{+}12, JWP^{+}16, JMCC18, JRCMWE23, KC22, LMdVT^{+}12, dSLBF15, NPLP12, PGAP17, PPdAPP21, PNL12, QB21, RCL19, SSM21, SAC21, Yag18, dSVB^{+}22]. PROMETHEE [El 18]. promotion [TWFF18]. prone [AAYK20, EAK13]. proneness [QT22, ZHC^{+}18, ZLW22]. propagation [BKZ13]. properties [KBHG17, SGD^{+}13]. property [CNS14]. Propinquity [Pri12]. Proposal [OKA14]. proposed [PGCMA15, SDWA20]. prospective [AMA^{+}20]. prospects [MCSGBSA20]. Protecting [IZJG18, KKW18]. protocols [BAAO^{+}20]. provenance [BNBA21]. provide [UWKRS21]. provider [VvV15]. proximity [SAK17]. Pthreads [FZW^{+}18]. published [LSBV17]. pull [Ano21s]. Purity [NSDD17]. Python [HZ23, SAD22, TFAL21]. Python/C [HZ23].

qualitative [EAD17, LHR18, SH23]. Quality [GKS^{+}12, RMP^{+}20, ST23, YT22, CN12, GLL^{+}21, HBH12, JWP^{+}16, LC23, LMOLG23, LML^{+}18, MRJ20, MDV^{+}18, MAL^{+}13, NBD17, Pot16, SSA15, UFF18, VMSH13, WWKC18, YFL^{+}22, ZKZH14]. quality-centric [NBD17]. quantification [ADCG12]. Quantifying [EKV15, Bol13, EAK13]. Quantitative [Ano21t, Gal20, LHR18]. quantum [HMB22]. question [MKPB23]. questions [Ano21c]. quickly [COGFCA22].

R [MCSGBSA20, Bor16, vdGK12]. R/3 [vdGK12]. radial [Ano21t]. radius [Ano21u]. radius-synthetic [Ano21u]. railway [BDS^{+}22]. ranking [SVSC20]. rate [HSvG17]. rating [DGVP21, Jun12]. Rationale [AEC23]. re [SG17, BCD13]. re-assessing [SG17]. readability [SLVOP18]. Readiness [AMH^{+}20, RYA^{+}21, NMA20, SG20]. real [BO21, SDWA20]. real-life [SDWA20]. Reality [LP17]. realize [DHH^{+}16]. Realizing [NGvD13]. really [GKS^{+}12]. reasoning [AIA20]. recognition [GDAG13]. recommend [CdSdAM16]. recommendation [GLL^{+}21, LLLL21, OKÓ^{+}17, WWQ^{+}21]. recommendations [Raj18b]. recommending [XLWZ15]. reconfiguration [HMB^{+}18]. record [AM16]. recovered [DMT13]. recovering [PCdGCP13]. recovery [BCD13, CDO^{+}13, QBO^{+}13, ZPFV14]. reduce [KTHS16]. reducing [ABS^{+}23]. Reengineering [FKW13, Ano13, CDF13]. Refactoring [KTHS16, AAJE13, CKS18, GJ13, GEK14, GKHE18, MRJ20, MAMN19, OKÓ^{+}17]. refactorings [CKS18]. refinement [DFGP18]. Reflecting [NPLP12]. reflections [BABW12, RTCL16]. reflexion [Kos13]. Regression [UY13, AMH15, AL15, IAA18b, KMB17, MKPB23]. regression-based [IAA18b]. regressions [BMZP14]. regular [PPPdG20]. regulated [DT18]. relate [MAD^{+}14]. related [KMP22, TBP^{+}17]. relating [RGS^{+}15]. relation [AEC23, DRM21, SAPF22]. relations [DGVP21]. relationship [AN17, KCH^{+}17, LSBV16, LSBV17, SP22a]. relative [BPKZ16]. Release [CNSR23]. releases [JPvdW^{+}17]. relevance [HCL^{+}17]. relevant [PJS^{+}17, ZCL^{+}18]. reliability [Ano21q, BDG^{+}23, CS19, KHZ21, Wan21]. reliable [BW18]. remediation [TFAL21]. repair [WWQ^{+}21]. repeat [DRM21]. replay [AM16]. replicated [FQA17]. Report [TKS^{+}14, AÖAG16, BDS^{+}22, KK14, PMT20, SV20, TP21, WWKC18, ZTGG16]. reported [DGVP21]. reports [ATR^{+}17, RHD^{+}16]. repositories [AIA20, CdMSNdCM^{+}14, JGHA14, TUF20]. representation [SLL^{+}21]. representatives [PJ16]. Reproducible [GS23]. reproduction [NHLTL17]. request [CdMSNdCM^{+}14]. requests [Ano21s, HZ14, KGPH12]. required [LP17]. Requirement [AKKM20, ASK^{+}21, AÖAG16, VMCN20, YAMM22, ZWK18]. requirement-oriented [ZWK18]. Requirements [DFGP18, ARD^{+}12, AMA^{+}20, AMH^{+}20, ASACG13, AL17, AHC21, BCPM20a, CKG15, GGPT12, HZ14, HBM^{+}16, JRSS13, JN12, KLWA20, KA20, KST14, KMH^{+}15, LHA^{+}16, OFAT^{+}18, PGCMA15, PVA^{+}21, SDWA20, SL20, TNLJ^{+}17, WCS14, WKB^{+}15, dGNA^{+}16, dGNFAT17]. resampling [ZLM^{+}21, ZHC^{+}18]. Research [RPV12, Ano21v, CDF13, CKR18, DFGP18, Dow14, KMP22, MKPB23, MGW^{+}17, RBE^{+}21, ST23, SBS20, SV21, SH23, UWKRS21, WWKC18, YFL^{+}22]. research-driven [WWKC18]. residual [HLWL15]. Resilient [LAF^{+}18]. resistance [CMCG^{+}12]. resolution [DRM21]. resolve [XLWZ15]. resolving [VvV15]. resource [AVC20]. resources [HMMR18]. responsibility [MSB^{+}14]. restructuring [MPSM22]. restructurings [NBCM17]. Results [BMCP16, BDB^{+}18, HSvG17, LSCP16, RM22, dSPF^{+}12]. retention [RCO20]. retraining [ST22]. retrieval [SAK17]. retrospective [CDR17, CDF13, EPG18, HBH12, Kuh15]. reusable [MGUA14]. reuse [AJP14, ADH20, AM16, PGCMA15, RCL19, SG17, dGNA^{+}16, dGNFAT17]. reuse-based [dGNA^{+}16, dGNFAT17]. Reusing [HAN18]. Reverse [OP14, PP14, GAA18, MBMR19, QSM15, PCCRP21]. reverse-engineered [GAA18]. review [AAG21, AJP14, AHC21, BNBA21, BGM18, BDB^{+}18, CGU22, CPSGAA20, GAA18, HHW12, JbAPT18, JW12, KKT12, KA20, MKA20, NBF18, PTK13, PDP18, RA15, RL15, RBG15, SK18, SSB^{+}20, VKC^{+}21, YT22, ZKN^{+}21]. Reviewer [Ano17m]. reviews [BKS^{+}17, GLL^{+}20, LLLW18]. rewards [SV21]. Rich [WL16]. rigorous [SLMW21]. risk [AAM^{+}20, BMM19, BJBM21, HMMR18, ISK15, LMdVT^{+}12, Lin15, dSLBF15, MLCW22, MMK16, ZIDM20]. risk-enabled [MLCW22]. Risks [BJBM21, Nia12, ZIDM21]. roadmap [ALR^{+}22, Ano21v, CKR18, FMC^{+}15, JPvdW^{+}17, KKN^{+}19]. robotic [JRCMWE23]. Robula [LSRT16]. robust [AKM^{+}21, LSRT16]. role [BCD16]. role-based [BCD16]. root [RF16]. ROS [LGNEAO20, LAF^{+}18]. ROS-based [LGNEAO20]. Rule [KST14, LMdVT^{+}12, OCCSBDBP20]. Rule-based [KST14, LMdVT^{+}12, OCCSBDBP20]. rules [yCP12, DMS12, HBG21, PJW17]. run [THP^{+}16]. run-time [THP^{+}16]. runs [SPC16]. runtime [SPC16].

S# [LHR18]. SaaS [MSK18]. Safe [BMS21, BW18, TMDE21]. safety [BM13, Gal20, GNIS18, JN12, LGNEAO20, LHR18, MBMCP18, MSBE20, PGM21, RGSR20, ST22, VMCN20, Wal18]. safety-critical [JN12, MSBE20, ST22, Wal18]. safety-oriented [Gal20]. same [ADCG12, CMQ^{+}20]. sampling [Ano21u]. SANER [SA17]. SAP [vdGK12]. Savalan [AFK22]. scalable [ACM^{+}14, BAAO^{+}20]. scale [ASA^{+}20, Kos14, PRNR14, SSM21, SSL15, WKB^{+}15]. Scaled [KMP22, TST17, TST17]. Scaling [KHR18, AAM^{+}21]. SCAM [KLG17, CLC14]. SCAN [MAD^{+}14]. scenarios [ISK15]. schema [ZPFV14]. scheme [KBM^{+}12, MK22]. science [JSI^{+}22, MDSSMS15, SV21]. Scorecard [dH13]. Scorecard-based [dH13]. scripted [SGAP14]. scripts [BDS^{+}22]. SCRUM [SDWA20, ÇD19, GP13, PGAP17, PCMCAC14]. Search [FTW16, GEK14, GKHE18, JKIU22, KÓ16, SKWD15]. Search-based [FTW16, GEK14, JKIU22, KÓ16, SKWD15]. Searching [CdSdAM16]. section [CT13, HH12, KR15]. sector [Ano21b, KO13, SSB^{+}20]. secure [BW18, BMS21]. Security [AYB^{+}18, BNBA21, CYK20, MSBE20, MSK18, ZIDM21, CG14, Gal20, GLL^{+}21, GNIS18, KKK^{+}21, LSFG20, PGM21, QHY^{+}22, UFF18, WSB^{+}21, ZIDM20]. Security- [MSBE20]. Security-aware [MSK18]. security-informed [Gal20]. Segmentation [WPVS14]. segments [MAD^{+}14]. Selected [BMP13, MBMCP18, KDK13]. Selecting [TUF20]. Selection [RAM^{+}22, Ano21r, CDF13, KMB17, aRL21, MLM16, OP14, PP14, UC12, UY13]. self [GWA15, MKMW18]. self-assessed [GWA15]. self-driving [MKMW18]. semantic [Ano21c, Ano21p, KKW18, YF22, YYF^{+}22, YZT22]. semantics [WZ17]. sensitive [ARSH16]. sensitivity [Jun12, LHA^{+}16]. sentence [MLM16]. sequence [GAA18, LCMCPA15, SNP22, YYF^{+}22]. series [HH12, RHL15]. Service [BMCP16, UC12, ARD^{+}12, AMB18, AMB22, Ano21p, AYB^{+}18, BCPM19, KO13, LM18, MBMCP18, MKMW18, NGvD13, RLR21, Sch12, UTZ13, VvV15, WRC^{+}13, ZDV22]. service-oriented [ARD^{+}12, RLR21]. Services [BM12b, ARD^{+}12, BH16, EKP18, FAT22, PCdGCP13, PGM21]. set [LHEG18]. setting [ETT^{+}21, HD17]. settings [HLP15, LS19]. SFLA [GA20].

SFLA-based [GA20]. shared [BMCP14c]. sharing [BABW12, CCUP21, GS12, LHEG18, VvV15]. similar [SBPS12]. Similarities [Cap13]. similarity [PCCP14]. Simple [ARSH16]. Simplified [YZT22, COGFCA22, MFCP14, PB17]. Simplifying [NBCM17]. Simulation [CLMZ13, BHZK12, BKS^{+}17, GP16, OCCSBDBP20, ZRB^{+}14]. simulator [MVP12]. since [OKA14]. single [KRR^{+}16]. single-objective [KRR^{+}16]. sites [AHC21]. situation [PKY14]. situation-aware [PKY14]. size [LLT20]. sized [LCMCPA15]. sizes [SP22b]. Skills [SRSEGA14, ISK15, MDSSMS15, RDD^{+}12, TWFF18]. slicing [ACM^{+}14, HF12, KFYI13, WZ17]. SLOC [LSBV16, LSBV17]. SLR [ASA^{+}20, KKK^{+}21]. small [Abu16, ABG18, BAB^{+}22, CG14, ECRJFZ13, ECRJFZ16, GPM^{+}12, LS19, LCMCPA15, MPS^{+}16, MPM^{+}20, PRGP12, SG20]. smaller [BPKZ16]. smart [AVC20, PKY14]. SME [BABW12]. smell [AAG21, MRJ20, RA15]. smells [MAMN19, OKÓ^{+}17, SLHS18, SP22a, SAPF22]. SMEs [Abu16, CO13, SS18]. Snapshot [BMCP14c, TPB^{+}17]. SOA [RL15]. Social [AHC21, MSB^{+}14, AY20, CPMS^{+}14, ECRJVA^{+}18, NMA^{+}17]. sociocultural [GSB^{+}20]. Software [AAG21, AL17, Ano21u, BM12b, BMCP14b, BMCP16, DS17, FKW13, HOJ15, IK17, aRL21, MPS^{+}16, McQ12, MTG22, MAL^{+}13, PF18, PCdGCP13, PDP18, PBDJ16, Rya16, TFAL21, VBMH22, WWKC18, ZRB^{+}14, ZHC^{+}18, ZLW22, AAM^{+}20, AIA20, ASA^{+}20, ARD^{+}12, ATR^{+}17, ALK^{+}17, AY20, AHA17, AMH^{+}20, ASK^{+}21, AKM^{+}21, AÖAG16, ASG18, ABB14, AM20, AHKL20, ADH20, ALK^{+}20, AG21, ABG18, AL15, AI21, Ano13, Ano21b, Ano21a, Ano21q, Ano21t, Ano21v, AFK22, AAA13, AN18, Bab12, BW18, BM12a, BHZK12, BAB^{+}22, BH16, BJBM21, BSEA14, BPKZ16, BMP13, BMCP14a, BCPM19, BCPM20a, BCPM20b, BSB16, BGM18, Bor16, BDB^{+}18, BCSS18, Cao12, CDF13, CDO^{+}13, CKR18, CPR15, CC13, CGU22, CN12, CdMSNdCM^{+}14, ÇD19, CNSR23, CAKA21, CKG15, CS19, CS13, CKS18, CGN17, CCUP21, CO13, CLMZ13, Dow14, ETT^{+}21]. software [EIA19, EAK13, ECRJFZ13, ECRJFZ16, FDM^{+}18, FQA17, FPA17, FHS13, GPM^{+}12, GP16, GFK^{+}20, GP13, GS18b, GA20, GKHE18, Ghe17, GJC^{+}18, GdOK18, GBRHO14, GKS^{+}12, GWA15, GD21, HAN18, HH12, dH13, HB16, HSvG17, HAS21, HLP15, HBM^{+}16, HIA19, HLWL15, HCL^{+}17, IAA18a, IAA18b, II17, JGHA14, JWP^{+}16, JPvdW^{+}17, JW12, JMCC18, JN12, KGPH12, KC22, KTYK16, KKT12, KÓ16, KKN^{+}19, KA20, KHZ21, KGW13, KCH^{+}17, KR15, Kos21, KMP22, KRG^{+}13, Kru13, Kuh15, KFT16, KFYI13, KO13, LSW13, LMB12, LP17, eLAA^{+}23, LMdVT^{+}12, LCMCPA15, LFT19, LCM^{+}15, LMC16, LYL^{+}12, LLSA22, Lin15, LLLW18, LHEG18, dSLBF15, LHIE18, MKA20, MBD^{+}22, MVRP20, MGUA14, MPB16, MBMCP18, MMC14, MDSSMS15, MGN18, MAD12, MFCP14, MA20, MP16, MAMN19, MPM^{+}20, NXC13, Nia12, Nia15, NMA^{+}17, NSDP15, NPLP12]. software [ND20, NBF18, OEC17, OMTR12, OKA14, OB20, Ost14, ÖTM18, PL14, PGCMA15, PJW17, PA14, PBSA17, PKY14, PRNR14, PPdAPP21, PRGP12, PNL12, PK14, PSMM19, PUWS21, Pri12, QVR^{+}21, RPZ14, RYA^{+}21, RLMC19, Raj18b, RCL19, RMOP20, RMAB15, RK18a, RK18b, RCO20, RF16, RTCL16, RPV12, RBE^{+}21, RBG15, RLTTC14, RMCLGAM14, RHL15, RFM16, ST23, SSM21, SDWA20, SKWD15, SVSC20, SK18, SRSEGA14, SKTW14, SGSGYO19, SAM13, SPC16, SSA15, SKK^{+}18, SJA^{+}14, SS18, SP22b, SAEH16, SLL^{+}21, SG20, ŠW12, SV20, SBPS12, SAC21, SG17, SH23, TP21, TGE17, TUF20, TWL^{+}21, TWFF18, THP^{+}16, THP13, UWGP15, VKPvV13, VKC^{+}21, Wan21, Yag18, YZT22, YAMM22, YFL^{+}22, YT22, ZKJ12, ZKZH14, ZWK18, ZWY^{+}20, dSPF^{+}12, dGNFAT17, dSVB^{+}22, GSB^{+}20, HOD^{+}14, OCCSBDBP20, RRDPD^{+}13, CDR12]. software-intensive [AAA13, OB20]. solution [AEC23, ETT^{+}21, GGGBEM18]. solutions [BW18, CdSdAM16, KJ14, KKT12, MVRP20, UWGP15]. Source [BHLO16, AVC20, Ano21s, AM16, Cap13, CLC14, CNSR23, DRM21, DRGP13, HZ14, KC22, KMH^{+}15, KFYI13, eLAA^{+}23, LGS^{+}17, MLM16, NXC13, NBCM17, QB21, RCO20, SAPF22, SAK17, Wan21, YT22, KLG17]. Sources [WL16]. sourcing [UC12]. Space [PBDJ16]. Spanish [GP13]. spatial [SAK17]. Special [Ano13, BMP13, CKR18, CT13, FKW13, KLG17, KDK13, KOPR17, MC14, MAD12, NSBR12, PF18, PKBT16, PP14, RPZ14, BMC18, BBD^{+}22, CLC14, HAS21, KÓ16, KR15, MTG22, MP16, SA17, RBM16]. specific [EMRF14]. specification [GDJV18, dGNA^{+}16]. specifications [Bol13, BGM20, PCL21, TG20]. spectrum [ZKK13]. spectrum-based [ZKK13]. speech [GDAG13]. speed [MPB16]. SPI [BMCP15, ECRJFZ16, KBM^{+}12, LS19, MFA12, MSB^{+}14, SS18, UGI^{+}14]. SPICE [Gal20, ISK15, MKMW18, NGAEE22, SVK^{+}14, SBPS12, VNM14]. SPICE-compliant [Gal20]. spin [Dow14]. spin-out [Dow14]. splitting [GDAG13]. Spreadsheet [KKW18, ADFT16]. spreadsheets [KKW18]. sprint [EPG18, PCMCAC14]. Square [eLAA^{+}23]. srcDiff [DCVM20]. srcSlice [ACM^{+}14]. stable [HG13]. stage [PBSA17]. staged [BDGR12]. stakeholder [BHZK12]. stakeholder-oriented [BHZK12]. stakeholders [LYL^{+}12]. standard [KK14]. standardized [IAA18a]. standards [ASACG13, KO13, PGM21]. standards-based [ASACG13]. State [BGM20, GS18b, PA14]. state-of-the-art [GS18b]. states [AHA17]. static [ACM^{+}14, BCD16, CNS14, WRC^{+}13, SJG^{+}14]. statistical [EPG18, NL18, TKK^{+}21]. stealth [BRDL21]. stochastic [BEB21]. stores [LLLL21]. story [NEU^{+}19]. strategic [OB17]. Strategy [OB17, ABQO14, BM12a, PCL21, PHJ15, RL14]. stream [KPG14]. structural [AAM^{+}21, AHKL20, EKG18, GA20, JKIU22, SS18, eLAA^{+}23]. structure [WCDY22, YYF^{+}22]. structures [FMR13]. student [MDSSMS15]. studies [BABW12, CAKA21, MKA20]. study [AAM^{+}20, ALK^{+}17, AY20, AHA17, AMA^{+}20, ABG18, AAYK20, Ano21b, BKZ13, BO21, BCSS18, CdMSNdCM^{+}14, CNSR23, CAKA21, CV21, CNS14, DRM21, EAD17, ECRJFZ16, EKV15, FQA17, GFK^{+}20, GAA^{+}22, GP13, GdOK18, HD17, HLP15, HZ23, JPvdW^{+}17, KTP22, KFT14, KFYI13, KO13, LSW13, LMB12, LS19, LGS^{+}17, LLSA22, LHIE18, LLJ^{+}14, MPSM22, MGUA14, MPB16, MDV^{+}18, MdlRT19, NH14, NXC13, Nia12, Nia15, NMA^{+}17, OCCSBDBP20, PMT20, PCP18, PCCRP21, PCMCAC14, QHY^{+}22, QT22, RGS^{+}15, RMOP20, RL14, RBE^{+}21, RHD^{+}16, RRDPD^{+}13, SP22a, SDG21, SJA^{+}14, SKR15, TFAL21, TGE17, TFPB14, TWL^{+}21, TMDE21, TBP^{+}17, VvV15, WPXZ13, WSB^{+}21, YAMM22, ZIDM21, ZKJ12, ZKZH14, dSPF^{+}12, dSVB^{+}22]. Studying [GK13, GBRHO14, KRG^{+}13, Ost14]. styles [CNPR13]. Success [RK18b, YAMM22, ASK^{+}21, BDB^{+}18, KHR18, LYL^{+}12, MSB^{+}14, Nia15, RGS^{+}15, Yag18]. success-critical [LYL^{+}12]. Successful [CKMN15]. Successfully [Dow14]. suffix [Kos14]. suite [EAK13, UY13]. summarization [MLM16]. summarizing [CKS18]. supplier [KJ14, NRB13]. Support [IAA18b, GP16, HB16, KTHS16, KL18, KMH^{+}15, LHEG18, dSLBF15, NSDP15, NGAEE22, RMAB15, SKTW14, Sch12, THP13, WRC^{+}13, ZCL^{+}18, ZPW12, dGNFAT17]. supported [Bab12, NRB13]. Supporting [CKG15, KLC^{+}14, Stu12, VKPvV13, ADH20, MSB^{+}14]. survey [AMB22, DRGP13, FMR13, LM18, MPS^{+}16, MAMN19, SK18, SAM13]. survivability [ZBGH22]. survival [SPC16]. Sustainability [OFAT^{+}18, BAB^{+}22, CGN17, LP17, SL20]. sustainable [PJW17, RK18a, RK18b, SK18]. swarm [MSK18]. symbolic [CsZZ^{+}13]. Symposium [KDK13]. synchronicity [NPLP12]. syndrome [ZDV22]. syntactic [DCVM20, NBCM17]. syntax [YZT22]. Synthesis [FAT22]. synthesize [GRK17]. synthesizing [WCDY22]. synthetic [Ano21u]. SysML [MCL^{+}18]. System [BMCP16, ASACG13, AAA13, BCPM19, CYK20, Kos14, aRL21, MBMCP18, MBMR19, NH14, PRNR14, SGD^{+}13, THP^{+}16, VKPvV13, dH13]. Systematic [KA20, AAG21, AJP14, AHC21, BNBA21, BGM18, BDB^{+}18, CGU22, CdMSNdCM^{+}14, GAA18, JbAPT18, JW12, JPW^{+}17, LLSA22, LHIE18, MKA20, MMPN20, MdlRT19, PTK13, PDP18, RL15, RBG15, RRDPD^{+}13, SK18, VKC^{+}21, YT22, ZIDM21, ZKN^{+}21, dSPF^{+}12, dSVB^{+}22]. systematization [Kuh15]. SystemC [NL18]. Systems [BM12b, BBD^{+}22, KDK13, dSLBF15, MCL^{+}18, RRDPD^{+}13, SDG21, ALR^{+}22, BW18, BM12a, BM13, BZ21, BGV18, BMS21, BGM18, BO21, CPR15, CYK20, DLM23, DP15, FLGB22, FHS13, GAA^{+}22, GdSC21, HF12, HAS21, HBM^{+}16, HMMR18, JN12, KRG^{+}13, LSW13, LdONB22, MLCW22, MSBE20, MA12, OB20, OCCSBDBP20, PMT20, PCCRP21, PRNR14, PDP18, RPZ14, RGSR20, RMP^{+}20, SJA^{+}14, SLMW21, ST22, TNLJ^{+}17, TMDE21, UFF18, Wal18, YF22, ZIDM20, ZIDM21, ZPFV14, vdGK12]. Systems-of-systems [MCL^{+}18, GdSC21].

table [WSB14]. tailoring [ABQO14, Gal20, Sir18]. TALAIA [MMJB16]. target [CLW22, RAS^{+}13]. task [FPA17, II17, LMdVT^{+}12, ZKN^{+}21]. tasks [CdSdAM16, WPXZ13]. Taxonomy [RLMC19, RLTTC14, AAJE13, DRGP13, KS20]. Taxonomy-based [RLMC19]. teaching [De 15]. Team [SSM21, ECRJVA^{+}18, GJC^{+}18, GGJ20, RMOP20, SRSEGA14]. Team-external [SSM21]. teams [ASA^{+}20, BCSS18, KTP22, PA14, Pri12]. teamwork [RLTTC14]. Technical [SN14, CN12, CWBC17, HBM^{+}16, KTHS16, MB17, TFAL21, ZBGH22]. technique [Ano21u, EKG18, GJC^{+}18, KMB17, RAS^{+}13]. techniques [AAG21, Ano21r, GDAG13, MMPN20, MBMR19, MAMN19, PPPdG20, RA15, dGNA^{+}16]. technology [BDB^{+}18, CPSGAA20, HD17]. TeLESM [OB17]. temporal [BGV18, LWY^{+}16]. term [HSvG17]. terminological [HBG21]. test [BLGZ18, BDS^{+}22, EMRF14, GAA^{+}22, GA20, GRK17, Hou14, JKIU22, KMB17, LdONB22, MMPN20, PPPdG20, QBO^{+}13, RAM^{+}22, SN14, SBPS12, UY13, UIK20, vdGK12, SAD22, SBPS12]. test-and-fix [Hou14]. test-to-code [QBO^{+}13]. Testing [BPKZ16, MTG22, MdlRT19, AMH15, ABS^{+}23, AMB18, AMB22, BDG^{+}23, BW22, CS19, DLM23, FDM^{+}18, FQA17, GP16, GFK^{+}20, GAA^{+}22, HH12, JKIU22, JSI^{+}22, JRCMWE23, KMB17, LSRT16, LYL^{+}12, LdONB22, MLCW22, MKPB23, QHY^{+}22, RLMC19, SKWD15, SGAP14]. TestSPICE [SN14]. text [DRW14, PJS^{+}17, ZTGG16]. text-based [DRW14]. textual [HBG21]. thef [PF18]. their [DGVP21, FMR13, SN14, VKPvV13]. them [BO18]. themes [BMCP14a]. theory [BPKZ16, NPLP12]. There [TPB^{+}17]. these [MVP12]. Things [BNBA21]. Thinking [SDG21, RTCL16]. Three [OB17, BCSS18, PJ16]. Three-Layer [OB17]. thresholds [Sha15]. throughout [AÖAG16, KKW18]. TIDIER [GDAG13]. tier [CLAD17]. Time [RHL15, ABS^{+}23, HH12, HZ14, LHA^{+}16, THP^{+}16, WKB^{+}15]. time-series [HH12]. timed [SNP22]. times [DRM21]. TIPA [CRP14]. TMAP [PJS^{+}17]. Token [WLL^{+}23]. Token-based [WLL^{+}23]. tolerance [LAF^{+}18]. tomorrow [RPZ14]. tool [ABB14, ADFT16, CsZZ^{+}13, CPMS^{+}14, GPM^{+}12, NSDP15, SV21, ZWW^{+}20]. tooled [KLC^{+}14]. tools [Bor16, MDSSMS15, NPLP12, PJ16]. top [AAA13]. top-down [AAA13]. Topics [BM12b]. topological [HMMR18]. TOPSIS [KKK^{+}21, RAM^{+}22]. Toshiba [OKA14]. trace [MAD^{+}14]. traceability [AÖAG16, CDO^{+}13, CC13, CAKA21, DHH^{+}16, HZ14, KMH^{+}15, QBO^{+}13, RBFM15, TWL^{+}21, TRC^{+}17]. traces [NHLTL17]. tracing [San14]. traditional [ECRJVA^{+}18, PSMM19, dGNA^{+}16, BW22]. traffic [HMB^{+}18]. train [MVP12]. trajectory [JPW^{+}17]. transaction [QB21]. transfer [BDB^{+}18]. transferring [ŠW12]. transformable [ZKJ12]. Transformation [SNP22, BO21, ECRJVA^{+}18, OB20, RRL22, ST23, SKWD15, Sha15, SGD^{+}13]. transformation-Action [ST23]. transformations [FTW16, JKIU22, NBCM17, PB17, WSB14]. Transition [HOD^{+}14, ARD^{+}12]. transitioning [Dow14]. traps [BO18]. tree [YZT22, WG18]. trees [Kos14]. Trenches [MTG22]. trends [RHL15]. Trustworthiness [ZWK18]. trustworthy [BHZK12, Cao12, FDM^{+}18, MA12, ZKJ12]. TuneR [Bor16]. tuning [Bor16]. tutorials [ND20]. two [BABW12, RTCL16]. types [PJ16].

UML [MMPN20, MAMN19]. Uncovering [CGN17, MB17]. understandability [DCVM20]. Understanding [DRM21, ECRJFZ16, JPW^{+}17, LLSA22, PCMCAC14, McQ12, NXC13]. unified [CS19]. unifying [SGAP14]. unit [GRK17, OCCSBDBP20, UIK20]. unit-level [UIK20]. university [De 15]. unofficial [SG20]. unsupervised [SLL^{+}21]. unusual [GFKH18]. update [BJBM21, CKG15]. Updating [GLL^{+}20, SAM13]. upgradation [Ano21q]. usage [Ano21v, CdSdAM16, EAD17, GRK17, NPLP12, SWT^{+}16, ZKN^{+}21]. Use [TG20, AN17, BGM18, DS17, KFT16, MKA20, MMC14, MKMW18, MBMR19, RCL19, ZWW^{+}20]. used [DT18]. User [SH23, VBMH22, DGVP21, GLL^{+}20, GD21, KLWA20, Lin15, LMOLG23, XBS^{+}21]. User-centered [SH23, GD21]. user-reported [DGVP21]. users [IZJG18]. Uses [RRDPD^{+}13]. Using [CPMS^{+}14, HCL^{+}17, MKPB23, MBMR19, RK18b, SAC21, AKM^{+}21, AMB22, AI21, Ano21p, APL^{+}14, BM13, BCC^{+}15, BKS^{+}17, BEB21, BAAO^{+}20, BCD14, CZY^{+}22, CKP15, CNPR13, Dow14, El 18, GEK14, GK13, GDAG13, JWP^{+}16, KLC^{+}14, KS20, Kos14, KRG^{+}13, KFYI13, KO13, eLAA^{+}23, LAF^{+}18, MGUA14, MRJ20, MK22, MSK18, MMPN20, MCL^{+}18, MPM^{+}20, NBD17, PVA^{+}21, PGAP17, QHY^{+}22, QT22, RAM^{+}22, SJG^{+}14, SNP22, SKK^{+}18, Sha15, SAEH16, SGD^{+}13, SKR15, UY13, Yag18]. utilization [YFL^{+}22]. UWA [BCD13].

V [KFT16, LC23]. V-model [LC23]. V-Modell [KFT16]. valid [ND20]. Validating [VMCN20]. validation [AG21, BGM20, GSB^{+}20, eLAA^{+}23, MMK16, OCCSBDBP20, RLMC19, VMSH13, BAAO^{+}20]. VALO [RGS^{+}15]. value [Ano21r, KGW13, KPG14, SG17]. values [LFT19]. variability [KFT16, LdONB22, NH14, NBD17]. variables [AAYK20]. variants [BDS^{+}22, VHBK22]. variation [EIA19]. various [BMCP14b]. VBA [ADFT16]. VBA-based [ADFT16]. vector [IAA18b]. vehicles [HMB^{+}18]. velocity [ASG18, PCMCAC14]. vendor [AMA^{+}20, KHZ21]. vendors [IK17, RK18a, RK18b, SK18]. verbal [FPA17]. verification [BEB21, BGV18, FZW^{+}18, LWY^{+}16, NL18]. verified [GDJV18]. verifying [CZY^{+}22]. versa [ARD^{+}12]. Version [MLCW22]. Version-based [MLCW22]. versions [ND20]. versus [BW22, MDSSMS15, dGNA^{+}16]. very [ACM^{+}14, MPM^{+}20, PRNR14]. via [AM16, CDO^{+}13, MLM16, NBCM17, QB21, Stu12, WWQ^{+}21]. vice [ARD^{+}12]. victimized [IZJG18]. view [ALK^{+}20, CDF13, THP13]. viewpoint [CPSGAA20, MCL^{+}18]. viewpoint-based [MCL^{+}18]. views [AAA13]. violations [yCP12]. visibility [VvV14]. VISSOFT [PF18]. Visualization [PF18, LHIE18, MLCW22, MGN18]. Visualizing [WG18]. vocabulary [BL15]. vs [DS17]. VSE [LSFG20]. vulnerable [SAC21].

warehouses [SZ12]. warehousing [Stu12]. waste [KPG14, RZL^{+}15]. wasted [LHA^{+}16]. ways [GGJ20]. Web [KDK13, Ano21p, DMT13, LSRT16, LMOLG23, XBS^{+}21, ZDV22, BCC^{+}15, BCD13, LB13, ND20, PCdGCP13]. WebKit [SJG^{+}14]. website [VMSH13]. weight [RAS^{+}13]. weighted [AL15, ST22]. weighting [EKG18]. well [RGSR20]. well-being [RGSR20]. Where [RZL^{+}15, dSVB^{+}22, BO18, RPV12]. WIFLF [CLW22]. Wikipedia [Cap13]. wild [DBP22]. windows [AL15, CsZZ^{+}13, MK22]. winners [LYL^{+}12]. within [Gal20]. without [CLMZ13]. Work [CSGR17, CLMZ13, RMCLGAM14]. work-in-process [CLMZ13]. workers [CPMS^{+}14]. Workflow [LWY^{+}16, BEB21, THP13, WCDY22]. Working [OP14, PF18, PP14, GGJ20, SRSEGA14]. world [RM22, Rya16]. WSE [KDK13].

XPath [LSRT16]. XT [KFT16].

year [CDR17, MMJB16].

Z [Bol13].

References

Arias:2013:TAC

[AAA13] Trosky B. Callo Arias, Pierre America, and Paris Avgeriou. A top-down approach to construct execution views of a large software-intensive system. Journal of Software: Evolution and Process, 25(3): 233–260, March 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

AbuHassan:2021:SSD

[AAG21] Amjad AbuHassan, Mohammad Alshayeb, and Lahouari Ghouti. Software smell detection techniques: a systematic literature review. Journal of Software: Evolution and Process, 33(3): e2320:1–e2320:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Alshayeb:2013:ETR

[AAJE13] Mohammad Alshayeb, Hamdi Al-Jamimi, and Mahmoud O. Elish. Empirical taxonomy of refactoring methods for aspectoriented programming. Journal of Software: Evolution and Process, 25(1): 1–25, January 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abioye:2020:TOB

[AAM^{+}20] Temitope Elizabeth Abioye, Oluwasefunmi Tale Arogundade, Sanjay Misra, Adio T. Akinwale, and Olusola John Adeniran. Toward ontology-based risk management framework for software projects: an empirical study. Journal of Software: Evolution and Process, 32(12): e2269:1–e2269:??, December 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abrar:2021:FMS

[AAM^{+}21] Muhammad Faisal Abrar, Sikandar Ali, Muhammad Faran Majeed, Sohail Khan, Muzamil Khan, Hamid Ullah, Mushtaq Ahmad Khan, Samad Baseer, and Muhammad Asshad. A framework for modeling structural association among De-Motivators of scaling agile. Journal of Software: Evolution and Process, 33(8): e2366:1–e2366:??, August 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Aman:2020:ESA

[AAYK20] Hirohisa Aman, Sousuke Amasaki, Tomoyuki Yokogawa, and Minoru Kawahara. Empirical study of abnormality in local variables and its application to fault-prone Java method analysis. Journal of Software: Evolution and Process, 32(4): e2220:1–e2220:??, April 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Alegria:2014:ATA

[ABB14] Julio A. Hurtado Alegría, María Cecilia Bastarrica, and Alexandre Bergel. Avispa: a tool for analyzing software process models. Journal of Software: Evolution and Process, 26(4):434–450, April 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Almomani:2018:ESS

[ABG18] Malek Ahmad Almomani, Shuib Basri, and Abdul Rehman Gilal. Empirical study of software process improvement in Malaysian small and medium enterprises: the human aspects. Journal of Software: Evolution and Process, 30(10): e1953:1–e1953:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Alegria:2014:MBP

[ABQO14] Julio A. Hurtado Alegría, María Cecilia Bastarrica, Alcides Quispe, and Sergio F. Ochoa. MDE-based process tailoring strategy. Journal of Software: Evolution and Process, 26(4): 386–403, April 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

AlGhamdi:2023:TRT

[ABS^{+}23] Hammam M. AlGhamdi, Cor-Paul Bezemer, Weiyi Shang, Ahmed E. Hassan, and Parminder Flora. Towards reducing the time needed for load testing. Journal of Software: Evolution and Process, 35(3): e2276:1–e2276:??, March 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abushama:2016:PSP

[Abu16] Hisham M. Abushama. PAM–SMEs: process assessment method for small to medium enterprises. Journal of Software: Evolution and Process, 28 (8):689–711, August 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Alomari:2014:SVE

[ACM^{+}14] Hakam W. Alomari, Michael L.

Collard, Jonathan I. Maletic, Nouh Alhindawi, and Omar Meqdadi. srcSlice: very efficient and scalable forward static slicing. Journal of Software: Evolution and Process, 26(11): 931–961, November 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abran:2012:MQS

[ADCG12] Alain Abran, Jean-Marc Desharnais, and Juan Jose Cuadrado-Gallego. Measurement and quantification are not the same: ISO 15939 and ISO 9126. Journal of Software: Evolution and Process, 24(5):585– 601, August 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Amalfitano:2016:ETC

[ADFT16]  Domenico  Amalfitano,

Vincenzo De Simone, Anna Rita Fasolino, and Porfirio Tramontana. EX-ACT: A tool for comprehending VBA-based Excel spreadsheet applications. Journal of Software: Evolution and Process, 28(6): 483–505, June 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2020:HDP

[ADH20] Nazakat Ali, Horn Daneth, and Jang-Eui Hong. A hybrid DevOps process supporting software reuse: a pilot project. Journal of Software: Evolution and Process, 32 (7):e2248:1–e2248:??, July 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Altamimi:2023:MRA

[AEC23] Emran Altamimi, Abdullah Elkawakjy, and Cagatay Catal. Metamorphic relation automation: Rationale, challenges, and solution directions. Journal of Software: Evolution and Process, 35(1): e2509:1–e2509:??, January 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Arasteh:2022:SMO

[AFK22] Bahman Arasteh, Ahmad Fatolahzadeh, and Farzad Kiani. Savalan: Multi objective and homogeneous method for software modules clustering. Journal of Software: Evolution and Process, 34(1): e2408:1–e2408:??, January 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2021:ECV

[AG21] Asad Ali and Carmine Gravino. An empirical comparison of validation methods for software prediction models. Journal of Software: Evolution and Process, 33(8): e2367:1–e2367:??, August 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2017:MMS

[AHA17] Rehan Akbar, Mohd Fadzil Hassan, and Azrai Abdullah. A meta-model of software development project states behavioral study of software projects. Journal of Software: Evolution and Process, 29(4):??, April 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2021:SNS

[AHC21] Nazakat Ali, Jang-Eui Hong, and Lawrence Chung. Social network sites and requirements engineering: a systematic literature review. Journal of Software: Evolution and Process, 33(4): e2332:1–e2332:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Aivaloglou:2017:PEF

[AHH17] Efthimia Aivaloglou, David Hoepelman, and Felienne Hermans. Parsing Excel formulas: A grammar and its application on 4 large datasets. Journal of Software: Evolution and Process, 29(12):??, December 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2020:FMS

[AHKL20] Sikandar Ali, Jiwei Huang, Siffat Ullah Khan, and Hongqi Li. A framework for modelling structural association amongst barriers to software outsourcing partnership formation: an interpretive structural modelling approach. Journal of Software: Evolution and Process, 32 (6):e2243:1–e2243:??, June 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Amazal:2021:ESD

[AI21] Fatima Azzahra Amazal and Ali Idri. Estimating software development effort using fuzzy clusteringbased analogy. Journal of Software: Evolution and Process, 33(4): e2324:1–e2324:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abnane:2020:FCB

[AIA20] Ibtissam Abnane, Ali Idri, and Alain Abran. Fuzzy case-based-reasoning-based imputation for incomplete data in software engineering repositories. Journal of Software: Evolution and Process, 32(9): e2260:1–e2260:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ahmad:2014:CCA

[AJP14] Aakash Ahmad, Pooyan Jamshidi, and Claus Pahl. Classification and comparison of architecture evolution reuse knowledge — a systematic review. Journal of Software: Evolution and Process, 26(7): 654–691, July 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2020:RCM

[AKKM20] Muhammad Azeem Akbar, Arif Ali Khan, Abdul Wahid Khan, and Sajjad Mahmood. Requirement change management challenges in GSD: an analytical hierarchy process approach. Journal of Software: Evolution and Process, 32 (7):e2246:1–e2246:??, July 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2021:RFC

[AKM^{+}21] Muhammad Azeem Akbar, Arif Ali Khan, Sajjad Mahmood, Ahmed Alsanad, and Abdu Gumaei. A robust framework for cloud-based software development outsourcing factors using analytical hierarchy process. Journal of Software: Evolution and Process, 33(2): e2275:1–e2275:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

AlDallal:2015:ADE

[Al 15] Jehad Al Dallal. Accounting for data encapsulation in the measurement of object-oriented class cohesion. Journal of Software: Evolution and Process, 27(5):373–400, May 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Amasaki:2015:EWM

[AL15] S. Amasaki and C. Lokan. On the effectiveness of weighted moving windows: Experiment on linear regression based software effort estimation. Journal of Software: Evolution and Process, 27(7): 488–507, July 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2017:MRE

[AL17] Naveed Ali and Richard Lai. A method of requirements elicitation and analysis for Global Software Development. Journal of Software: Evolution and Process, 29(4): ??, April 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ahmad:2017:ESP

[ALK^{+}17] Muhammad Ovais Ahmad, Lucy Ellen Lwakatare, Pasi Kuvaja, Markku Oivo, and Jouni Markkula. An empirical study of portfolio management and Kanban in agile and lean software companies. Journal of Software: Evolution and Process, 29(6):??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2020:PVB

[ALK^{+}20] Sikandar Ali, Hongqi Li, Siffat Ullah Khan, Muhammad Faisal Abrar, and Yanhong Zhao. Practitioner’s view of barriers to software outsourcing partnership formation: an empirical exploration. Journal of Software: Evolution and Process, 32 (5):e2233:1–e2233:??, May 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2022:TRI

[ALR^{+}22] Muhammad Azeem Akbar, Víctor Leiva, Saima Rafi, Syed Furqan Qadri, Sajjad Mahmood, and Ahmed Alsanad. Towards roadmap to implement blockchain in healthcare systems based on a maturity model. Journal of Software: Evolution and Process, 34(12): e2500:1–e2500:??, December 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Armaly:2016:PSC

[AM16] Ameer Armaly and Collin McMillan. Pragmatic source code reuse via execution record and replay. Journal of Software: Evolution and Process, 28 (8):642–664, August 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Alhammad:2020:CGS

[AM20] Manal M. Alhammad and Ana M. Moreno. Challenges of gamification in software process improvement. Journal of Software: Evolution and Process, 32 (6):e2231:1–e2231:??, June 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2020:MSI

[AMA^{+}20] Muhammad Azeem Akbar, Sajjad Mahmood, Ahmed Alsanad, Abeer Abdul-Aziz Alsanad, Abdu Gumaei, and Syed Furqan Qadri. A multivocal study to improve the implementation of global requirements change management process: a clientvendor prospective. Journal of Software: Evolution and Process, 32(8): e2252:1–e2252:??, August 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2018:APG

[AMB18] Amira Ali, Huda Amin Maghawry, and Nagwa Badr. Automated parallel GUI testing as a service for mobile applications. Journal of Software: Evolution and Process, 30(10): e1963:1–e1963:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ali:2022:PTS

[AMB22] Amira Ali, Huda Amin Maghawry, and Nagwa Badr. Performance testing as a service using cloud computing environment: a survey. Journal of Software: Evolution and Process, 34(12): e2492:1–e2492:??, December 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abadeh:2015:DBR

[AMH15] Maryam Nooraei Abadeh and Seyed-Hassan Mirian-Hosseinabadi. Delta-based regression testing: a formal framework towards model-driven regression testing. Journal of Software: Evolution and Process, 27(12):913–952, December 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2020:RMR

[AMH^{+}20] Muhammad Azeem Akbar, Sajjad Mahmood, Zhiqiu Huang, Arif Ali Khan, and Mohammad Shameem. Readiness model for requirements change management in global software development. Journal of Software: Evolution and Process, 32 (10):e2264:1–e2264:??, October 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Azzeh:2017:ARB

[AN17] Mohammad Azzeh and Ali Bou Nassif. Analyzing the relationship between project productivity and environment factors in the use case points method. Journal of Software: Evolution and Process, 29(9): ??, September 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Azzeh:2018:PPE

[AN18] Mohammad Azzeh and Ali Bou Nassif. Project productivity evaluation in early software effort estimation. Journal of Software: Evolution and Process, 30(12):e2110:1– e2110:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2013:ESI

[Ano13]  Anonymous.  Editorial:

Special issue: software maintenance and reengineering (CSMR 2011). Journal of Software: Evolution and Process, 25(10): 1061–1062, October 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIa

[Ano14a] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(1):i–iii, January 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIb

[Ano14b] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(2):i–iii, February 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIc

[Ano14c] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(3):i–iii, March 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IId

[Ano14d] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(4):i–iii, April 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIe

[Ano14e] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(5):i–iii, May 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIf

[Ano14f] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(6):i–iii, June 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIg

[Ano14g] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(7):i–iii, July 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIh

[Ano14h] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(8):i–iii, August 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIi

[Ano14i] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(9):i–iii, September 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIj

[Ano14j] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(10):i–iii, October 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIk

[Ano14k] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(11):i–iii, November 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2014:IIl

[Ano14l] Anonymous. Issue information. Journal of Software: Evolution and Process, 26(12):i–iii, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIa

[Ano15a] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(1):i–iii, January 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIb

[Ano15b] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(2):i–iii, February 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIc

[Ano15c] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(3):i–iii, March 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IId

[Ano15d] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(4):i–iii, April 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIe

[Ano15e] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(5):i–iii, May 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIf

[Ano15f] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(6):i–iii, June 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIg

[Ano15g] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(7):i–iii, July 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIh

[Ano15h] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(8):i–iii, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIi

[Ano15i] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(9):i–iii, September 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIj

[Ano15j] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(10):i–iii, October 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIk

[Ano15k] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(11):i–iii, November 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2015:IIl

[Ano15l] Anonymous. Issue information. Journal of Software: Evolution and Process, 27(12):i–iii, December 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIa

[Ano16a] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(1):1–3, January 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIb

[Ano16b] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(2):79–81, February 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIc

[Ano16c] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(3):147–149, March 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IId

[Ano16d] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(4):229–231, April 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIe

[Ano16e] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(5):303–305, May 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIf

[Ano16f] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(6):391–393, June 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIg

[Ano16g] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(7):507–509, July 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIh

[Ano16h] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(8):619–621, August 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIi

[Ano16i] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(9):713–715, September 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIj

[Ano16j] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(10):835–837, October 2016. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIk

[Ano16k] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(11):943–945, November 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2016:IIl

[Ano16l] Anonymous. Issue information. Journal of Software: Evolution and Process, 28(12):1021–1023, December 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIa

[Ano17a] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(1):??, January 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIb

[Ano17b] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(2):??, February 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIc

[Ano17c] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(3):??, March 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IId

[Ano17d] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(4):??, April 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIe

[Ano17e] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(5):??, May 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIf

[Ano17f] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(6):??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIg

[Ano17g] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(7):??, July 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIh

[Ano17h] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(8):??, August 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIi

[Ano17i] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(9):??, September 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIj

[Ano17j] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(10):??, October 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIk

[Ano17k] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(11):??, November 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:IIl

[Ano17l] Anonymous. Issue information. Journal of Software: Evolution and Process, 29(12):??, December 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2017:RA

[Ano17m] Anonymous. Reviewer acknowledgements. Journal of Software: Evolution and Process, 29(1):??, January 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(1):??, January 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(2):e1900:1– e1900:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(3):e1901:1– e1901:??, March 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(4):e1902:1– e1902:??, April 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(5):e1903:1– e1903:??, May 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(6):e1904:1– e1904:??, June 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIg

[Ano18g] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(7):e1905:1– e1905:??, July 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIh

[Ano18h] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(8):e1906:1– e1906:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIi

[Ano18i] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(9):e1907:1– e1907:??, September 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIj

[Ano18j] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(10):e1908:1– e1908:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIk

[Ano18k] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(11):e1909:1– e1909:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2018:IIl

[Ano18l] Anonymous. Issue information. Journal of Software: Evolution and Process, 30(12):e1910:1– e1910:??, December 2018.

CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2019:IIa

[Ano19a] Anonymous. Issue information. Journal of Software: Evolution and Process, 31(1):e1966:1– e1966:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2019:IIb

[Ano19b] Anonymous. Issue information. Journal of Software: Evolution and Process, 31(2):e1967:1– e1967:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2019:IIc

[Ano19c] Anonymous. Issue information. Journal of Software: Evolution and Process, 31(3):e1968:1– e1968:??, March 2019. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(1):e2185:1– e2185:??, January 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(2):e2186:1– e2186:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(3):e2187:1– e2187:??, March 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(4):e2188:1– e2188:??, April 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(5):e2189:1– e2189:??, May 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(6):e2190:1– e2190:??, June 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIg

[Ano20g] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(7):e2191:1– e2191:??, July 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIh

[Ano20h] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(8):e2192:1– e2192:??, August 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIi

[Ano20i] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(9):e2193:1– e2193:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIj

[Ano20j] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(10):e2194:1– e2194:??, October 2020.

CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIk

[Ano20k] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(11):e2195:1– e2195:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2020:IIl

[Ano20l] Anonymous. Issue information. Journal of Software: Evolution and Process, 32(12):e2196:1– e2196:??, December 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:EPS

[Ano21a] Anonymous. Effort prediction for the software project construction phase. Journal of Software: Evolution and Process, 33(7):e2365:1– e2365:??, July 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:EAS

[Ano21b]  Anonymous.  Enhanc-

ing agile software development in the banking sector — a comprehensive case study at LHV.

Journal of Software: Evolution and Process, 33 (7):e2363:1–e2363:??, July 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:ESA

[Ano21c]  Anonymous.  An ex-

ploratory semantic analysis of logging questions. Journal of Software: Evolution and Process, 33 (7):e2361:1–e2361:??, July 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIa

[Ano21d] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(1):e2278:1– e2278:??, January 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIb

[Ano21e] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(2):e2279:1– e2279:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIc

[Ano21f] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(3):e2280:1– e2280:??, March 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IId

[Ano21g] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(4):e2281:1– e2281:??, April 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIe

[Ano21h] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(5):e2282:1– e2282:??, May 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIf

[Ano21i] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(6):e2283:1– e2283:??, June 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIg

[Ano21j] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(7):e2284:1– e2284:??, July 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIh

[Ano21k] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(8):e2285:1– e2285:??, August 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIi

[Ano21l] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(9):e2286:1– e2286:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIj

[Ano21m] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(10):e2287:1– e2287:??, October 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIk

[Ano21n] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(11):e2288:1– e2288:??, November 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:IIl

[Ano21o] Anonymous. Issue information. Journal of Software: Evolution and Process, 33(12):e2289:1– e2289:??, December 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:MDA

[Ano21p] Anonymous. A modeldriven approach for semantic web service modeling using web service modeling languages. Journal of Software: Evolution and Process, 33 (7):e2364:1–e2364:??, July 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:MUS

[Ano21q]  Anonymous.  Multi-

upgradation software reliability growth model with dependency of faults under change point and imperfect debugging. Journal of Software: Evolution and Process, 33 (6):e2344:1–e2344:??, June 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:VFF

[Ano21r]  Anonymous.  On the

value of filter feature selection techniques in homogeneous ensembles effort estimation. Journal of Software: Evolution and Process, 33 (6):e2343:1–e2343:??, June 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:PLP

[Ano21s] Anonymous. Predicting the lifetime of pull requests in open-source projects. Journal of Software: Evolution and Process, 33 (6):e2337:1–e2337:??, June 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:QSP

[Ano21t] Anonymous. Quantitative software project management with mixed data: a comparison of radial, nonradial, and ensemble data envelopment analysis models. Journal of Software: Evolution and Process, 33 (6):e2348:1–e2348:??, June 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:SDP

[Ano21u] Anonymous. Software defect prediction with imbalanced distribution by radius-synthetic minority over-sampling technique. Journal of Software: Evolution and Process, 33 (7):e2362:1–e2362:??, July 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2021:TEB

[Ano21v]  Anonymous.  Towards

evidence-based decisionmaking for identification and usage of assets in composite software: a research roadmap. Journal of Software: Evolution and Process, 33 (6):e2345:1–e2345:??, June 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIa

[Ano22a] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(1):e2349:1– e2349:??, January 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIb

[Ano22b] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(2):e2350:1– e2350:??, February 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIc

[Ano22c] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(3):e2351:1– e2351:??, March 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IId

[Ano22d] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(4):e2352:1– e2352:??, April 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIe

[Ano22e] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(5):e2353:1– e2353:??, May 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIf

[Ano22f] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(6):e2354:1– e2354:??, June 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIg

[Ano22g] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(7):e2355:1– e2355:??, July 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIh

[Ano22h] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(8):e2356:1– e2356:??, August 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIi

[Ano22i] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(9):e2357:1– e2357:??, September 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIj

[Ano22j] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(10):e2358:1– e2358:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIk

[Ano22k] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(11):e2359:1– e2359:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2022:IIl

[Ano22l] Anonymous. Issue information. Journal of Software: Evolution and Process, 34(12):e2360:1– e2360:??, December 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2023:IIa

[Ano23a] Anonymous. Issue information. Journal of Software: Evolution and Process, 35(1):e2468:1– e2468:??, January 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2023:IIb

[Ano23b] Anonymous. Issue information. Journal of Software: Evolution and Process, 35(2):e2469:1– e2469:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Anonymous:2023:IIc

[Ano23c] Anonymous. Issue information. Journal of Software: Evolution and Process, 35(3):e2470:1– e2470:??, March 2023. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akman:2016:EER

[AÖAG16] Suha Akman, Mert Özmut, Burak Aydın, and Serhat Göktürk. EUROSPI 2015: Experience report: implementing requirement traceability throughout the software development life cycle. Journal of Software: Evolution and Process, 28(11):950–954, November 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Aryani:2014:PDU

[APL^{+}14] Amir Aryani, Fabrizio Perin, Mircea Lungu, Abdun Naser Mahmood, and Oscar Nierstrasz. Predicting dependences using domain-based coupling. Journal of Software: Evolution and Process, 26(1): 50–76, January 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Adam:2012:BPS

[ARD^{+}12] Sebastian Adam, Norman Riegel, Joerg Doerr, Oezguer Uenalan, and Daniel Kerkow. From business processes to software services and vice versa — an improved transition through service-oriented requirements engineering. Journal of Software: Evolution and Process, 24(3): 237–258, April 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Litoriya:2021:SPS

[aRL21] Prateek Pandey and. Ratnesh Litoriya. Software process selection system based on multicriteria decision making. Journal of Software: Evolution and Process, 33(2): e2305:1–e2305:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Asaduzzaman:2016:SEC

[ARSH16] Muhammad Asaduzzaman, Chanchal K. Roy, Kevin A. Schneider, and Daqing Hou. A simple, efficient, context-sensitive approach for code completion. Journal of Software: Evolution and Process, 28(7):512–541, July 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Abrar:2020:MAA

[ASA^{+}20] Muhammad Faisal Abrar, M. Sohail, Sikandar Ali, Muhammad Faran Majeed, Ibrar Ali Shah, Nasir Rashid, and Naimat Ullah. De-motivators for the adoption of agile methodologies for large-scale software development teams: An SLR from management perspective. Journal of Software: Evolution and Process, 32(12): e2268:1–e2268:??, December 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Al-Sarayreh:2013:SBM

[ASACG13] Khalid T. Al-Sarayreh, Alain Abran, and Juan J. Cuadrado-Gallego. A standards-based model of system maintainability requirements. Journal of Software: Evolution and Process, 25(5):459–505, May 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Al-Sabbagh:2018:CBG

[ASG18] Khaled Walid Al-Sabbagh and Lucas Gren. The connections between group maturity, software development velocity, and planning effectiveness. Journal of Software: Evolution and Process, 30(1):??, January 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akbar:2021:FAH

[ASK^{+}21] Muhammad Azeem Akbar, Mohammad Shameem, Arif Ali Khan, Mohammad Nadeem, Ahmed Alsanad, and Abdu Gumaei. A fuzzy analytical hierarchy process to prioritize the success factors of requirement change management in global software development. Journal of Software: Evolution and Process, 33(2): e2292:1–e2292:??, February 2021. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Aggarwal:2017:DDB

[ATR^{+}17] Karan Aggarwal, Finbarr Timbers, Tanner Rutgers, Abram Hindle, Eleni Stroulia, and Russell Greiner. Detecting duplicate bug reports with software engineering domain knowledge. Journal of Software: Evolution and Process, 29(3):??, March 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ajienka:2020:EAS

[AVC20] Nemitari Ajienka, Peter Vangorp, and Andrea Capiluppi. An empirical analysis of source code metrics and smart contract resource consumption. Journal of Software: Evolution and Process, 32 (10):e2267:1–e2267:??, October 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Akarsu:2020:MSA

[AY20] Zulal Akarsu and Murat Yilmaz. Managing the social aspects of software development ecosystems: an industrial case study on personality. Journal of Software: Evolution and Process, 32(11): e2277:1–e2277:??, November 2020. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Asim:2018:SPM

[AYB^{+}18] Muhammad Asim, Artsiom Yautsiukhin, Achim D. Brucker, Thar Baker, Qi Shi, and Brett Lempereur. Security policy monitoring of BPMNbased service compositions. Journal of Software: Evolution and Process, 30(9):e1944:1– e1944:??, September 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Benyagoub:2020:FDS

[BAAO^{+}20] Sarah Benyagoub, Yamine Aït-Ameur, Meriem Ouederni, Atif Mashkoor, and Ahmed Medeghri. Formal design of scalable conversation protocols using Event-B: Validation, experiments, and benchmarks. Journal of Software: Evolution and Process, 32(2):e2209:1– e2209:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Babar:2012:FGS

[Bab12] Muhammad Ali Babar. A framework for groupwaresupported software architecture evaluation process in global software development. Journal of Software: Evolution and Process, 24(2):207–229, March 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Balogun:2022:TSS

[BAB^{+}22] Abdullateef Oluwagbemiga Balogun, Malek Almomani, Shuib Basri, Omar Almomani, Luiz Fernando Capretz, Arif Ali Khan, Abdul Rehman Gilal, and Yahia Baashar. Towards the sustainability of small and medium software enterprises through the implementation of software process improvement: Empirical investigation. Journal of Software: Evolution and Process, 34 (8):e2466:1–e2466:??, August 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Boden:2012:KSP

[BABW12] Alexander Boden, Gabriela Avram, Liam Bannon, and Volker Wulf. Knowledge sharing practices and the impact of cultural factors: reflections on two case studies of offshoring in SME. Journal of Software: Evolution and Process, 24(2):139–152, March 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Braione:2022:ASI

[BBD^{+}22] Pietro Braione, Daniela Briola, Guglielmo De Angelis, Francesco Gallo, Francesco Poggi, and Giovanni Quattrocchi. About the special issue on: “Distributed Complex Systems: Governance, Engineering, and Maintenance”. Journal of Software: Evolution and Process, 34(10):e2459:1– e2459:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Barabino:2015:WFP

[BCC^{+}15] Giulio Barabino, Giulio Concas, Erika Corona, Daniele Grechi, Michele Marchesi, and Danilo Tigano. Web framework points: an effort estimation methodology for Web application development using a content management framework. Journal of Software: Evolution and Process, 27(9):603– 624, September 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bernardi:2013:WAD

[BCD13] Mario Luca Bernardi, Marta Cimitile, and Damiano Distante. Web applications design recovery and evolution with RE– UWA. Journal of Software: Evolution and Process, 25(8):789–814, August 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bernardi:2014:DPD

[BCD14] Mario Luca Bernardi, Marta Cimitile, and Giuseppe Di Lucca. Design pattern detection using a DSL-driven graph matching approach. Journal of Software: Evolution and Process, 26(12):1233–1266, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bernardi:2016:MSD

[BCD16] Mario Luca Bernardi, Marta Cimitile, and Giuseppe Di Lucca. Mining static and dynamic crosscutting concerns: a role-based approach. Journal of Software: Evolution and Process, 28(5):306–339, May 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2019:ASS

[BCPM19] Miklós Biró, Ricardo Colomo-Palacios, and Richard Messnarz. Advances in system, software and service process improvement and innovation. Journal of Software: Evolution and Process, 31(1): e2146:1–e2146:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2020:AER

[BCPM20a] Miklós Biró, Ricardo Colomo-Palacios, and Richard Messnarz. Addressing evolving requirements faced by the software industry. Journal of Software: Evolution and Process, 32(3):e2237:1– e2237:??, March 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2020:CPS

[BCPM20b] Miklós Biró, Ricardo Colomo-Palacios, and Richard Messnarz. Current perspectives on the software engineering process. Journal of Software: Evolution and Process, 32(11): e2313:1–e2313:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Britto:2018:OSD

[BCSS18] Ricardo Britto, Daniela S. Cruzes, Darja Smite, and Aivars Sablis. Onboarding software developers and teams in three globally distributed legacy projects: a multi-case study. Journal of Software: Evolution and Process, 30(4): e1921:1–e1921:??, April 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Brings:2018:ASF

[BDB^{+}18] Jennifer Brings, Marian Daun, Sarah Brinckmann, Kevin Keller, and Thorsten Weyer. Approaches, success factors, and barriers for technology transfer in software engineering — results of a systematic literature review. Journal of Software: Evolution and Process, 30(11): e1981:1–e1981:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bertolino:2023:DCR

[BDG^{+}23] Antonia Bertolino, Guglielmo De Angelis, Antonio Guerriero, Breno Miranda, Roberto Pietrantuono, and Stefano Russo. DevOpRET: Continuous reliability testing in DevOps. Journal of Software: Evolution and Process, 35(3): e2298:1–e2298:??, March 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bagheri:2012:FIS

[BDGR12] Ebrahim Bagheri, Tommaso Di Noia, Dragan Gasevic, and Azzurra Ragone. Formalizing interactive staged feature model configuration. Journal of Software: Evolution and Process, 24(4): 375–400, June 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bucaioni:2022:MBG

[BDS^{+}22] Alessio Bucaioni, Fabio Di Silvestro, Inderjeet Singh, Mehrdad Saadatmand, and Henry Muccini. Model-based generation of test scripts across product variants: an experience report from the railway industry. Journal of Software: Evolution and Process, 34(11): e2498:1–e2498:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

BenMesmia:2021:DWV

[BEB21] Walid Ben Mesmia, Mohamed Escheikh, and Kamel Barkaoui. DevOps workflow verification and duration prediction using non-Markovian stochastic Petri nets. Journal of Software: Evolution and Process, 33(3): e2329:1–e2329:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bonfanti:2018:SLR

[BGM18] Silvia Bonfanti, Angelo Gargantini, and Atif Mashkoor. A systematic literature review of the use of formal methods in medical software systems. Journal of Software: Evolution and Process, 30 (5):e1943:1–e1943:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bonfanti:2020:DVC

[BGM20] Silvia Bonfanti, Angelo Gargantini, and Atif Mashkoor. Design and validation of a C++ code generator from Abstract State Machines specifications. Journal of Software: Evolution and Process, 32(2): e2205:1–e2205:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bersani:2018:OVC

[BGV18] Marcello M. Bersani and Marisol García-Valls. Online verification in cyberphysical systems: Practical bounds for meaningful temporal costs. Journal of Software: Evolution and Process, 30(3), March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bardsiri:2016:DEB

[BH16] Amid Khatibi Bardsiri and Seyyed Mohsen Hashemi. A differential evolutionbased model to estimate the software services development effort. Journal of Software: Evolution and Process, 28(1):57–77, January 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Binkley:2016:SCA

[BHLO16] David Binkley, Daniel Heinz, Dawn Lawrie, and Justin Overfelt. Source code analysis with LDA. Journal of Software: Evolution and Process, 28(10): 893–920, October 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bai:2012:HMS

[BHZK12] Xu Bai, LiGuo Huang, He Zhang, and Supannika Koolmanojwong. Hybrid modeling and simulation for trustworthy software process management: a stakeholderoriented approach. Journal of Software: Evolution and Process, 24(7):721– 740, November 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Barros-Justo:2021:RRM

[BJBM21] José L. Barros-Justo, Fabiane B. V. Benitti, and Jefferson S. Molleri. Risks and risk mitigation in global software development: an update. Journal of Software: Evolution and Process, 33(11): e2370:1–e2370:??, November 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bendraou:2012:APM

[BJF12] Reda Bendraou, Jean-Marc Jezéquél, and Franck Fleurey. Achieving process modeling and execution through the combination of aspect and model-driven engineering approaches. Journal of Software: Evolution and Process, 24(7): 765–781, November 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Baum:2017:CPC

[BKS^{+}17] Tobias Baum, Fabian Kortum, Kurt Schneider, Arthur Brack, and Jens Schauder. Comparing precommit reviews and postcommit reviews using process simulation. Journal of Software: Evolution and Process, 29(11):??, November 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Barbour:2013:ESF

[BKZ13] Liliane Barbour, Foutse Khomh, and Ying Zou. An empirical study of faults in late propagation clone genealogies. Journal of Software: Evolution and Process, 25(11):1139–1165, November 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Binkley:2015:IVN

[BL15] Dave Binkley and Dawn Lawrie. The impact of vocabulary normalization. Journal of Software: Evolution and Process, 27(4): 255–273, April 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bian:2018:CHF

[BLGZ18] Yi Bian, Zheng Li, Junxia Guo, and Ruilian Zhao. Concrete hyperheuristic framework for test case prioritization. Journal of Software: Evolution and Process, 30(11):e1992:1– e1992:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bachmann:2012:ISD

[BM12a] Volver Bachmann and Richard Messnarz. Improving the software development for multiple projects by applying a platform strategy for mechatronic systems. Journal of Software: Evolution and Process, 24(5):541– 549, August 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2012:ETS

[BM12b] Miklós Biró and Richard Messnarz. Editorials: Topics in software, systems and services process improvement. Journal of Software: Evolution and Process, 24(5):455–458, August 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bachmann:2013:ISA

[BM13] Volker Bachmann and Richard Messnarz. Improving safety and availability of complex systems by using an integrated design approach in development. Journal of Software: Evolution and Process, 25(4):341–349, April 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bavota:2018:ISI

[BMC18] Gabriele Bavota, Jonathan I.

Maletic, and Michael L. Collard. Introduction to the special issue on program comprehension. Journal of Software: Evolution and Process, 30 (6):e1962:1–e1962:??, June 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2014:EAT

[BMCP14a] Miklós Biró, Richard Messnarz, and Ricardo Colomo-Palacios. Editorial: Application themes of software process assessment and improvement. Journal of Software: Evolution and Process, 26(3): 281–283, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2014:ESP

[BMCP14b] Miklós Biró, Richard Messnarz, and Ricardo Colomo-Palacios. Editorial: Software process improvement leveraged in various application domains. Journal of Software: Evolution and Process, 26(5):465–467, May 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2014:ESI

[BMCP14c] Miklós Biró, Richard Messnarz, and Ricardo Colomo-Palacios. Editorials: Snapshot of industrial experiences shared at the 20th Anniversary EuroSPI Conference. Journal of Software: Evolution and Process, 26(9):771– 775, September 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2015:EPI

[BMCP15] Miklós Biró, Richard Messnarz, and Ricardo Colomo-Palacios. Editorial: Process improvement approaches fertilised by advances in SPI. Journal of Software: Evolution and Process, 27(8):509– 513, August 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2016:ERL

[BMCP16] Miklós Biró, Richard Messnarz, and Ricardo Colomo-Palacios. Editorial: Results and lessons learned in system, software & service process improvement & innovation. Journal of Software: Evolution and Process, 28(11):946– 949, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Barafort:2019:IBI

[BMM19] Béatrix Barafort, Antoni-Lluís Mesquida, and Antònia Mas. ISO 31000-based integrated risk management process assessment model for IT organizations. Journal of Software: Evolution and Process, 31(1): e1984:1–e1984:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2013:EIE

[BMP13] Miklós Biró, Richard Messnarz, and Ricardo Colomo Palacios. Editorial: Industrial experiences with software process assessment and improvement: (special issue with selected industrial experience papers of EuroSPI^{2} 2010). Journal of Software: Evolution and Process, 25(4): 325–327, April 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Biro:2021:SSC

[BMS21] Miklós Biró, Atif Mashkoor, and Johannes Sametinger. Safe and secure cyberphysical systems. Journal of Software: Evolution and Process, 33(9): e2340:1–e2340:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bezemer:2014:DAP

[BMZP14] C. Bezemer, E. Milon, A. Zaidman, and J. Pouwelse. Detecting and analyzing I/O performance regressions. Journal of Software: Evolution and Process, 26 (12):1193–1212, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bai:2021:SPI

[BNBA21] Baogang Bai, Shah Nazir, Yuhe Bai, and Amir Anees. Security and provenance for Internet of Health Things: a systematic literature review. Journal of Software: Evolution and Process, 33 (5):e2335:1–e2335:??, May 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bosch:2018:ETW

[BO18] Jan Bosch and Helena Holmström Olsson. Ecosystem traps and where to find them. Journal of Software: Evolution and Process, 30(11):e1961:1– e1961:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bosch:2021:DRM

[BO21] Jan Bosch and Helena H.

Olsson. Digital for real: a multicase study on the digital transformation of companies in the embedded systems domain. Journal of Software: Evolution and Process, 33 (5):e2333:1–e2333:??, May 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bollin:2013:MQE

[Bol13] Andreas Bollin. Metrics for quantifying evolutionary changes in Z specifications. Journal of Software: Evolution and Process, 25 (9):1027–1059, September 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Borg:2016:TFT

[Bor16] Markus Borg. TuneR: a framework for tuning software engineering tools with hands-on instructions in R. Journal of Software: Evolution and Process, 28(6):427–459, June 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Bian:2016:TTR

[BPKZ16] Yixin Bian, Mohammed Aziz Parande, Gunes Koru, and Song Zhao. Testing the theory of relative dependency from an evolutionary perspective: higher dependencies concentration in smaller modules over the lifetime of software products. Journal of Software: Evolution and Process, 28(5):340–371, May 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Banach:2021:FMS

[BRDL21] Richard Banach, Joseph Razavi, Oivier Debicki, and Suzanne Lesecq. Formal methods by stealth: The INSPEX experience. Journal of Software: Evolution and Process, 33(12): e2383:1–e2383:??, December 2021. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Blum:2016:ADS

[BSB16] Fabian Rojas Blum, Jocelyn Simmonds, and María Cecilia Bastarrica. The v-algorithm for discovering software process lines. Journal of Software: Evolution and Process, 28 (9):783–799, September 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Belategi:2014:ESP

[BSEA14] Lorea Belategi, Goiuria Sagardui, Leire Etxeberria, and Maider Azanza. Embedded software product lines: domain and application engineering model-based analysis processes. Journal of Software: Evolution and Process, 26(4):419–433, April 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Babiceanu:2018:EHA

[BW18] Radu Babiceanu and Hélène Waeselynck. Emerging high assurance solutions for safe, secure, and reliable software systems. Journal of Software: Evolution and Process, 30(3): e1950:1–e1950:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Betka:2022:TPA

[BW22] Maik Betka and Stefan Wagner. Towards practical application of mutation testing in industry — Traditional versus extreme mutation testing. Journal of Software: Evolution and Process, 34(11): e2450:1–e2450:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Banach:2021:LEH

[BZ21] Richard Banach and Huibiao Zhu. Language evolution and healthiness for critical cyber-physical systems. Journal of Software: Evolution and Process, 33(9):e2301:1– e2301:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Charalampidou:2021:ESS

[CAKA21] Sofia Charalampidou, Apostolos Ampatzoglou, Evangelos Karountzos, and Paris Avgeriou. Empirical studies on software traceability: a mapping study. Journal of Software: Evolution and Process, 33(2): e2294:1–e2294:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cao:2012:DCT

[Cao12] Lan Cao. Dynamic capability for trustworthy software development. Journal of Software: Evolution and Process, 24(7): 837–850, November 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Capiluppi:2013:SCO

[Cap13] Andrea Capiluppi. Similarities, challenges and opportunities of Wikipedia content and open source projects. Journal of Software: Evolution and Process, 25(9):891–914, September 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Casey:2013:LTA

[CC13] Valentine Casey and Fergal Mc Caffery. A lightweight traceability assessment method for medical device software. Journal of Software: Evolution and Process, 25(4): 363–372, April 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chugh:2021:ACK

[CCUP21] Mitali Chugh, Nitin Chanderwal, Rajesh Kumar Upadhyay, and Devendra Kumar Punia. Antecedents and consequences of knowledge sharing for software process improvement in the Indian software industry. Journal of Software: Evolution and Process, 33(2): e2291:1–e2291:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cetin:2019:BSM

[ÇD19] Esra Çetin and Pinar Onay Durdu. Blended Scrum model for software development organizations. Journal of Software: Evolution and Process, 31(2): e2147:1–e2147:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Capilla:2013:ERV

[CDF13] Rafael Capilla, Juan C. Dueñas, and Rudolf Ferenc. Editorial: A retrospective view of software maintenance and reengineering research — a selection of papers from European Conference on Software Maintenance and Reengineering 2010. Journal of Software: Evolution and Process, 25(6): 569–574, June 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cavalcanti:2014:COS

[CdMSNdCM^{+}14] Yguaratã Cerqueira Cavalcanti, Paulo Anselmo da Mota Silveira Neto, Ivan do Carmo Machado, Tassio Ferreira Vale, Eduardo Santana de Almeida, and Silvio Romero de Lemos Meira. Challenges and opportunities for software change request repositories: a systematic mapping study. Journal of Software: Evolution and Process, 26(7): 620–653, July 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Capobianco:2013:IIB

[CDO^{+}13] Giovanni Capobianco, Andrea De Lucia, Rocco Oliveto, Annibale Panichella, and Sebastiano Panichella. Improving IR-based traceability recovery via nounbased indexing of software artifacts. Journal of Software: Evolution and Process, 25(7):743–762, July 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Canfora:2012:ESE

[CDR12] Gerardo Canfora, Darren Dalcher, and David Raffo. Editorials: Software: evolution and process. A new journal is born. Journal of Software: Evolution and Process, 24(1):1, January 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Canfora:2017:EFY

[CDR17] Gerardo Canfora, Darren Dalcher, and David Raffo. Editorial: A five year retrospective. Journal of Software: Evolution and Process, 29(1):??, January 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Campos:2016:SCK

[CdSdAM16] Eduardo C. Campos, Lucas B. L. de Souza, and Marcelo de A. Maia. Searching crowd knowledge to recommend solutions for API usage tasks. Journal of Software: Evolution and Process, 28(10): 863–892, October 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cholez:2014:MAP

[CG14] Hervé Cholez and Frédéric Girard. Maturity assessment and process improvement for information security management in small and medium enterprises. Journal of Software: Evolution and Process, 26(5): 496–503, May 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chitchyan:2017:USC

[CGN17] Ruzanna Chitchyan, Iris Groher, and Joost Noppen. Uncovering sustainability concerns in software product lines. Journal of Software: Evolution and Process, 29(2):??, February 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

CastellanosArdila:2022:CCS

[CGU22] Julieth Patricia Castellanos Ardila, Barbara Gallina, and Faiz Ul Muram. Compliance checking of software processes: a systematic literature review. Journal of Software: Evolution and Process, 34 (5):e2440:1–e2440:??, May 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Charrada:2015:SRU

[CKG15] Eya Ben Charrada, Anne Koziolek, and Martin Glinz. Supporting requirements update during software evolution. Journal of Software: Evolution and Process, 27(3): 166–194, March 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Calderon:2015:SPI

[CKMN15] Natalja Nikitina Calderon, Mira Kajko-Mattsson, and Andrew James Nolan. Successful process improvement projects are no accidents. Journal of Software: Evolution and Process, 27 (11):896–911, November 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Choi:2015:ICF

[CKP15] Su-Jin Choi, Dae-Kyoo Kim, and Sooyong Park. Identifying correlations of findings for building process improvement packages using graph clustering. Journal of Software: Evolution and Process, 27 (8):514–527, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Carroll:2018:SIS

[CKR18] Noel Carroll, Craig Kuziemsky, and Ita Richardson. Special issue on software engineering for Connected Health: Challenges and research roadmap. Journal of Software: Evolution and Process, 30 (7):e1941:1–e1941:??, July 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chen:2018:CRI

[CKS18] Zhiyuan Chen, Young-Woo Kwon, and Myoungkyu Song. Clone refactoring inspection by summarizing clone refactorings and detecting inconsistent changes during software evolution. Journal of Software: Evolution and Process, 30(10): e1951:1–e1951:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chinenyeze:2017:BBD

[CLAD17] Samuel J. Chinenyeze, Xiaodong Liu, and Ahmed Al-Dubai. BEFTIGRE: Behaviour-driven full-tier green evaluation of mobile cloud applications. Journal of Software: Evolution and Process, 29(2):??, February 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ceccato:2014:EGE

[CLC14] Mariano Ceccato, Zheng Li, and James R. Cordy. Editorial: Guest editorial for the special issue on source code analysis and manipulation, SCAM 2012. Journal of Software: Evolution and Process, 26(6):531–532, June 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Concas:2013:SSM

[CLMZ13] Giulio Concas, Maria Ilaria Lunesu, Michele Marchesi, and Hongyu Zhang. Simulation of software maintenance process, with and without a work-in-process limit. Journal of Software: Evolution and Process, 25 (12):1225–1248, December 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cui:2022:WAI

[CLW22] Can Cui, Bin Liu, and Shihai Wang. WIFLF: an approach independent of the target project for cross-project defect prediction. Journal of Software: Evolution and Process, 34 (12):e2497:1–e2497:??, December 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Calvo-Manzano:2012:MPI

[CMCG^{+}12] Jose A. Calvo-Manzano, Gonzalo Cuevas, Gerzón Gómez, Jezreel Mejia, Mirna Muñoz, and Tomás San Feliu. Methodology for process improvement through basic components and focusing on the resistance to change. Journal of Software: Evolution and Process, 24(5):511– 523, August 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chen:2020:DDC

[CMQ^{+}20] Xiang Chen, Yanzhou Mu, Yubin Qu, Chao Ni, Meng Liu, Tong He, and Shangqing Liu. Do different cross-project defect prediction methods identify the same defective modules? Journal of Software: Evolution and Process, 32 (5):e2234:1–e2234:??, May 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cataldo:2012:IGD

[CN12] Marcelo Cataldo and Sangeeth Nambiar. The impact of geographic distribution and the nature of technical coupling on the quality of global software development projects. Journal of Software: Evolution and Process, 24(2):153– 168, March 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cuesta:2013:ESU

[CNPR13] Carlos E. Cuesta, Elena Navarro, Dewayne E. Perry, and Cristina Roda. Evolution styles: using architectural knowledge as an evolution driver. Journal of Software: Evolution and Process, 25(9):957– 980, September 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Crabtree:2014:ESP

[CNS14] Carlton A. Crabtree, Anthony F. Norcio, and Carolyn B. Seaman. An empirical study of process knowledge: coherence as a static process property. Journal of Software: Evolution and Process, 26(4): 451–463, April 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chakroborti:2023:RCO

[CNSR23]  Debasish  Chakroborti,

Sristy Sumana Nath, Kevin A. Schneider, and Chanchal K. Roy. Release conventions of open-source software: an exploratory study. Journal of Software: Evolution and Process, 35 (1):e2499:1–e2499:??, January 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Clarke:2013:EEE

[CO13] Paul Clarke and Rory V. O’Connor. An empirical examination of the extent of software process improvement in software SMEs. Journal of Software: Evolution and Process, 25(9):981–998, September 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cervantes-Ojeda:2022:CND

[COGFCA22] Jorge Cervantes-Ojeda, María Gómez-Fuentes, and Guillermo Chacón-Acosta. Can non-developers learn a simplified modeling notation quickly? Journal of Software: Evolution and Process, 34(8): e2481:1–e2481:??, August 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Colomo-Palacios:2014:USM

[CPMS^{+}14] Ricardo Colomo-Palacios, Richard Messnarz, Kerstin Siakas, Daniel Palosi, and Darragh Coakley. Using social media as a tool for business improvement and certification of knowledge workers. Journal of Software: Evolution and Process, 26(9): 791–798, September 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Carrozza:2015:DAM

[CPR15] Gabriella Carrozza, Roberto Pietrantuono, and Stefano Russo. Defect analysis in mission-critical software systems: a detailed investigation. Journal of Software: Evolution and Process, 27(1):22–49, January 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Colomo-Palacios:2020:CRB

[CPSGAA20] Ricardo Colomo-Palacios, Mary Sánchez-Gordón, and Daniel Arias-Aranda. A critical review on blockchain assessment initiatives: a technology evolution viewpoint. Journal of Software: Evolution and Process, 32(11): e2272:1–e2272:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cortina:2014:APO

[CRP14] Stéphane Cortina, Alain Renault, and Michel Picard. Assessing partially outsourced processes — lessons learned from TIPA assessments. Journal of Software: Evolution and Process, 26(9):799–807, September 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chatzigeorgiou:2013:CMS

[CS13] Alexander Chatzigeorgiou and Emmanouil Stiakakis. Combining metrics for software evolution assessment by means of Data Envelopment Analysis. Journal of Software: Evolution and Process, 25(3): 303–324, March 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chatterjee:2019:UAT

[CS19] Subhashis Chatterjee and Ankur Shukla. A unified approach of testing coverage-based software reliability growth modelling with fault detection probability, imperfect debugging, and change point. Journal of Software: Evolution and Process, 31(3): e2150:1–e2150:??, March 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cruz:2017:WFD

[CSGR17] Luis C. Cruz, Heider Sanchez, Víctor M. González, and Romain Robbes. Work fragmentation in developer interaction data. Journal of Software: Evolution and Process, 29(3):??, March 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chen:2013:DID

[CsZZ^{+}13] Ting Chen, Xiao song Zhang, Cong Zhu, Xiao li Ji, Shi ze Guo, and Yue Wu. Design and implementation of a dynamic symbolic execution tool for windows executables. Journal of Software: Evolution and Process, 25 (12):1249–1272, December 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Cordy:2013:ESS

[CT13] James R. Cordy and Paolo Tonella. Editorial: Special section on ICSM 2011. Journal of Software: Evolution and Process, 25 (11):1137–1138, November 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Couckuyt:2021:ESG

[CV21] Dries Couckuyt and Amy Van Looy. An empirical study on Green BPM adoption: Contextual factors and performance. Journal of Software: Evolution and Process, 33(3): e2299:1–e2299:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Codabux:2017:EAT

[CWBC17] Zadia Codabux, Byron J. Williams, Gary L. Bradshaw, and Murray Cantor. An empirical assessment of technical debt practices in industry. Journal of Software: Evolution and Process, 29(10):??, October 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chaudhry:2020:SAD

[CYK20] Natalia Chaudhry, Muhammad Murtaza Yousaf, and Muhammad Taimoor Khan. Security assessment of data management systems for cyber physical system applications. Journal of Software: Evolution and Process, 32(2): e2241:1–e2241:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chen:2022:MVN

[CZY^{+}22] Ningning Chen, Huibiao Zhu, Jiaqi Yin, Yuan Fei, Lili Xiao, and Minghua Zhu. Modeling and verifying NDN-based IoV using CSP. Journal of Software: Evolution and Process, 34 (10):e2371:1–e2371:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DaSilva:2022:BCW

[DBP22] Léuson Da Silva, Paulo Borba, and Arthur Pires. Build conflicts in the wild. Journal of Software: Evolution and Process, 34(4): e2441:1–e2441:??, April 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Decker:2020:SSD

[DCVM20] Michael John Decker, Michael L. Collard, L. Gwenn Volkert, and Jonathan I. Maletic. srcDiff: a syntactic differencing approach to improve the understandability of deltas. Journal of Software: Evolution and Process, 32(4): e2226:1–e2226:??, April 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DeMan:2015:ETA

[De 15] Jozef De Man. Experience with teaching and applying process appraisals at a university. Journal of Software: Evolution and Process, 27(8):539–544, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DeMan:2016:EOD

[De 16] Jozef De Man. EUROSPI 2015: Objective-driven process appraisals: an educational experiment. Journal of Software: Evolution and Process, 28(11):955– 959, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DeAngelis:2018:RER

[DFGP18] Guglielmo De Angelis, Alessio Ferrari, Stefania Gnesi, and Andrea Polini. Requirements elicitation and refinement in collaborative research projects. Journal of Software: Evolution and Process, 30(12): e1990:1–e1990:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

deGea:2016:CLD

[dGNA^{+}16] Juan M. Carrillo de Gea, Joaquín Nicolás, José L. Fernández Alemán, Ambrosio Toval, Sofia Ouhbi, and Ali Idri. Colocated and distributed natural-language requirements specification: traditional versus reuse-based techniques. Journal of Software: Evolution and Process, 28(3):205–227, March 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

deGea:2017:ASR

[dGNFAT17] Juan Manuel Carrillo de Gea, Joaquín Nicolás, José L. Fernández-Alemán, and Ambrosio Toval. Automated support for reusebased requirements engineering in global software engineering. Journal of Software: Evolution and Process, 29(8):??, August 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DiSorbo:2021:ICU

[DGVP21] Andrea Di Sorbo, Giovanni Grano, Corrado Aaron Visaggio, and Sebastiano Panichella. Investigating the criticality of userreported issues through their relations with app rating. Journal of Software: Evolution and Process, 33(3):e2316:1– e2316:??, March 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hamouda:2013:BBS

[dH13] Alaa El deen Hamouda. The BSCBAS: a Balanced Scorecard-based Appraisal System for improving the performance of software organizations. Journal of Software: Evolution and Process, 25(7):763– 780, July 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Dreves:2016:EMR

[DHH^{+}16] Rainer Dreves, Frank Hällmayer, Lutz Haunert, Bernhard Sechser, and Armin Rieß. EUROSPI 2015: A method to realize traceability in development processes. Journal of Software: Evolution and Process, 28(11):1011– 1019, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Daoudagh:2023:AFC

[DLM23] Said Daoudagh, Francesca Lonetti, and Eda Marchetti. An automated framework for continuous development and testing of access control systems. Journal of Software: Evolution and Process, 35(3): e2306:1–e2306:??, March 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DeNicola:2012:TMB

[DMS12] Antonio De Nicola, Michele Missikoff, and Fabrizio Smith. Towards a method for business process and informal business rules compliance. Journal of Software: Evolution and Process, 24(3):341–360, April 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DiFrancescomarino:2013:CBM

[DMT13] Chiara Di Francescomarino, Alessandro Marchetto, and Paolo Tonella. Clusterbased modularization of processes recovered from web applications. Journal of Software: Evolution and Process, 25(2):113– 138, February 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Dou:2022:JJA

[DMZ^{+}22] Hui Dou, Jicheng Mei, Yiwen Zhang, Pengfei Chen, and Zibin Zheng. Joint-Conf: Jointly autotuning configuration parameters for modularized graph databases. Journal of Software: Evolution and Process, 34(12):e2495:1– e2495:??, December 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Dowling:2014:STR

[Dow14] Phelim Dowling. Successfully transitioning a research project to a commercial spin-out using an agile software process. Journal of Software: Evolution and Process, 26(5): 468–475, May 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

DiRuscio:2015:MDA

[DP15] Davide Di Ruscio and Patrizio Pelliccione. A modeldriven approach to detect faults in FOSS systems. Journal of Software: Evolution and Process, 27(4): 294–318, April 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Dit:2013:FLS

[DRGP13] Bogdan Dit, Meghan Revelle, Malcom Gethers, and Denys Poshyvanyk. Feature location in source code: a taxonomy and survey. Journal of Software: Evolution and Process, 25(1):53–95, January 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Datta:2021:URB

[DRM21] Subhajit Datta, Reshma Roychoudhuri, and Subhashis Majumder. Understanding the relation between repeat developer interactions and bug resolution times in large open source ecosystems: a multisystem study. Journal of Software: Evolution and Process, 33(4): e2317:1–e2317:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Davies:2014:CTB

[DRW14] Steven Davies, Marc Roper, and Murray Wood. Comparing text-based and dependence-based approaches for determining the origins of bugs. Journal of Software: Evolution and Process, 26(1):107–139, January 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Diebold:2017:SPM

[DS17] P. Diebold and S. A.

Scherr. Software process models vs descriptions: What do practitioners use and need? Journal of Software: Evolution and Process, 29(11): ??, November 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lopes:2015:SDM

[dSLBF15] Jhoney da. Silva Lopes, José Luis Braga, and Moisés A. Resende Filho. Systems dynamics model for decision support in risk assessment in software projects. Journal of Software: Evolution and Process, 27(12):976–989, December 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

daSilva:2012:EBM

[dSPF^{+}12] Fabio Q. B. da Silva, Rafael Prikladnicki, A. César C. França, Cleviton V. F. Monteiro, Catarina Costa, and Rodrigo Rocha. An evidence-based model of distributed software development project management: results from a systematic mapping study. Journal of Software: Evolution and Process, 24(6): 625–642, October 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

dosSantos:2022:SMS

[dSVB^{+}22] Rodrigo Augusto dos Santos, Darli Vieira, Alencar Bravo, Larissa Suzuki, and Fadiah Qudah. A systematic mapping study on the employment of neural networks on software engineering projects: Where to go next? Journal of Software: Evolution and Process, 34(3): e2402:1–e2402:??, March 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Diebold:2018:HAD

[DT18] Philipp Diebold and Sven Theobald. How is agile development currently being used in regulated embedded domains? Journal of Software: Evolution and Process, 30(8): e1935:1–e1935:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Erich:2017:QSD

[EAD17] F. M. A. Erich, C. Amrit, and M. Daneva. A qualitative study of DevOps usage in practice. Journal of Software: Evolution and Process, 29(6): ??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Elish:2013:SMQ

[EAK13] Mahmoud O. Elish and Mojeeb Al-Rahman Al-Khiaty. A suite of metrics for quantifying historical changes to predict future change-prone classes in object-oriented software. Journal of Software: Evolution and Process, 25(5):407–437, May 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Espinosa-Curiel:2013:FEC

[ECRJFZ13] Ismael Edrein Espinosa-Curiel, Josefina Rodríguez-Jacobo, and José Alberto Fernández-Zepeda. A framework for evaluation and control of the factors that influence the software process improvement in small organizations. Journal of Software: Evolution and Process, 25(4): 393–406, April 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Espinosa-Curiel:2016:USS

[ECRJFZ16] Ismael Edrein Espinosa-Curiel, Josefina Rodríguez-Jacobo, and José Alberto Fernández-Zepeda. Understanding SPI in small organizations: a study of Mexican software enterprises. Journal of Software: Evolution and Process, 28(5): 372–390, May 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Espinosa-Curiel:2018:ACC

[ECRJVA^{+}18] Ismael Edrein Espinosa-Curiel, Josefina Rodríguez-Jacobo, Erika Vázquez-Alfaro, José Alberto Fernández-Zepeda, and Daniel Fajardo-Delgado. Analysis of the changes in communication and social interactions during the transformation of a traditional team into an agile team. Journal of Software: Evolution and Process, 30(9): e1946:1–e1946:??, September 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

ElKoutbi:2019:EEE

[EIA19] Salma El Koutbi, Ali Idri, and Alain Abran. Empirical evaluation of an entropy-based approach to estimation variation of software development effort. Journal of Software: Evolution and Process, 31(3):e2149:1– e2149:??, March 2019. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Eddy:2018:ISW

[EKG18] Brian P. Eddy, Nicholas A.

Kraft, and Jeff Gray. Impact of structural weighting on a latent Dirichlet allocation-based feature location technique. Journal of Software: Evolution and Process, 30(1): ??, January 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Eze:2018:CBP

[EKP18] Benjamin Eze, Craig Kuziemsky, and Liam Peyton. Cloud-based performance management of community care services. Journal of Software: Evolution and Process, 30 (7):e1897:1–e1897:??, July 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Eveleens:2015:QFP

[EKV15] Laurenz Eveleens, Peter Kampstra, and Chris Verhoef. Quantifying fair payment after outsourcing — a case study. Journal of Software: Evolution and Process, 27(3): 147–165, March 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

ElBeggar:2018:MDA

[El 18] Omar El Beggar. Multicriteria decision aid for agile methods evaluation using fuzzy PROMETHEE. Journal of Software: Evolution and Process, 30(12): e2108:1–e2108:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Laila:2023:MCO

[eLAA^{+}23] Umm e Laila, Najeed Ahmed, Asad Arfeen, Agha Yasir Ali, Mohammad Khurram, and Muzammil Ahmed Khan. Mission-critical open-source software adoption model validation using Partial Least Square — Structural Equation Modeling. Journal of Software: Evolution and Process, 35(2): e2514:1–e2514:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Elberzhager:2014:IIT

[EMRF14] Frank Elberzhager, Jürgen Münch, Dieter Rombach, and Bernd Freimut. Integrating inspection and test processes based on context-specific assumptions. Journal of Software: Evolution and Process, 26(4):371–385, April 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Erdogan:2018:MES

[EPG18] Onur Erdogan, Muhammed Emre Pekkaya, and Halime Gök. More effective sprint retrospective with statistical analysis. Journal of Software: Evolution and Process, 30(5):e1933:1– e1933:??, May 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Eken:2021:DCL

[ETT^{+}21] Beyza Eken, Selda Tufan, Alper Tunaboylu, Tevfik Guler, Rifat Atar, and Ayse Tosun. Deployment of a change-level software defect prediction solution into an industrial setting. Journal of Software: Evolution and Process, 33(11): e2381:1–e2381:??, November 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Feitosa:2017:IED

[FAA^{+}17] Daniel Feitosa, Rutger Alders, Apostolos Ampatzoglou, Paris Avgeriou, and Elisa Yumi Nakagawa. Investigating the effect of design patterns on energy consumption. Journal of Software: Evolution and Process, 29(2):??, February 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Filippone:2022:SCA

[FAT22] Gianluca Filippone, Marco Autili, and Massimo Tivoli. Synthesis of context-aware business-to-business processes for location-based services through choreographies. Journal of Software: Evolution and Process, 34(10):e2416:1– e2416:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Fabijan:2018:EGE

[FDM^{+}18] Aleksander Fabijan, Pavel Dmitriev, Colin McFarland, Lukas Vermeer, Helena Holmström Olsson, and Jan Bosch. Experimentation growth: Evolving trustworthy A/B testing capabilities in online software companies. Journal of Software: Evolution and Process, 30(12): e2113:1–e2113:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Frey:2013:ACC

[FHS13] Sören Frey, Wilhelm Hasselbring, and Benjamin Schnoor. Automatic conformance checking for migrating software systems to cloud infrastructures and platforms. Journal of Software: Evolution and Process, 25(10):1089–1115, October 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ferenc:2013:EIS

[FKW13] Rudolf Ferenc, Jens Knodel, and Andreas Winter. Editorial: Introduction to the special issue of the 13th European Conference on Software Maintenance and Reengineering (CSMR 2009). Journal of Software: Evolution and Process, 25(2): 111–112, February 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Forti:2022:OMD

[FLGB22] Stefano Forti, Isaac Lera, Carlos Guerrero, and Antonio Brogi. Osmotic management of distributed complex systems: a declarative decentralised approach. Journal of Software: Evolution and Process, 34(10):e2405:1– e2405:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Feng:2023:FGA

[FM23] Qiong Feng and Ran Mo. Fine-grained analysis of dependency cycles among classes. Journal of Software: Evolution and Process, 35(1): e2496:1–e2496:??, January 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Flood:2015:RII

[FMC^{+}15] Derek Flood, Fergal Mc-Caffery, Valentine Casey, Ruth McKeever, and Peter Rust. A roadmap to ISO 14971 implementation. Journal of Software: Evolution and Process, 27(5):319–336, May 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Fontana:2013:DPS

[FMR13] Francesca Arcelli Fontana, Stefano Maggioni, and Claudia Raibulet. Design patterns: a survey on their micro-structures. Journal of Software: Evolution and Process, 25(1):27–52, January 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Filho:2017:ATA

[FPA17] Marum Simão Filho, Plácido Rogério Pinheiro, and Adriano Bessa Albuquerque. Analysis of task allocation in distributed software development through a hybrid methodology of verbal decision analysis. Journal of Software: Evolution and Process, 29(7):??, July 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Farooq:2017:RES

[FQA17] Sheikh Umar Farooq, S. M. K. Quadri, and Nesar Ahmad. A replicated empirical study to evaluate software testing methods. Journal of Software: Evolution and Process, 29(9): ??, September 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Fleck:2016:SBM

[FTW16] Martin Fleck, Javier Troya, and Manuel Wimmer.

Search-based model transformations. Journal of Software: Evolution and Process, 28(12):1081–1117, December 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Fei:2018:CMV

[FZW^{+}18] Yuan Fei, Huibiao Zhu, Xi Wu, Huixing Fang, and Shengchao Qin. Comparative modelling and verification of Pthreads and Dthreads. Journal of Software: Evolution and Process, 30(3): e1919:1–e1919:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ghaemi:2020:SBH

[GA20] Amir Ghaemi and Bahman Arasteh. SFLA-based heuristic method to generate software structural test data. Journal of Software: Evolution and Process, 32 (1):e2228:1–e2228:??, January 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ghaleb:2018:PCT

[GAA18] Taher Ahmed Ghaleb, Musab A. Alturki, and Khalid Aljasser. Program comprehension through reverse-engineered sequence diagrams: a systematic review. Journal of Software:

Evolution and Process, 30(11):e1965:1–e1965:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gartziandia:2022:MLB

[GAA^{+}22] Aitor Gartziandia, Aitor Arrieta, Jon Ayerdi, Miren Illarramendi, Aitor Agirre, Goiuria Sagardui, and Maite Arratibel. Machine learning-based test oracles for performance testing of cyber-physical systems: an industrial case study on elevators dispatching algorithms. Journal of Software: Evolution and Process, 34(11): e2465:1–e2465:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gallina:2020:QET

[Gal20] Barbara Gallina. Quantitative evaluation of tailoring within SPICEcompliant security-informed safety-oriented process lines. Journal of Software: Evolution and Process, 32(3):e2212:1– e2212:??, March 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gonzalez-Barahona:2014:SLS

[GBRHO14] Jesus M. Gonzalez-Barahona, Gregorio Robles, Israel Herraiz, and Felipe Ortega. Studying the laws of software evolution in a long-lived FLOSS project. Journal of Software: Evolution and Process, 26(7): 589–612, July 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Grossi:2014:HML

[GCMS14] Lucas Grossi, Jose A. Calvo-Manzano, and Tomas San Feliu. High-maturity levels: achieving CMMI ML-5 in a consultancy company. Journal of Software: Evolution and Process, 26(9):808–817, September 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gonzalez:2019:AGB

[GCPM19] Anaisa Hernández González, Coral Calero, Dianelys Pérez Parra, and Javier Mancebo. Approaching Green BPM characterisation. Journal of Software: Evolution and Process, 31(2): e2145:1–e2145:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Guncan:2021:UCB

[GD21] Damla Güncan and Pinar Onay Durdu. A user-centered behavioral software development model. Journal of Software: Evolution and Process, 33(2): e2274:1–e2274:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Guerrouj:2013:TIS

[GDAG13] Latifa Guerrouj, Massimiliano Di Penta, Giuliano Antoniol, and Yann-Gaël Guéhéneuc. TIDIER: an identifier splitting approach using speech recognition techniques. Journal of Software: Evolution and Process, 25(6): 575–599, June 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ge:2018:CCS

[GDJV18] Ning Ge, Arnaud Dieumegard, Eric Jenn, and Laurent Voisin. Correct-by-construction specification to verified code. Journal of Software: Evolution and Process, 30(10): e1959:1–e1959:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Goncalves:2018:HPE

[GdOK18] Taisa Guidini Gonçalves, Káthia Marçal de Oliveira, and Christophe Kolski. HCI in practice: an empirical study with software process capability maturity model consultants in Brazil. Journal of Software: Evolution and Process, 30(11): e2109:1–e2109:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gharib:2021:MDA

[GdSC21] Mohamad Gharib, Leandro Dias da Silva, and Andrea Ceccarelli. A model to discipline autonomy in cyber-physical systems-of-systems and its application. Journal of Software: Evolution and Process, 33(9):e2328:1– e2328:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ghannem:2014:MRU

[GEK14] Adnane Ghannem, Ghizlane El Boussaidi, and Marouane Kessentini. Model refactoring using examples: a search-based approach. Journal of Software: Evolution and Process, 26(7):692–713, July 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garousi:2020:EIC

[GFK^{+}20] Vahid Garousi, Michael Felderer, Marco Kuhrmann, Kadir Herkiloglu, and Sigrid Eldh. Exploring the industry’s challenges in software testing: an empirical study. Journal of Software: Evolution and Process, 32(8): e2251:1–e2251:??, August 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Goyal:2018:IUC

[GFKH18] Raman Goyal, Gabriel Ferreira, Christian Kästner, and James Herbsleb. Identifying unusual commits on GitHub. Journal of Software: Evolution and Process, 30(1):??, January 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Godfrey:2014:ELL

[GG14] Michael W. Godfrey and Daniel M. German. On the evolution of Lehman’s Laws. Journal of Software: Evolution and Process, 26(7):613–619, July 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garcia-Garcia:2018:MBS

[GGGBEM18] Julián Alberto García-García, Laura García-Borgoñón, María José Escalona, and Manuel Mejías. A model-based solution for process modeling in practice environments: PLM_{4} BS. Journal of Software: Evolution and Process, 30(12): e1982:1–e1982:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gren:2020:AWW

[GGJ20] Lucas Gren, Alfredo Goldman, and Christian Jacobsson. Agile ways of working: a team maturity perspective. Journal of Software: Evolution and Process, 32 (6):e2244:1–e2244:??, June 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gorschek:2012:IPM

[GGPT12] Tony Gorschek, Andrigo Gomes, Andreas Pettersson, and Richard Torkar. Introduction of a process maturity model for market-driven product management and requirements engineering. Journal of Software: Evolution and Process, 24(1):83– 113, January 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ghezzi:2017:SC

[Ghe17] C. Ghezzi. Of software and change. Journal of Software: Evolution and Process, 29(9):??, September 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garrido:2013:ECP

[GJ13] Alejandra Garrido and Ralph Johnson. Embracing the C preprocessor during refactoring. Journal of Software: Evolution and Process, 25(12):1285– 1304, December 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gilal:2018:FEC

[GJC^{+}18] Abdul Rehman Gilal, Jafreezal Jaafar, Luiz Fernando Capretz, Mazni Omar, Shuib Basri, and Izzatdin Abdul Aziz. Finding an effective classification technique to develop a software team composition model. Journal of Software: Evolution and Process, 30(1):??, January 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gode:2013:SCE

[GK13] Nils Göde and Rainer Koschke. Studying clone evolution using incremental clone detection. Journal of Software: Evolution and Process, 25(2):165– 192, February 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ghannem:2018:MRE

[GKHE18] Adnane Ghannem, Marouane Kessentini, Mohammad Salah Hamdi, and Ghizlane El Boussaidi. Model refactoring by example: a multi-objective search based software engineering approach. Journal of Software: Evolution and Process, 30(4):e1916:1– e1916:??, April 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gotel:2012:QIG

[GKS^{+}12] Olly Gotel, Vidya Kulkarni, Moniphal Say, Christelle Scharff, and Thanwadee Sunetnanta. Quality indicators on global software development projects: does ‘getting to know you’ really matter? Journal of Software: Evolution and Process, 24(2):169– 184, March 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gao:2020:UGM

[GLL^{+}20] Shanquan Gao, Lei Liu, Yuzhou Liu, Huaxiao Liu, and Yihui Wang. Updating the goal model with user reviews for the evolution of an app. Journal of Software: Evolution and Process, 32(8): e2257:1–e2257:??, August 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gao:2021:ARB

[GLL^{+}21] Shanquan Gao, Lei Liu, Yuzhou Liu, Huaxiao Liu, Yihui Wang, and Peixun Liu. App recommendation based on both quality and security. Journal of Software: Evolution and Process, 33(3): e2325:1–e2325:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Geist:2018:TFS

[GNIS18] Verena Geist, Christine Natschläger, Christa Illibauer, and Klaus-Dieter Schewe. Towards functional safety and security for adaptive and flexible business processes. Journal of Software: Evolution and Process, 30 (5):e1952:1–e1952:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garzas:2013:CSS

[GP13] Javier Garzás and Mark C.

Paulk. A case study of software process improvement with CMMI– DEV and Scrum in Spanish companies. Journal of Software: Evolution and Process, 25(12):1325–1333, December 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garousi:2016:WAS

[GP16] Vahid Garousi and Dietmar Pfahl. When to automate software testing? a decision-support approach based on process simulation. Journal of Software: Evolution and Process, 28(4):272–285, April 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garcia:2012:MSP

[GPM^{+}12] Iván García, Carla Pacheco, Eloy Mendoza, Jose A. Calvo-Manzano, Gonzalo Cuevas, and Tomás San Feliu. Managing the software process with a software process improvement tool in a small enterprise. Journal of Software: Evolution and Process, 24 (5):481–491, August 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ghafari:2017:MUT

[GRK17] Mohammad Ghafari, Konstantin Rubinov, and Mohammad Mehdi Pourhashem K. Mining unit test cases to synthesize API usage examples. Journal of Software: Evolution and Process, 29(12):??, December 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Georgiadou:2012:KID

[GS12] Elli Georgiadou and Kerstin V. Siakas. KASL-II: a dynamic four-loop model for knowledge sharing and learning. Journal of Software: Evolution and Process, 24(5):573–583, August 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gautam:2018:CI

[GS18a] Swarnima Singh Gautam and Vrijendra Singh. Cover image. Journal of Software: Evolution and Process, 30(12):e2152:1– e2152:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Gautam:2018:SAS

[GS18b] Swarnima Singh Gautam and Vrijendra Singh. The state-of-the-art in software development effort estimation. Journal of Software: Evolution and Process, 30 (12):e1983:1–e1983:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Goyal:2023:IHA

[GS23] Anjali Goyal and Neetu Sardana. IM_{NRFixer}: a hybrid approach to alleviate class-imbalance problem for predicting the fixability of Non-Reproducible bugs. Journal of Software: Evolution and Process, 35(3): e2290:1–e2290:??, March 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Georgiadou:2020:SDS

[GSB^{+}20] Elli Georgiadou, Kerstin Siakas, Eleni Berki, John Estdale, Harjinder Rahanu, and Margaret Ross. The sociocultural dimension of the Software Process Improvement Manifesto: Pilot validation by experts. Journal of Software: Evolution and Process, 32(11):e2304:1– e2304:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Garousi:2012:EGE

[GW12] Vahid Garousi and Qing Wang. Editorial: Guest editors’ introduction. Journal of Software: Evolution and Process, 24(7):719– 720, November 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Graziotin:2015:DFM

[GWA15] Daniel Graziotin, Xiaofeng Wang, and Pekka Abrahamsson. Do feelings matter? On the correlation of affects and the self-assessed productivity in software engineering. Journal of Software: Evolution and Process, 27(7): 467–487, July 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Huang:2014:EGE

[HAE14] LiGuo Huang, Ove Armbrust, and Guest Editors. Editorial: Guest editors’ introduction. Journal of Software: Evolution and Process, 26(10):867– 868, October 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hachemi:2018:RPP

[HAN18] Asma Hachemi and Mohamed Ahmed-Nacer. Reusing process patterns in software process models modification. Journal of Software: Evolution and Process, 30(8):e1938:1– e1938:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hachemi:2022:PPO

[HAN22] Asma Hachemi and Mohamed Ahmed-Nacer. PO-EML: a Process Orchestration, Execution, and Modeling Language. Journal of Software: Evolution and Process, 34 (6):e2456:1–e2456:??, June 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hebig:2021:HEP

[HAS21] Regina Hebig, Ove Armbrust, and Stanley M. Sutton, Jr. Hybrid and evolving processes for software and systems — ICSSP 2019 special issue. Journal of Software: Evolution and Process, 33(1): e2318:1–e2318:??, January 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hartmann:2016:TMC

[HB16] Herman Hartmann and Jan Bosch. Towards a multi-criteria decision support method for consumer electronics software ecosystems. Journal of Software: Evolution and Process, 28(6):460–482, June 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Haj:2021:AGT

[HBG21] Abdellatif Haj, Youssef Balouki, and Taoufiq Gadi. Automated generation of terminological dictionary from textual business rules. Journal of Software: Evolution and Process, 33(5):e2339:1– e2339:??, May 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Houston:2012:DPD

[HBH12] Dan Houston, Douglas Buettner, and Myron Hecht. Defectivity profiling with dynamic CO-QUALMO: An explication and product quality retrospective. Journal of Software: Evolution and Process, 24(7):803– 814, November 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Holtmann:2016:IIS

[HBM^{+}16] Jörg Holtmann, Ruslan Bernijazov, Matthias Meyer, David Schmelter, and Christian Tschirner. Integrated and iterative systems engineering and software requirements engineering for technical systems. Journal of Software: Evolution and Process, 28 (9):722–743, September 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Huang:2017:UDF

[HCL^{+}17] Yuan Huang, Xiangping Chen, Zhiyong Liu, Xiaonan Luo, and Zibin Zheng. Using discriminative feature in software entities for relevance identification of code changes. Journal of Software: Evolution and Process, 29(7):??, July 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hebig:2017:CBT

[HD17] Regina Hebig and Jesper Derehag. The changing balance of technology and process: A case study on a combined setting of modeldriven development and classical c coding. Journal of Software: Evolution and Process, 29(11): ??, November 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hajnal:2012:DDA

[HF12] Ákos Hajnal and István Forgács. A demand-driven approach to slicing legacy COBOL systems. Journal of Software: Evolution and Process, 24(1):67–82, January 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Harder:2013:CCS

[HG13] Jan Harder and Nils Göde. Cloned code: stable code. Journal of Software: Evolution and Process, 25(10): 1063–1088, October 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Heredia:2014:APA

[HGGASVD14] Alberto Heredia, Javier Garcia-Guzman, Antonio Amescua-Seco, and Manuel Velasco-Diego. Agile practices adapted to mass-market application development. Journal of Software: Evolution and Process, 26(9):818– 828, September 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hale:2012:ETE

[HH12] Joanne E. Hale and David P. Hale. Evaluating testing effectiveness during software evolution: a time-series cross-section approach. Journal of Software: Evolution and Process, 24(1):35–49, January 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Helgesson:2012:RME

[HHW12] Yeni Yuqin Li Helgesson, Martin Höst, and Kim Weyns. A review of methods for evaluation of maturity models for process improvement. Journal of Software: Evolution and Process, 24(4): 436–454, June 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hosni:2019:EFF

[HIA19] Mohamed Hosni, Ali Idri, and Alain Abran. Evaluating filter fuzzy analogy homogeneous ensembles for software development effort estimation. Journal of Software: Evolution and Process, 31(2): e2117:1–e2117:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Holtkamp:2015:HSD

[HLP15] Philipp Holtkamp, Ivan Lau, and Jan Martin Pawlowski. How software development competences change in global settings — an explorative study. Journal of Software: Evolution and Process, 27(1): 50–72, January 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Huang:2015:ISP

[HLWL15] Fuqun Huang, Bin Liu, Shihai Wang, and Qiuying Li. The impact of software process consistency on residual defects. Journal of Software: Evolution and Process, 27(9): 625–646, September 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Horcas:2018:CDR

[HMB^{+}18] José Miguel Horcas, Julien Monteil, Mélanie Bouroche, Mónica Pinto, Lidia Fuentes, and Siobhán Clarke. Contextdependent reconfiguration of autonomous vehicles in mixed traffic. Journal of Software: Evolution and Process, 30(4): e1926:1–e1926:??, April 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hedhli:2022:PBM

[HMB22] Ameni Hedhli, Haithem Mezni, and Lamjed Ben Said. Predictive BPaaS management with quantum and neural computing. Journal of Software: Evolution and Process, 34(2):e2421:1– e2421:??, February 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hyden:2018:MRH

[HMMR18] Paul Hyden, Richard G.

McGrath, Ira S. Moskowitz, and Stephen Russell. Managing risk in high assurance systems by optimizing topological resources. Journal of Software: Evolution and Process, 30(3): e1887:1–e1887:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hegarty:2014:TPL

[HOD^{+}14] Paraic Hegarty, Eoghan O’Leary, Danny Dineen, David Bustard, and Adrian Moore. Transition to a product line infrastructure: experience at Akari Software. Journal of Software: Evolution and Process, 26(9):784– 790, September 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Heidrich:2015:ESP

[HOJ15] Jens Heidrich, Markku Oivo, and Andreas Jedlitschka. Editorial: Software productivity and effort estimation. Journal of Software: Evolution and Process, 27(7): 465–466, July 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Houston:2014:GDF

[Hou14] Dan Houston. A generalized duration forecasting model of test-and-fix cycles. Journal of Software: Evolution and Process, 26(10):877–889, October 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hatton:2017:LTG

[HSvG17] Les Hatton, Diomidis Spinellis, and Michiel van Genuchten. The long-term growth rate of evolving software: Empirical results and implications. Journal of Software: Evolution and Process, 29(5):??, May 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Heck:2014:HTJ

[HZ14] Petra Heck and Andy Zaidman. Horizontal traceability for just-in-time requirements: the case for open source feature requests. Journal of Software: Evolution and Process, 26(12):1280–1296, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Hu:2023:ESP

[HZ23] Mingzhe Hu and Yu Zhang.

An empirical study of the Python/C API on evolution and bug patterns. Journal of Software: Evolution and Process, 35(2): e2507:1–e2507:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Huang:2016:EGE

[HZR16] LiGuo Huang, He Zhang, and Ita Richardson. Editorial: Guest editors’ introduction. Journal of Software: Evolution and Process, 28(4):232–235, April 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Idri:2018:EPS

[IAA18a] Ali Idri, Ibtissam Abnane, and Alain Abran. Evaluating Pred(p) and standardized accuracy criteria in software development effort estimation. Journal of Software: Evolution and Process, 30(4): e1925:1–e1925:??, April 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Idri:2018:SVR

[IAA18b] Ali Idri, Ibtissam Abnane, and Alain Abran. Support vector regression-based imputation in analogybased software development effort estimation. Journal of Software: Evolution and Process, 30(12): e2114:1–e2114:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Imtiaz:2017:DTA

[II17] Salma Imtiaz and Naveed Ikram. Dynamics of task allocation in global software development. Journal of Software: Evolution and Process, 29(1):??, January 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ito:2014:AMC

[IK14] Masao Ito and Koichi Kishida. An approach to manage the concept phase of ISO 26262. Journal of Software: Evolution and Process, 26(9):829– 836, September 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ilyas:2017:SIG

[IK17] Muhammad Ilyas and Siffat Ullah Khan. Software integration in global software development: Challenges for GSD vendors. Journal of Software: Evolution and Process, 29(8): ??, August 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ivanyos:2015:EGS

[ISK15] János Ivanyos and Éva Sándor-Kriszt. ECQA Governance SPICE assessor skills for evaluating integrated risk management scenarios. Journal of Software: Evolution and Process, 27(8):545–554, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Iqbal:2018:PIU

[IZJG18] Md Shahrear Iqbal, Mohammad Zulkernine, Fehmi Jaafar, and Yuan Gu. Protecting Internet users from becoming victimized attackers of click-fraud. Journal of Software: Evolution and Process, 30(3): e1871:1–e1871:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jabbari:2018:TBD

[JbAPT18] Ramtin Jabbari, Nauman bin Ali, Kai Petersen, and Binish Tanveer. Towards a benefits dependency network for DevOps based on a systematic literature review. Journal of Software: Evolution and Process, 30(11): e1957:1–e1957:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jaafar:2014:DAD

[JGHA14] Fehmi Jaafar, Yann-Gaël Guéhéneuc, Sylvie Hamel, and Giuliano Antoniol. Detecting asynchrony and dephase change patterns by mining software repositories. Journal of Software: Evolution and Process, 26(1):77–106, January 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jilani:2022:ASB

[JKIU22] Atif Aftab Jilani, Muhammad Uzair Khan, Muhammad Zohaib Iqbal, and Muhammad Usman. An automated search-based test model generation approach for structural testing of model transformations. Journal of Software: Evolution and Process, 34(11):e2461:1–e2461:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jia:2018:GEF

[JMCC18] Jingdong Jia, Hanlin Mo, Luiz Fernando Capretz, and Zupeng Chen. Grouping environmental factors influencing individual decision-making behavior in software projects: a cluster analysis. Journal of Software: Evolution and Process, 30(1):??, January 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jalali:2018:HAA

[JMR18] Amin Jalali, Fabrizio Maria Maggi, and Hajo A. Reijers. A hybrid approach for aspect-oriented business process modeling. Journal of Software: Evolution and Process, 30(8): e1931:1–e1931:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Johansson:2012:ARP

[JN12] Mika Johansson and Risto Nevalainen. Additional requirements for process assessment in safety-critical software and systems domain. Journal of Software: Evolution and Process, 24 (5):501–510, August 2012.

CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jagroep:2017:EEP

[JPvdW^{+}17] Erik Jagroep, Giuseppe Procaccianti, Jan Martijn van der Werf, Sjaak Brinkkemper, Leen Blom, and Rob van Vliet. Energy efficiency on the product roadmap: An empirical study across releases of a software product. Journal of Software: Evolution and Process, 29(2): ??, February 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jiang:2017:USC

[JPW^{+}17] Qingtao Jiang, Xin Peng, Hai Wang, Zhenchang Xing, and Wenyun Zhao. Understanding systematic and collaborative code changes by mining evolutionary trajectory patterns. Journal of Software: Evolution and Process, 29 (3):??, March 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jimenez-Ramirez:2023:ATR

[JRCMWE23] Andres Jiménez-Ramírez, Jesús Chacón-Montero, Tomasz Wojdynsky, and José González Enríquez. Automated testing in robotic process automation projects. Journal of Software: Evolution and Process, 35(3):e2259:1– e2259:??, March 2023. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Janes:2013:MCR

[JRSS13] Andrea Janes, Tadas Remencius, Alberto Sillitti, and Giancarlo Succi. Managing changes in requirements: an empirical investigation. Journal of Software: Evolution and Process, 25(12):1273–1283, December 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jilani:2022:DEF

[JSI^{+}22] Atif Aftab Jilani, Salman Sherin, Sidra Ijaz, Muhammad Zohaib Iqbal, and Muhammad Uzair Khan. Deriving and evaluating a fault model for testing data science applications. Journal of Software: Evolution and Process, 34 (5):e2449:1–e2449:??, May 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jung:2012:PAR

[Jun12] Ho-Won Jung. Process attribute rating and sensitivity analysis in process assessment. Journal of Software: Evolution and Process, 24(4):401–419, June 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jalali:2012:GSE

[JW12] Samireh Jalali and Claes Wohlin. Global software engineering and agile practices: a systematic review. Journal of Software: Evolution and Process, 24(6):643–659, October 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Jabangwe:2016:MIQ

[JWP^{+}16] Ronald Jabangwe, Claes Wohlin, Kai Petersen, Darja Šmite, and Jürgen Börstler. A method for investigating the quality of evolving object-oriented software using defects in global software development projects. Journal of Software: Evolution and Process, 28(8):622– 641, August 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khan:2020:SLR

[KA20] Arif Ali Khan and Muhammad Azeem Akbar. Systematic literature review and empirical investigation of motivators for requirements change management process in global software development. Journal of Software: Evolution and Process, 32(4):e2242:1– e2242:??, April 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khelladi:2017:CMP

[KBHG17] Djamel Eddine Khelladi, Reda Bendraou, Regina Hebig, and Marie-Pierre Gervais. Coadapting multidimension process properties. Journal of Software: Evolution and Process, 29(11):??, November 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Korsaa:2012:SME

[KBM^{+}12] Morten Korsaa, Miklos Biro, Richard Messnarz, Jörn Johansen, Detlef Vohwinkel, Risto Nevalainen, and Tomas Schweigert. The SPI manifesto and the ECQA SPI manager certification scheme. Journal of Software: Evolution and Process, 24(5):525– 540, August 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kaur:2022:EFA

[KC22] Rajdeep Kaur and Kuljit Kaur Chahal. Exploring factors affecting developer abandonment of open source software projects. Journal of Software: Evolution and Process, 34(9): e2484:1–e2484:??, September 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kirbas:2017:RBE

[KCH^{+}17] Serkan Kirbas, Bora Caglayan, Tracy Hall, Steve Counsell, David Bowes, Alper Sen, and Ayse Bener. The relationship between evolutionary coupling and defects in large industrial software. Journal of Software: Evolution and Process, 29(4):??, April 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kienle:2013:ESI

[KDK13] Holger M. Kienle, Giuseppe Di Lucca, and Kostas Kontogiannis. Editorial: Special issue: selected papers from the 12th International Symposium on Web Systems Evolution (WSE 2010). Journal of Software: Evolution and Process, 25(8):781–787, August 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kuhrmann:2014:MSF

[KFT14] Marco Kuhrmann, Daniel Méndez Fernández, and Michaela Tiessler. A mapping study on the feasibility of method engineering. Journal of Software: Evolution and Process, 26(12):1053– 1073, December 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kuhrmann:2016:UVO

[KFT16] Marco Kuhrmann, Daniel Méndez Fernández, and Thomas Ternité. On the use of variability operations in the V-Modell XT software process line. Journal of Software: Evolution and Process, 28(4):241–253, April 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kula:2013:MPA

[KFYI13] Raula Gaikovina Kula, Kyohei Fushida, Norihiro Yoshida, and Hajimu Iida. Micro process analysis of maintenance effort: an open source software case study using metrics based on program slicing. Journal of Software: Evolution and Process, 25(9):935– 955, September 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kagdi:2012:ACR

[KGPH12] Huzefa Kagdi, Malcom Gethers, Denys Poshyvanyk, and Maen Hammad. Assigning change requests to software developers. Journal of Software: Evolution and Process, 24(1):3–33, January 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khurum:2013:SVM

[KGW13] Mahvish Khurum, Tony Gorschek, and Magnus Wilson. The software value map — an exhaustive collection of value aspects for the development of software intensive products. Journal of Software: Evolution and Process, 25(7): 711–741, July 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kalenda:2018:SAL

[KHR18] Martin Kalenda, Petr Hyna, and Bruno Rossi. Scaling agile in large organizations: Practices, challenges, and success factors. Journal of Software: Evolution and Process, 30(10): e1954:1–e1954:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khan:2021:AHP

[KHZ21] Abdul Wahid Khan, Imtiaz Hussain, and Muhammad Zamir. Analytic hierarchy process-based prioritization framework for vendor’s reliability challenges in global software development. Journal of Software: Evolution and Process, 33(3):e2310:1– e2310:??, March 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kasse:2014:MDB

[KJ14] Tim Kasse and Jørn Johansen. Maturity differences between customer and supplier — challenges, problems, and possible solutions. Journal of Software: Evolution and Process, 26(3):295–305, March 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Korsaa:2013:PAM

[KJS^{+}13] Morten Korsaa, Jörn Johansen, Tomas Schweigert, Detlef Vohwinkel, Richard Messnarz, Risto Nevalainen, and Miklos Biro. The people aspects in modern process improvement management approaches. Journal of Software: Evolution and Process, 25(4): 381–391, April 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kaynak:2014:ERI

[KK14] Onur Kaynak and N. Alpay Karagöz. Experience report: implementation of a multi-standard compliant process improvement program. Journal of Software: Evolution and Process, 26(5):488–495, May 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khan:2021:IPS

[KKK^{+}21] Abdul Wahid Khan, Maseeh Ullah Khan, Javed Ali Khan, Javed Khan, and Wresham Gul. Identification and prioritization of security challenges of big data on cloud computing based on SLR: a fuzzy-TOPSIS analysis approach. Journal of Software: Evolution and Process, 33(12): e2387:1–e2387:??, December 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khan:2019:GRS

[KKN^{+}19] Arif Ali Khan, Jacky Keung, Mahmood Niazi, Shahid Hussain, and Mohammad Shameem. GSEPIM: a roadmap for software process assessment and improvement in the domain of global software development. Journal of Software: Evolution and Process, 31(1): e1988:1–e1988:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kelemen:2012:ICM

[KKT12] Zádor Dániel Kelemen, Rob Kusters, and Jos Trienekens. Identifying criteria for multimodel software process improvement solutions — based on a review of current problems and initiatives. Journal of Software: Evolution and Process, 24(8):895– 909, December 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kulesz:2018:SGA

[KKW18] Daniel Kulesz, Verena Käfer, and Stefan Wagner. Spreadsheet guardian: an approach to protecting semantic correctness throughout the evolution of spreadsheets. Journal of Software: Evolution and Process, 30(9): e1934:1–e1934:??, September 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khalsa:2018:ECP

[KL18] Sunint Kaur Khalsa and Yvan Labiche. Extending category partition’s base choice criterion to better support constraints. Journal of Software: Evolution and Process, 30(3): e1868:1–e1868:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kedji:2014:SCD

[KLC^{+}14] Komlan Akpédjé Kedji, Redouane Lbath, Bernard Coulette, Mahmoud Nassar, Laurent Baresse, and Florin Racaru. Supporting collaborative development using process models: a tooled integrationfocused approach. Journal of Software: Evolution and Process, 26(10):890– 909, October 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khomh:2017:EGE

[KLG17] Foutse Khomh, David Lo, and Michael W. Godfrey. Editorial: Guest editor’s introduction to the special issue on Source Code Analysis and Manipulation (SCAM 2015). Journal of Software: Evolution and Process, 29(12): ??, December 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khan:2020:CEA

[KLWA20] Javed Ali Khan, Lin Liu, Lijie Wen, and Raian Ali. Conceptualising, extracting and analysing requirements arguments in users’ forums: The CrowdRE-Arg framework. Journal of Software: Evolution and Process, 32(12): e2309:1–e2309:??, December 2020. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Kandil:2017:CBT

[KMB17] Passant Kandil, Sherin Moussa, and Nagwa Badr. Cluster-based test cases prioritization and selection technique for agile regression testing. Journal of Software: Evolution and Process, 29(6):??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kuang:2015:CMD

[KMH^{+}15] Hongyu Kuang, Patrick Mäder, Hao Hu, Achraf Ghabi, LiGuo Huang, Jian Lü, and Alexander Egyed. Can method data dependencies support the assessment of traceability between requirements and source code? Journal of Software: Evolution and Process, 27(11):838– 866, November 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kowalczyk:2022:SAF

[KMP22] Michał Kowalczyk, Bartosz Marcinkowski, and Adam Przybyłek. Scaled agile framework. Dealing with software processrelated challenges of a financial group with the action research approach. Journal of Software: Evolution and Process, 34 (6):e2455:1–e2455:??, June 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kurtel:2013:CSM

[KO13] Kaan Kurtel and Murat Ozemre. Cohesive software measurement planning framework using ISO standards: a case study from logistics service sector. Journal of Software: Evolution and Process, 25(7):663–679, July 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kessentini:2016:EIS

[KÓ16] Marouane Kessentini and Mel Ó Cinnéide. Editorial: Introduction to the special issue on search-based software engineering (Nas-BASE 2015). Journal of Software: Evolution and Process, 28(12):1024, December 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kuhrmann:2017:ISI

[KOPR17] Marco Kuhrmann, Rory V. O’Connor, Dewayne E. Perry, and David Raffo. ICSSP 2016 — special issue introduction. Journal of Software: Evolution and Process, 29(11):??, November 2017. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Koschke:2013:IRA

[Kos13] Rainer Koschke. Incremental reflexion analysis. Journal of Software: Evolution and Process, 25(6): 601–637, June 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Koschke:2014:LSI

[Kos14] Rainer Koschke. Largescale inter-system clone detection using suffix trees and hashing. Journal of Software: Evolution and Process, 26(8):747– 769, August 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kose:2021:BPM

[Kos21] Busra Ozdenizci Kose.

Business process management approach for improving agile software process and agile maturity. Journal of Software: Evolution and Process, 33(4): e2331:1–e2331:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khurum:2014:EVS

[KPG14] Mahvish Khurum, Kai Petersen, and Tony Gorschek. Extending value stream mapping through waste definition beyond customer perspective. Journal of Software: Evolution and Process, 26(12):1074– 1105, December 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Koschke:2015:PSS

[KR15] Rainer Koschke and Juergen Rilling. Preface special section on software clones (IWSC’13). Journal of Software: Evolution and Process, 27(6):401, June 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kramer:2013:AMC

[Kra13] Anne Kramer. Automotive and medical: can we learn from each other? Journal of Software: Evolution and Process, 25(4): 373–379, April 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kpodjedo:2013:SSE

[KRG^{+}13] Segla Kpodjedo, Filippo Ricca, Philippe Galinier, Giuliano Antoniol, and Yann-Gaël Guéhéneuc. Studying software evolution of large objectoriented software systems using an ETGM algorithm. Journal of Software: Evolution and Process, 25(2):139–163, February 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Karim:2016:EIS

[KRR^{+}16] Muhammad Rezaul Karim, Günther Ruhe, Md. Mainur Rahman, Vahid Garousi, and Thomas Zimmermann. An empirical investigation of single-objective and multiobjective evolutionary algorithms for developer’s assignment to bugs. Journal of Software: Evolution and Process, 28 (12):1025–1060, December 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kruchten:2013:CAS

[Kru13] Philippe Kruchten. Contextualizing agile software development. Journal of Software: Evolution and Process, 25(4):351– 361, April 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Khan:2020:MDM

[KS20] Arif Ali Khan and Mohammad Shameem. Multicriteria decision-making taxonomy for DevOps challenging factors using analytical hierarchy process. Journal of Software: Evolution and Process, 32(10): e2263:1–e2263:??, October 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kotejoshyer:2014:RBR

[KST14] Badari Kotejoshyer, Birendra Singh, and Suhas Tittibha. Rule-based requirements management methodology. Journal of Software: Evolution and Process, 26(3):321– 328, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kelemen:2016:ERT

[KTHS16] Zádor Dániel Kelemen, Balázs Tódor, Sándor Hodosi, and Ákos Somfai. EUROSPI 2015: Refactoring technical support to reduce interrupts of developers. Journal of Software: Evolution and Process, 28 (11):960–968, November 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Klunder:2022:HAC

[KTP22] Jil Klünder, Felix Trommer, and Nils Prenner. How agile coaches create an agile mindset in development teams: Insights from an interview study. Journal of Software: Evolution and Process, 34(12): e2491:1–e2491:??, December 2022. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Kaushik:2016:IFA

[KTYK16] Anupama Kaushik, Devendra Kr. Tayal, Kalpana Yadav, and Arvinder Kaur. Integrating firefly algorithm in artificial neural network models for accurate software cost predictions. Journal of Software: Evolution and Process, 28(8):665–688, August 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Kuhrmann:2015:CSP

[Kuh15] Marco Kuhrmann. Crafting a software process improvement approach — a retrospective systematization. Journal of Software: Evolution and Process, 27(2):114–145, February 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lauer:2018:RCR

[LAF^{+}18] Michael Lauer, Matthieu Amy, Jean-Charles Fabre, Matthieu Roy, William Excoffon, and Miruna Stoicescu. Resilient computing on ROS using adaptive fault tolerance. Journal of Software: Evolution and Process, 30(3): e1917:1–e1917:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Litoiu:2013:PEF

[LB13] Marin Litoiu and Cornel Barna. A performance evaluation framework for Web applications. Journal of Software: Evolution and Process, 25(8):871– 890, August 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lim:2023:VMF

[LC23] Chong Hon Lim and Jeng Feng Chin. V-model with fuzzy quality function deployments for mobile application development. Journal of Software: Evolution and Process, 35 (1):e2518:1–e2518:??, January 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lepmets:2015:DMM

[LCM^{+}15] Marion Lepmets, Paul Clarke, Fergal McCaffery, Anita Finnegan, and Alec Dorling. Development of MDevSPICE(R) — the medical device software process assessment framework. Journal of Software: Evolution and Process, 27(8):565–572, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lema:2015:ISM

[LCMCPA15] Lohana Lema, José-Antonio Calvo-Manzano, Ricardo Colomo-Palacios, and Magdalena Arcilla. ITIL in small to medium-sized enterprises software companies: towards an implementation sequence. Journal of Software: Evolution and Process, 27(8):528– 538, August 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lonetti:2022:DTS

[LdONB22] Francesca Lonetti, Vânia de Oliveira Neves, and Antonia Bertolino. Designing and testing systems of systems: From variability models to test cases passing through desirability assessment. Journal of Software: Evolution and Process, 34(10): e2427:1–e2427:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lenberg:2019:MVS

[LFT19] Per Lenberg, Robert Feldt, and Lars Göran Wallgren Tengberg. Misaligned values in software engineering organizations. Journal of Software: Evolution and Process, 31(3): e2148:1–e2148:??, March 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Larrucea:2020:AEM

[LGNEAO20] Xabier Larrucea, Pablo González-Nalda, Ismael Etxeberria-Agiriano, and Mari Carmen Otero. Analysing encryption mechanisms and functional safety in a ROS-based architecture. Journal of Software: Evolution and Process, 32(2): e2224:1–e2224:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Li:2017:EOS

[LGS^{+}17] Deguang Li, Bing Guo, Yan Shen, Junke Li, and Yanhui Huang. The evolution of open-source mobile applications: An empirical study. Journal of Software: Evolution and Process, 29 (7):??, July 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Li:2016:ELW

[LHA^{+}16] Wenbin Li, Jane Huffman Hayes, Giulio Antoniol, Yann-Gaël Guéhéneuc, and Bram Adams. Error leakage and wasted time: sensitivity and effort analysis of a requirements consistency checking process. Journal of Software: Evolution and Process, 28 (12):1061–1080, December 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Liu:2018:DCD

[LHEG18] Xiaoyu Liu, LiGuo Huang, Alexander Egyed, and Jidong Ge. Do code data sharing dependencies support an early prediction of software actual change impact set? Journal of Software: Evolution and Process, 30(11): e1960:1–e1960:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lopez-Herrejon:2018:SMS

[LHIE18] Roberto Erick Lopez-Herrejon, Sheny Illescas, and Alexander Egyed. A systematic mapping study of information visualization for software product line engineering. Journal of Software: Evolution and Process, 30(2): e1912:1–e1912:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Leupolz:2018:QQS

[LHR18] Johannes Leupolz, Axel Habermaier, and Wolfgang Reif. Quantitative and qualitative safety analysis of a hemodialysis machine with S#. Journal of Software: Evolution and Process, 30 (5):e1942:1–e1942:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lindholm:2015:IUP

[Lin15] Christin Lindholm. Involving user perspective in a software risk management process. Journal of Software: Evolution and Process, 27(12):953–975, December 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lucia:2014:ECS

[LLJ^{+}14] Lucia Lucia, David Lo, Lingxiao Jiang, Ferdian Thung, and Aditya Budi. Extended comprehensive study of association measures for fault localization. Journal of Software: Evolution and Process, 26(2): 172–219, February 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Liu:2021:API

[LLLL21] Lei Liu, Xun Li, Yuzhou Liu, and Huaxiao Liu. Application programming interface recommendation according to the knowledge indexed by app feature mined from app stores. Journal of Software: Evolution and Process, 33(11):e2380:1– e2380:??, November 2021.

CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Liu:2018:ARG

[LLLW18] Yuzhou Liu, Lei Liu, Huaxiao Liu, and Xiaoyu Wang. Analyzing reviews guided by App descriptions for the software development and evolution. Journal of Software: Evolution and Process, 30(12): e2112:1–e2112:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Li:2022:USA

[LLSA22] Ruiyin Li, Peng Liang, Mohamed Soliman, and Paris Avgeriou. Understanding software architecture erosion: a systematic mapping study. Journal of Software: Evolution and Process, 34(3): e2423:1–e2423:??, March 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Liu:2020:EFS

[LLT20] Geng Liu, Luigi Lavazza, and Davide Tosi. Evolution of functional size measures through ICONIX process phases. Journal of Software: Evolution and Process, 32 (5):e2240:1–e2240:??, May 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lahmar:2018:MSC

[LM18] Fatma Lahmar and Haithem Mezni. Multicloud service composition: a survey of current approaches and issues. Journal of Software: Evolution and Process, 30 (10):e1947:1–e1947:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lago:2012:ESL

[LMB12] Patricia Lago, Henry Muccini, and Muhammad Ali Babar. An empirical study of learning by osmosis in global software engineering. Journal of Software: Evolution and Process, 24(6):693–706, October 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lepmets:2016:DBM

[LMC16] Marion Lepmets, Fergal McCaffery, and Paul Clarke. Development and benefits of MDevSPICE(R) the medical device software process assessment framework. Journal of Software: Evolution and Process, 28(9):800– 816, September 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lamersdorf:2012:RBM

[LMdVT^{+}12] Ansgar Lamersdorf, Jürgen Münch, Alicia Fernández del Viso Torre, Carlos Rebate Sánchez, Markus Heinz, and Dieter Rombach. A rule-based model for customized risk identification and evaluation of task assignment alternatives in distributed software development projects. Journal of Software: Evolution and Process, 24(6):661– 675, October 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lu:2018:IQA

[LML^{+}18] Yao Lu, Xinjun Mao, Zude Li, Yang Zhang, Tao Wang, and Gang Yin. Internal quality assurance for external contributions in GitHub: an empirical investigation. Journal of Software: Evolution and Process, 30(4): e1918:1–e1918:??, April 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lizcano:2023:EUM

[LMOLG23] David Lizcano, Andrés-Leonardo Martínez-Ortíz, Genoveva López, and Arnaud Grignard. End-user modeling of quality for web components. Journal of Software: Evolution and Process, 35(3): e2256:1–e2256:??, March 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lago:2017:ERC

[LP17] Patricia Lago and Birgit Penzenstadler. Editorial: Reality check for software engineering for sustainability — pragmatism required. Journal of Software: Evolution and Process, 29(2):??, February 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Larrucea:2014:IAM

[LS14] Xabier Larrucea and Izaskun Santamaria. An industrial assessment for a multimodel framework. Journal of Software: Evolution and Process, 26(9): 837–845, September 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Larrucea:2019:CSC

[LS19] Xabier Larrucea and Izaskun Santamaría. Correlations study and clustering from SPI experiences in small settings. Journal of Software: Evolution and Process, 31(1): e1989:1–e1989:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Landman:2016:EAR

[LSBV16] Davy Landman, Alexander Serebrenik, Eric Bouwers, and Jurgen J. Vinju. Empirical analysis of the relationship between CC and SLOC in a large corpus of Java methods and C functions. Journal of Software: Evolution and Process, 28(7):589–618, July 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic). See corrigendum [LSBV17].

Landman:2017:CEA

[LSBV17] Davy Landman, Alexander Serebrenik, Eric Bouwers, and Jurgen Vinju. Corrigendum: Empirical analysis of the relationship between CC and SLOC in a large corpus of Java methods and C functions published on 9 December 2015. Journal of Software: Evolution and Process, 29(10):??, October 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic). See [LSBV16].

Larrucea:2016:EAI

[LSCP16] Xabier Larrucea, Izaskun Santamaría, and Ricardo Colomo-Palacios. EU-ROSPI 2015: Assessing ISO/IEC 29110 by means of ITMark: results from an experience factory. Journal of Software: Evolution and Process, 28(11):969– 980, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Larrucea:2020:MSD

[LSFG20] Xabier Larrucea, Izaskun Santamaria, and Borja Fernandez-Gauna. Managing security debt across PLC phases in a VSE context. Journal of Software: Evolution and Process, 32(3):e2214:1– e2214:??, March 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Leotta:2016:RAG

[LSRT16] Maurizio Leotta, Andrea Stocco, Filippo Ricca, and Paolo Tonella. Robula+: an algorithm for generating robust XPath locators for web testing. Journal of Software: Evolution and Process, 28(3): 177–204, March 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Lagerstrom:2013:ISD

[LSW13] Robert Lagerström, Ulf Sporrong, and Anders Wall. Increasing software development efficiency and maintainability for complex industrial systems — a case study. Journal of Software: Evolution and Process, 25(3): 285–301, March 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Liu:2016:WTV

[LWY^{+}16] Xiao Liu, Dingxian Wang, Dong Yuan, Futian Wang, and Yun Yang. Workflow temporal verification for monitoring parallel business processes. Journal of Software: Evolution and Process, 28(4): 286–302, April 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Li:2012:IST

[LYL^{+}12] Qi Li, Ye Yang, Mingshu Li, Qing Wang, Barry W. Boehm, and Chenyong Hu. Improving software testing process: feature prioritization to make winners of success-critical stakeholders. Journal of Software: Evolution and Process, 24(7):783–801, November 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mohammad:2012:CBD

[MA12] Mubarak Mohammad and Vangalur Alagar. A component-based development process for trustworthy systems. Journal of Software: Evolution and Process, 24(7):815– 835, November 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mittas:2020:DDB

[MA20] Nikolaos Mittas and Lefteris Angelis. Datadriven benchmarking in software development effort estimation: the few define the bulk. Journal of Software: Evolution and Process, 32(9): e2258:1–e2258:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Milewski:2012:ESI

[MAD12] Allen Milewski, Alberto Avritzer, and Yael Dubinsky. Editorial: Special issue on global software engineering. Journal of Software: Evolution and Process, 24(6):603–604, October 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Medini:2014:SAL

[MAD^{+}14] Soumaya Medini, Venera Arnaoudova, Massimiliano Di Penta, Giuliano Antoniol, Yann-Gaël Guéhéneuc, and Paolo Tonella. SCAN: an approach to label and relate execution trace segments. Journal of Software: Evolution and Process, 26(11): 962–995, November 2014.

CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mordal:2013:SQM

[MAL^{+}13] Karine Mordal, Nicolas Anquetil, Jannik Laval, Alexander Serebrenik, Bogdan Vasilescu, and Stéphane Ducasse. Software quality metrics aggregation in industry. Journal of Software: Evolution and Process, 25(10):1117–1135, October 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mumtaz:2019:SUM

[MAMN19] Haris Mumtaz, Mohammad Alshayeb, Sajjad Mahmood, and Mahmood Niazi. A survey on UML model smells detection techniques for software refactoring. Journal of Software: Evolution and Process, 31(3): e2154:1–e2154:??, March 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Miranskyy:2018:DEE

[MAzGB18] A. V. Miranskyy, Z. Alzanbouri, D. Godwin, and A. B. Bener. Database engines: Evolution of greenness. Journal of Software: Evolution and Process, 30(4):e1915:1– e1915:??, April 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Martini:2017:IAT

[MB17] Antonio Martini and Jan Bosch. On the interest of architectural technical debt: Uncovering the contagious debt phenomenon. Journal of Software: Evolution and Process, 29(10): ??, October 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Majumdar:2022:AEC

[MBD^{+}22] Srijoni Majumdar, Ayush Bansal, Partha Pratim Das, Paul D. Clough, Kausik Datta, and Soumya Kanti Ghosh. Automated evaluation of comments to aid software maintenance. Journal of Software: Evolution and Process, 34 (7):e2463:1–e2463:??, July 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mashkoor:2018:SFS

[MBMCP18] Atif Mashkoor, Miklós Biró, Richard Messnarz, and Ricardo Colomo-Palacios. Selected functional safety and cybersecurity concerns in system, software, and service process improvement and innovation. Journal of Software: Evolution and Process, 30 (5):e1955:1–e1955:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Miranda:2019:URE

[MBMR19] Enrique A. Miranda, Mario Berón, Germán Montejano, and Daniel Riesco. Using reverse engineering techniques to infer a system use case model. Journal of Software: Evolution and Process, 31(2): e2121:1–e2121:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mens:2014:ECS

[MC14] T. Mens and A. Cleve. Editorial: CSMR 2012 special issue — guest editorial. Journal of Software: Evolution and Process, 26 (8):729, August 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mori:2018:SSM

[MCL^{+}18] Marco Mori, Andrea Ceccarelli, Paolo Lollini, Bernhard Frömel, Francesco Brancati, and Andrea Bondavalli. Systems-of-systems modeling using a comprehensive viewpointbased SysML profile. Journal of Software: Evolution and Process, 30(3): e1878:1–e1878:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

McQuaid:2012:SDU

[McQ12] Patricia A. McQuaid. Software disasters — understanding the past, to improve the future. Journal of Software: Evolution and Process, 24(5):459– 470, August 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mora-Cantallops:2020:EPC

[MCSGBSA20] Marçal Mora-Cantallops, Miguel-Ángel Sicilia, Elena García-Barriocanal, and Salvador Sánchez-Alonso. Evolution and prospects of the Comprehensive R Archive Network (CRAN) package ecosystem. Journal of Software: Evolution and Process, 32(11): e2270:1–e2270:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Moraga:2017:GGC

[MdGC^{+}17] Ma Ángeles Moraga, Ignacio García-Rodríguez de Guzmán, Coral Calero, Timo Johann, Gianantonio Me, Harald Münzel, and Julia Kindelsberger. GreCo: Green code of ethics. Journal of Software: Evolution and Process, 29(2):??, February 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Moran:2019:TMP

[MdlRT19] Jesús Morán, Claudio de la Riva, and Javier Tuya. Testing MapReduce programs: a systematic mapping study. Journal of Software: Evolution and Process, 31(3):e2120:1– e2120:??, March 2019. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Medina-Dominguez:2015:CTC

[MDSSMS15] Fuensanta Medina-Dominguez, Maria-Isabel Sanchez-Segura, Ana Moreno, and Daniel Santin. Collaborative tools: computer science students’ skills versus software industry needs. Journal of Software: Evolution and Process, 27(3): 221–235, March 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mercaldo:2018:ESE

[MDV^{+}18] Francesco Mercaldo, Andrea Di Sorbo, Corrado Aaron Visaggio, Aniello Cimitile, and Fabio Martinelli. An exploratory study on the evolution of Android malware quality. Journal of Software: Evolution and Process, 30(11): e1978:1–e1978:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mas:2012:LLI

[MFA12] Antonia Mas, Bartomeu Fluxà, and Esperança Amengual. Lessons learned from an ISO/IEC 15504 SPI programme in a company. Journal of Software: Evolution and Process, 24 (5):493–500, August 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Misra:2014:SMS

[MFCP14] Sanjay Misra, Luis Fernández, and Ricardo Colomo-Palacios. A simplified model for software inspection. Journal of Software: Evolution and Process, 26 (12):1297–1315, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Merino:2018:TAV

[MGN18] Leonel Merino, Mohammad Ghafari, and Oscar Nierstrasz. Towards actionable visualization for software developers. Journal of Software: Evolution and Process, 30(2): e1923:1–e1923:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Martin:2014:MSD

[MGUA14] Diego Martín, Javier García Guzmán, Julián Urbano, and Antonio Amescua. Modeling software development practices using reusable project patterns: a case study. Journal of Software: Evolution and Process, 26(3):339– 349, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mishra:2017:EFD

[MGW^{+}17] Alok Mishra, Juan Garbajosa, Xiaofeng Wang, Jan Bosch, and Pekka Abrahamsson. Editorial: Future directions in Agile research: Alignment and divergence between research and practice. Journal of Software: Evolution and Process, 29(6):??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mensah:2022:CSI

[MK22] Solomon Mensah and Patrick Kwaku Kudjo. A classification scheme to improve conclusion instability using Bellwether moving windows. Journal of Software: Evolution and Process, 34(9): e2488:1–e2488:??, September 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mahmood:2020:SRS

[MKA20] Yasir Mahmood, Nazri Kama, and Azri Azmi. A systematic review of studies on use case points and expert-based estimation of software development effort. Journal of Software: Evolution and Process, 32 (7):e2245:1–e2245:??, July 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Messnarz:2018:EAS

[MKMW18] Richard Messnarz, Christian Kreiner, Georg Macher, and Alastair Walker. Extending automotive SPICE 3.0 for the use in ADAS and future self-driving service architectures. Journal of Software: Evolution and Process, 30 (5):e1948:1–e1948:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Minhas:2023:UGQ

[MKPB23] Nasir Mehmood Minhas, Thejendar Reddy Koppula, Kai Petersen, and Jürgen Börstler. Using goal-question-metric to compare research and practice perspectives on regression testing. Journal of Software: Evolution and Process, 35(2): e2506:1–e2506:??, February 2023. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Ma:2022:VBR

[MLCW22] Shang-Pin Ma, I-Hsiu Liu, Chun-Yu Chen, and Yu-Te Wang. Version-based and risk-enabled testing, monitoring, and visualization of microservice systems. Journal of Software: Evolution and Process, 34(10): e2429:1–e2429:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

McBurney:2016:AFD

[MLM16] Paul W. McBurney, Cheng Liu, and Collin McMillan. Automated feature discovery via sentence selection and source code summarization. Journal of Software: Evolution and Process, 28(2):120– 145, February 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mesquida:2014:PMI

[MM14] Antoni-Lluís Mesquida and Antonia Mas. A project management improvement program according to ISO/IEC 29110 and PMBOK(R). Journal of Software: Evolution and Process, 26(9):846– 854, September 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

McHugh:2014:AAP

[MMC14] Martin McHugh, Fergal McCaffery, and Valentine Casey. Adopting agile practices when developing software for use in the medical domain. Journal of Software: Evolution and Process, 26(5): 504–512, May 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mas:2016:ELL

[MMJB16] Antonia Mas, Antoni-Lluís Mesquida, Miloš Jovanović, and Joan Barceló. EUROSPI 2015: Lessons learned from 4year empowering project management with TA-LAIA OpenPPM. Journal of Software: Evolution and Process, 28(11):981– 989, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

MacMahon:2016:MAF

[MMK16] Silvana Togneri MacMahon, Fergal McCaffery, and Frank Keenan. The MedITNet assessment framework: development and validation of a framework for improving risk management of medical IT networks. Journal of Software: Evolution and Process, 28(9):817– 834, September 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Minhas:2020:SMT

[MMPN20] Nasir Mehmood Minhas, Sohaib Masood, Kai Petersen, and Aamer Nadeem. A systematic mapping of test case generation techniques using UML interaction diagrams. Journal of Software: Evolution and Process, 32(6):e2235:1– e2235:??, June 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Moonen:2016:EIS

[MP16] Leon Moonen and Lori Pollock. Editorial: Introduction to the special issue on software maintenance and evolution. Journal of Software: Evolution and Process, 28(7): 510–511, July 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Martini:2016:MCS

[MPB16] Antonio Martini, Lars Pareto, and Jan Bosch. A multiple case study on the inter-group interaction speed in large, embedded software companies employing agile. Journal of Software: Evolution and Process, 28(1):4–26, January 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Munoz:2020:AIS

[MPM^{+}20] Mirna Muñoz, Adriana Peña, Jezreel Mejía, Gloria Piedad Gasca-Hurtado, María Clara Gómez-Alvarez, and Claude Y. Laporte. Analysis of 13 implementations of the software engineering management and engineering basic profile guide of ISO/IEC 29110 in very small entities using different life cycles. Journal of Software: Evolution and Process, 32(11): e2300:1–e2300:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Majchrowski:2016:ESD

[MPS^{+}16] Annick Majchrowski, Christophe Ponsard, Sanae Saadaoui, Jacques Flamand, and Jean-Christophe Deprez. EUROSPI 2015: Software development practices in small entities: an ISO 29110-based survey. Journal of Software: Evolution and Process, 28(11):990– 999, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Machado:2022:ABC

[MPSM22] João Paulo L. Machado, Elder V. P. Paula-Sobrinho, and Marcelo A. Maia. Anti-bloater class restructuring: an exploratory study. Journal of Software: Evolution and Process, 34(3):e2431:1– e2431:??, March 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mayvan:2020:BSD

[MRJ20] Bahareh Bafandeh Mayvan, Abbas Rasoolzadegan, and Abbas Javan Jafari. Bad smell detection using quality metrics and refactoring opportunities. Journal of Software: Evolution and Process, 32(8): e2255:1–e2255:??, August 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Messnarz:2014:SRA

[MSB^{+}14] Richard Messnarz, Miguel-Angel Sicilia, Miklos Biro, Elena García-Barriocanal, Miguel Garre-Rubio, Kerstin Siakas, and Adrienne Clarke. Social responsibility aspects supporting the success of SPI. Journal of Software: Evolution and Process, 26(3): 284–294, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mashkoor:2020:SSC

[MSBE20] Atif Mashkoor, Johannes Sametinger, Miklós Biro, and Alexander Egyed. Security- and safety-critical cyber-physical systems. Journal of Software: Evolution and Process, 32(2): e2239:1–e2239:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mezni:2018:SAS

[MSK18] Haithem Mezni, Mokhtar Sellami, and Jaber Kouki. Security-aware SaaS placement using swarm intelligence. Journal of Software: Evolution and Process, 30(8):e1932:1– e1932:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Miranda:2022:GEI

[MTG22] Breno Miranda, Javier Tuya, and Alejandra Garrido. Guest editors’ introduction to the special issue “Automatic Software Testing from the Trenches”. Journal of Software: Evolution and Process, 34(11): e2510:1–e2510:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

McBride:2014:MAO

[MV14] Tom McBride and Timo Varkoi. A method for aggregating ordinal process assessment measures. Journal of Software: Evolution and Process, 26 (12):1267–1279, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Monasor:2012:CLP

[MVP12] Miguel J. Monasor, Aurora Vizcaíno, and Mario Piattini. Cultural and linguistic problems in GSD: a simulator to train engineers in these issues. Journal of Software: Evolution and Process, 24(6):707– 717, October 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Manjavacas:2020:GSD

[MVRP20] Antonio Manjavacas, Aurora Vizcaíno, Francisco Ruiz, and Mario Piattini. Global software development governance: Challenges and solutions. Journal of Software: Evolution and Process, 32(10): e2266:1–e2266:??, October 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Newman:2017:SCS

[NBCM17] Christian D. Newman, Brian Bartman, Michael L. Collard, and Jonathan I. Maletic. Simplifying the construction of source code transformations via automatic syntactic restructurings. Journal of Software: Evolution and Process, 29 (4):??, April 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Noorian:2017:TAQ

[NBD17] Mahdi Noorian, Ebrahim Bagheri, and Weichang Du. Toward automated quality-centric product line configuration using intentional variability. Journal of Software: Evolution and Process, 29(9): ??, September 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nurdiani:2018:LRF

[NBF18] Indira Nurdiani, Jürgen Börstler, and Samuel A. Fricker. Literature review of flexibility attributes: a flexibility framework for software developing organization. Journal of Software: Evolution and Process, 30(9):e1937:1– e1937:??, September 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nishi:2020:AIV

[ND20] Manziba Akanda Nishi and Kostadin Damevski. Automatically identifying valid API versions for software development tutorials on the Web. Journal of Software: Evolution and Process, 32(4): e2227:1–e2227:??, April 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Norimatsu:2019:FPI

[NEU^{+}19] So Norimatsu, Kiyoshi Endo, Makoto Usugi, Aiichiro Niwa, and Koichi Tange. Formulation of process improvement knowledge: 7 components of a good PI story. Journal of Software: Evolution and Process, 31(1): e2107:1–e2107:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Noureldin:2022:BAM

[NGAEE22] Menna Noureldin, Atef Ghalwash, Laila Abd-Ellatif, and Mohamed El-Gazzar. Blending agile methodologies to support automotive SPICE compliance. Journal of Software: Evolution and Process, 34 (1):e2391:1–e2391:??, January 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nasr:2013:RSM

[NGvD13] Khalid Adam Nasr, Hans-Gerhard Gross, and Arie van Deursen. Realizing service migration in industry — lessons learned. Journal of Software: Evolution and Process, 25(6): 639–661, June 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nadi:2014:LKC

[NH14] Sarah Nadi and Ric Holt. The Linux kernel: a case study of build system variability. Journal of Software: Evolution and Process, 26(8):730–746, August 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nayrolles:2017:BRA

[NHLTL17] Mathieu Nayrolles, Abdelwahab Hamou-Lhadj, Sofiène Tahar, and Alf Larsson. A bug reproduction approach based on directed model checking and crash traces. Journal of Software: Evolution and Process, 29(3):??, March 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Niazi:2012:ESS

[Nia12] Mahmood Niazi. An exploratory study of software process improvement implementation risks. Journal of Software: Evolution and Process, 24(8): 877–894, December 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Niazi:2015:CSS

[Nia15] Mahmood Niazi. A comparative study of software process improvement implementation success factors. Journal of Software: Evolution and Process, 27 (9):700–722, September 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ngo:2018:FVP

[NL18] Van Chan Ngo and Axel Legay. Formal verification of probabilistic SystemC models with statistical model checking. Journal of Software: Evolution and Process, 30(3): e1890:1–e1890:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Naeem:2020:MLA

[NLNL20] Muhammad Rashid Naeem, Tao Lin, Hamad Naeem, and Hailu Liu. A machine learning approach for classification of equivalent mutants. Journal of Software: Evolution and Process, 32 (5):e2238:1–e2238:??, May 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Niazi:2017:MAS

[NMA^{+}17] Mahmood Niazi, Sajjad Mahmood, Mohammad Alshayeb, Abdulrahman Ahmed Bobakr Baqais, and Asif Qumer Gill. Motivators for adopting social computing in global software development: An empirical study. Journal of Software: Evolution and Process, 29(8):??, August 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Niazi:2020:GGP

[NMA20] Mahmood Niazi, Sajjad Mahmood, and Mohammad Alshayeb. GLOB: a global project management readiness framework. Journal of Software: Evolution and Process, 32(12): e2302:1–e2302:??, December 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Niinimaki:2012:RCU

[NPLP12] Tuomas Niinimäki, Arttu Piri, Casper Lassenius, and Maria Paasivaara. Reflecting the choice and usage of communication tools in global software development projects with media synchronicity theory. Journal of Software: Evolution and Process, 24(6):677–692, October 2012. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Neumann:2013:ISI

[NRB13] Martin Neumann, Andreas Riel, and Daniel Brissaud. IT-supported innovation management in the automotive supplier industry to drive idea generation and leverage innovation. Journal of Software: Evolution and Process, 25(4):329–339, April 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nurcan:2012:ESI

[NSBR12] Selmin Nurcan, Rainer Schmidt, Ilia Bider, and Gil Regev. Editorials: Special issue on business process lifecycle: design, deployment, operation and evaluation. Journal of Software: Evolution and Process, 24(3):231–236, April 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nicolay:2017:PAJ

[NSDD17] Jens Nicolay, Quentin Stiévenart, Wolfgang De Meuter, and Coen De Roover. Purity analysis for JavaScript through abstract interpretation. Journal of Software: Evolution and Process, 29(12): ??, December 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nicoletti:2015:OBT

[NSDP15] Matias Nicoletti, Silvia Schiaffino, and J. Andres Diaz-Pace. An optimization-based tool to support the cost-effective production of software architecture documentation. Journal of Software: Evolution and Process, 27(9): 674–699, September 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Neamtiu:2013:TBU

[NXC13] Iulian Neamtiu, Guowu Xie, and Jianbo Chen. Towards a better understanding of software evolution: an empirical study on open-source software. Journal of Software: Evolution and Process, 25(3): 193–218, March 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Olsson:2017:AHS

[OB17] Helena Holmström Olsson and Jan Bosch. From ad hoc to strategic ecosystem management: the “Three-Layer Ecosystem Strategy Model” (TeLESM). Journal of Software: Evolution and Process, 29(7):??, July 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Olsson:2020:GDD

[OB20] Helena Holmström Olsson and Jan Bosch. Going digital: Disruption and transformation in softwareintensive embedded systems ecosystems. Journal of Software: Evolution and Process, 32 (6):e2249:1–e2249:??, June 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ortega-Cabezas:2020:ARB

[OCCSBDBP20] Pedro Miguel Ortega-Cabezas, Antonio Colmenar-Santos, David Borge-Diez, and Jorge Juan Blanes-Peiró. Application of rulebased expert systems in hardware-in-the-loop simulation case study: Software and performance validation of an engine electronic control unit. Journal of Software: Evolution and Process, 32(1): e2223:1–e2223:??, January 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

OConnor:2017:CSE

[OEC17] Rory V. O’Connor, Peter Elger, and Paul M. Clarke. Continuous software engineering — a microservices architecture perspective. Journal of Software: Evolution and Process, 29(11):??, November 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ouhbi:2018:SRC

[OFAT^{+}18] Sofia Ouhbi, José Luis Fernández-Alemán, Ambrosio Toval, José Rivera Pozo, and Ali Idri. Sustainability requirements for connected health applications. Journal of Software: Evolution and Process, 30(7):e1922:1– e1922:??, July 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ogasawara:2014:PPS

[OKA14] Hideto Ogasawara, Takumi Kusanagi, and Minoru Aizawa. Proposal and practice of software process improvement framework — Toshiba’s software process improvement history since 2000. Journal of Software: Evolution and Process, 26(5): 521–529, May 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ouni:2017:MMO

[OKÓ^{+}17] Ali Ouni, Marouane Kessentini, Mel Ó Cinnéide, Houari Sahraoui, Kalyanmoy Deb, and Katsuro Inoue. MORE: A multiobjective refactoring recommendation approach to introducing design patterns and fixing code smells. Journal of Software: Evolution and Process, 29(5):??, May 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

OLeary:2012:APM

[OMTR12] Pádraig O’Leary, Fergal McCaffery, Steffen Thiel, and Ita Richardson. An agile process model for product derivation in software product line engineering. Journal of Software: Evolution and Process, 24(5):561–571, August 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Oliveto:2014:SDP

[OP14] Rocco Oliveto and Denys Poshyvanyk. A selection of distinguished papers from the 19th Working Conference on Reverse Engineering 2012. Journal of Software: Evolution and Process, 26(11):929– 930, November 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Osterweil:2014:LAA

[Ost14] Leon J. Osterweil. Learning about application software by studying processes. Journal of Software: Evolution and Process, 26(10):869–876, October 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ozcan-Top:2018:HAA

[ÖTM18] Özden Özcan-Top and Fergal McCaffery. A hybrid assessment approach for medical device software development companies. Journal of Software: Evolution and Process, 30 (7):e1929:1–e1929:??, July 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Papatheocharous:2014:EES

[PA14] Efi Papatheocharous and Andreas S. Andreou. Empirical evidence and state of practice of software agile teams. Journal of Software: Evolution and Process, 26(9):855–866, September 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Perez:2016:FPC

[PA16] Alexandre Perez and Rui Abreu. Framing program comprehension as fault localization. Journal of Software: Evolution and Process, 28(10):840–862, October 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pracner:2017:ECT

[PB17] Doni Pracner and Zoran Budimac. Enabling code transformations with FermaT on simplified bytecode. Journal of Software: Evolution and Process, 29 (5):??, May 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Prause:2016:SPA

[PBDJ16] Christian R. Prause, Markus Bibus, Carsten Dietrich, and Wolfgang Jobi. Software product assurance at the German Space Agency. Journal of Software: Evolution and Process, 28(9): 744–761, September 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Papatheocharous:2017:IED

[PBSA17] Efi Papatheocharous, Stamatia Bibi, Ioannis Stamelos, and Andreas S. Andreou. An investigation of effort distribution among development phases: A four-stage progressive software cost estimation model. Journal of Software: Evolution and Process, 29(10):??, October 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Perez-Castillo:2014:OBS

[PCCP14] Ricardo Pérez-Castillo, Danilo Caivano, and Mario Piattini. Ontology-based similarity applied to business process clustering. Journal of Software: Evolution and Process, 26 (12):1128–1149, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Perez-Castillo:2021:ARE

[PCCRP21] Ricardo Pérez-Castillo, Danilo Caivano, Francisco Ruiz, and Mario Piattini. ArchiRev-Reverse engineering of information systems toward Archi-Mate models. An industrial case study. Journal of Software: Evolution and Process, 33(2): e2314:1–e2314:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Perez-Castillo:2013:SMR

[PCdGCP13] Ricardo Pérez-Castillo, Ignacio García-Rodríguez de Guzmán, Ismael Caballero, and Mario Piattini. Software modernization by recovering Web services from legacy databases. Journal of Software: Evolution and Process, 25(5):507–533, May 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Paul:2021:NDL

[PCL21] T. K. Paul, M. J. M.

Chowdhury, and M. F. Lau. A new disjunctive literal insertion fault detection strategy in boolean specifications. Journal of Software: Evolution and Process, 33(5):e2336:1– e2336:??, May 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pomar:2014:USV

[PCMCAC14] Filipe Albero Pomar, Jose A. Calvo-Manzano, Edgar Caballero, and Magdalena Arcilla-Cobián. Understanding sprint velocity fluctuations for improved project plans with Scrum: a case study. Journal of Software: Evolution and Process, 26(9):776– 783, September 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Perez-Castillo:2018:ESH

[PCP18] Ricardo Perez-Castillo and Mario Piattini. An empirical study on how project context impacts on code cloning. Journal of Software: Evolution and Process, 30(12): e2115:1–e2115:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Plaza:2018:SAH

[PDP18] Andrea M. Plaza, Jessica Díaz, and Jennifer Pérez. Software architectures for health care cyber-physical systems: a systematic literature review. Journal of Software: Evolution and Process, 30 (7):e1930:1–e1930:??, July 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Parnin:2018:SEJ

[PF18] Christopher Parnin and Johan Fabry. Special edition of the Journal of Software: Evolution and Process for thef 4th IEEE Working Conference on Software Visualization (VISSOFT 2016). Journal of Software: Evolution and Process, 30(2): e1940:1–e1940:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Perkusich:2017:ACI

[PGAP17] Mirko Perkusich, Kyller Costa Gorgônio, Hyggo Almeida, and Angelo Perkusich. Assisting the continuous improvement of Scrum projects using metrics and Bayesian networks. Journal of Software: Evolution and Process, 29(6):??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pacheco:2015:PMR

[PGCMA15] C. L. Pacheco, I. A. Garcia, J. A. Calvo-Manzano, and M. Arcilla. A proposed model for reuse of software requirements in requirements catalog. Journal of Software: Evolution and Process, 27(1): 1–21, January 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ponsard:2021:GDA

[PGM21] Christophe Ponsard, Jeremy Grandclaudon, and Philippe Massonet. A goal-driven approach for the joint deployment of safety and security standards for operators of essential services. Journal of Software: Evolution and Process, 33(9): e2338:1–e2338:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pries-Heje:2015:CCS

[PHJ15] Jan Pries-Heje and Jørn Johansen. Choosing change strategy for ISO/ IEC 33014. Journal of Software: Evolution and Process, 27(8):573–583, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Polancic:2016:IRT

[PJ16] Gregor Polančič and Gregor Jošt. The impact of the representatives of three types of process modeling tools on modeler’s perceptions and performance. Journal of Software: Evolution and Process, 28(1): 27–56, January 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pandita:2017:TDR

[PJS^{+}17] Rahul Pandita, Raoul Jetley, Sithu Sudarsan, Timothy Menzies, and Laurie Williams. TMAP: Discovering relevant API methods through text mining of API documentation. Journal of Software: Evolution and Process, 29(12): ??, December 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pahl:2017:CAC

[PJW17] Claus Pahl, Pooyan Jamshidi, and Danny Weyns. Cloud architecture continuity: Change models and change rules for sustainable cloud software architectures. Journal of Software: Evolution and Process, 29(2): ??, February 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Polgar:2014:PCM

[PK14] Péter Balázs Polgár and Ferenc Kazinci. Process communication in medical software development. Journal of Software: Evolution and Process, 26(3): 313–320, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pfahl:2016:ISI

[PKBT16] Dietmar Pfahl, Marco Kuhrmann, Reda Bendraou, and Richard Turner. ICSSP 2015 — special issue introduction. Journal of Software: Evolution and Process, 28(9):716– 721, September 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Park:2014:MSA

[PKY14] Joonseok Park, Taejun Kang, and Keunhyuk Yeom. Mobile situationaware framework for developing smart mobile software. Journal of Software: Evolution and Process, 26 (12):1213–1232, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Paasivaara:2014:ACG

[PL14] Maria Paasivaara and Casper Lassenius. Agile coaching for global software development. Journal of Software: Evolution and Process, 26(4): 404–418, April 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pagliari:2020:ECP

[PMT20] Lorenzo Pagliari, Raffaela Mirandola, and Catia Trubiani. Engineering cyberphysical systems through performance-based modelling and analysis: a case study experience report. Journal of Software: Evolution and Process, 32(1): e2179:1–e2179:??, January 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Piri:2012:FDG

[PNL12] Arttu Piri, Tuomas Niinimäki, and Casper Lassenius. Fear and distrust in global software engineering projects. Journal of Software: Evolution and Process, 24(2): 185–205, March 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Poth:2016:EEE

[Pot16]  Alexander Poth.  EU-

ROSPI 2015: Effectivity and economical aspects for agile quality assurance in large enterprises. Journal of Software: Evolution and Process, 28(11):1000– 1004, November 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pinzger:2014:ESI

[PP14] Martin Pinzger and Denys Poshyvanyk. Editorial: Special issue: a selection of distinguished papers from the 18th Working Conference on Reverse Engineering 2011. Journal of Software: Evolution and Process, 26(1):1–2, January 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

PhD:2021:KGM

[PPdAPP21] Maria-Isabel Sanchez-Segura PhD, Fuensanta Medina-Dominguez PhD, Antonio de Amescua PhD, and German-Lenin Dugarte-Peña PhD. Knowledge governance maturity assessment can help software engineers during the design of business digitalization projects. Journal of Software: Evolution and Process, 33(4): e2326:1–e2326:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Polo:2020:AGO

[PPPdG20] Macario Polo, Oscar Pedreira, Ángeles S. Places, and Ignacio García Rodríguez de Guzmán. Automated generation of oracled test cases with regular expressions and combinatorial techniques. Journal of Software: Evolution and Process, 32(12):e2273:1– e2273:??, December 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pino:2012:SMM

[PRGP12] Francisco J. Pino, Francisco Ruiz, Félix García, and Mario Piattini. A software maintenance methodology for small organizations: Agile MANTEMA. Journal of Software: Evolution and Process, 24(8): 851–876, December 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Prikladnicki:2012:PGS

[Pri12] Rafael Prikladnicki. Propinquity in global software engineering: examining perceived distance in globally distributed project teams. Journal of Software: Evolution and Process, 24(2): 119–137, March 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Petersen:2014:EIB

[PRNR14] Kai Petersen, Peter Roos, Staffan Nyström, and Per Runeson. Early identification of bottlenecks in very large scale system of systems software development. Journal of Software: Evolution and Process, 26 (12):1150–1171, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Poth:2019:LAS

[PSMM19] Alexander Poth, Susumu Sasabe, Antònia Mas, and Antoni-Lluís Mesquida. Lean and agile software process improvement in traditional and agile environments. Journal of Software: Evolution and Process, 31(1):e1986:1– e1986:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Pate:2013:CES

[PTK13] Jeremy R. Pate, Robert Tairas, and Nicholas A. Kraft. Clone evolution: a systematic review. Journal of Software: Evolution and Process, 25(3): 261–283, March 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Prenner:2021:GCH

[PUWS21] Nils Prenner, Carolin Unger-Windeler, and Kurt Schneider. Goals and challenges in hybrid software development approaches.

Journal of Software: Evolution and Process, 33(11): e2382:1–e2382:??, November 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Paschali:2021:IGR

[PVA^{+}21] Maria-Eleni Paschali, Christina Volioti, Apostolos Ampatzoglou, Anastasios Gkagkas, Ioannis Stamelos, and Alexander Chatzigeorgiou. Implementing game requirements using design patterns. Journal of Software: Evolution and Process, 33(12): e2399:1–e2399:??, December 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Qi:2021:EEO

[QB21] Kan Qi and Barry Boehm. Effort estimation of open source Android projects via transaction analysis. Journal of Software: Evolution and Process, 33(1): e2253:1–e2253:??, January 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Qusef:2013:ETC

[QBO^{+}13] Abdallah Qusef, Gabriele Bavota, Rocco Oliveto, Andrea De Lucia, and David Binkley. Evaluating test-to-code traceability recovery methods through controlled experiments. Journal of Software: Evolution and Process, 25(11):1167–1191, November 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Qasaimeh:2022:AST

[QHY^{+}22] Malik Qasaimeh, Rand Abu Hammour, Muneer Bani Yassein, Raad S. Al-Qassas, Juan Alfonso Lara Torralbo, and David Lizcano. Advanced security testing using a cyberattack forecasting model: a case study of financial institutions. Journal of Software: Evolution and Process, 34(11):e2489:1– e2489:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Qiu:2015:IFB

[QSM15] Jing Qiu, Xiaohong Su, and Peijun Ma. Identifying functions in binary code with reverse extended control flow graphs. Journal of Software: Evolution and Process, 27(10):793– 820, October 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Quattrocchi:2022:PMI

[QT22] Giovanni Quattrocchi and Damian Andrew Tamburri. Predictive maintenance of infrastructure code using “fluid” datasets: an exploratory study on Ansible defect proneness. Journal of Software: Evolution and Process, 34(11):e2480:1– e2480:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Quinton:2021:EDS

[QVR^{+}21] Clément Quinton, Michael Vierhauser, Rick Rabiser, Luciano Baresi, Paul Grünbacher, and Christian Schuhmayer. Evolution in dynamic software product lines. Journal of Software: Evolution and Process, 33(2): e2293:1–e2293:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rasool:2015:RCS

[RA15] Ghulam Rasool and Zeeshan Arshad. A review of code smell mining techniques. Journal of Software: Evolution and Process, 27(11):867– 895, November 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rajlich:2018:CI

[Raj18a] Václav Rajlich. Cover image. Journal of Software: Evolution and Process, 30 (9):i, September 2018. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rajlich:2018:FRS

[Raj18b]  Václav Rajlich.  Five

recommendations for software evolvability. Journal of Software: Evolution and Process, 30(9): e1949:1–e1949:??, September 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rafi:2022:SDB

[RAM^{+}22] Saima Rafi, Muhammad Azeem Akbar, Sajjad Mahmood, Ahmed Alsanad, and Abdulrahman Alothaim. Selection of DevOps best test practices: a hybrid approach using ISM and fuzzy TOP-SIS analysis. Journal of Software: Evolution and Process, 34(5):e2448:1– e2448:??, May 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Raninen:2013:LLW

[RAS^{+}13] Anu Raninen, Jarmo J.

Ahonen, Hanna-Miina Sihvonen, Paula Savolainen, and Sarah Beecham. LAPPI: A light-weight technique to practical process modeling and improvement target identification. Journal of Software: Evolution and Process, 25(9):915– 933, September 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rico:2021:CSI

[RBE^{+}21] Sergio Rico, Elizabeth Bjarnason, Emelie Engström, Martin Höst, and Per Runeson. A case study of industry-academia communication in a joint software engineering research project. Journal of Software: Evolution and Process, 33(10):e2372:1– e2372:??, October 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Regan:2015:ATP

[RBFM15] Gilbert Regan, Miklos Biro, Derek Flood, and Fergal McCaffery. Assessing traceability — practical experiences and lessons learned. Journal of Software: Evolution and Process, 27(8):591–601, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rizvi:2015:SRD

[RBG15] Buturab Rizvi, Ebrahim Bagheri, and Dragan Gasevic. A systematic review of distributed Agile software engineering. Journal of Software: Evolution and Process, 27(10):723– 762, October 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Roy:2016:EGE

[RBM16] Chanchal K. Roy, Andrew Begel, and Leon Moonen. Editorial: Guest editor’s introduction to the Special Issue on Program Comprehension (ICPC 2014). Journal of Software: Evolution and Process, 28(10): 838–839, October 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rak:2019:EEM

[RCL19] Katija Rak, Zeljka Car, and Ignac Lovrek. Effort estimation model for software development projects based on use case reuse. Journal of Software: Evolution and Process, 31(2): e2119:1–e2119:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rashid:2020:MEP

[RCO20] Mehvish Rashid, Paul M.

Clarke, and Rory V. O’Connor. A mechanism to explore proactive knowledge retention in open source software communities. Journal of Software: Evolution and Process, 32(3):e2198:1– e2198:??, March 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Riel:2012:PPI

[RDD^{+}12] Andreas Riel, Anca Draghici, George Draghici, Damian Grajewski, and Richard Messnarz. Process and product innovation needs integrated engineering collaboration skills. Journal of Software: Evolution and Process, 24(5):551– 560, August 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Raza:2016:MAP

[RF16] M. Raza and J. P. Faria. A model for analyzing performance problems and root causes in the personal software process. Journal of Software: Evolution and Process, 28(4): 254–271, April 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rust:2016:ECI

[RFM16] Peter Rust, Derek Flood, and Fergal McCaffery. EU-ROSPI 2015: Creation of an IEC 62304 compliant software development plan. Journal of Software: Evolution and Process, 28 (11):1005–1010, November 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rahanu:2015:TRD

[RGS^{+}15] Harjinder Rahanu, Elli Georgiadou, Kerstin Siakas, Damjan Ekert, Richard Messanarz, and Geetha Abeysinghe. Towards relating delivery methods and examination success: lessons learned from the VALO LLP project case study. Journal of Software: Evolution and Process, 27(8):555–564, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rahanu:2020:HSW

[RGSR20] Harjinder Rahanu, Elli Georgiadou, Kerstin Siakas, and Margaret Ross. Health, safety, and well-being through ethical behaviours in computer and information systems development and deployment. Journal of Software: Evolution and Process, 32(3): e2225:1–e2225:??, March 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rodeghero:2016:ESH

[RHD^{+}16] Paige Rodeghero, Da Huo, Tao Ding, Collin McMillan, and Malcom Gethers. An empirical study on how expert knowledge affects bug reports. Journal of Software: Evolution and Process, 28(7): 542–564, July 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ruohonen:2015:TST

[RHL15] Jukka Ruohonen, Sami Hyrynsalmi, and Ville Leppänen. Time series trends in software evolution. Journal of Software: Evolution and Process, 27 (12):990–1015, December 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rashid:2018:APG

[RK18a] Nasir Rashid and Siffat Ullah Khan. Agile practices for global software development vendors in the development of green and sustainable software. Journal of Software: Evolution and Process, 30(10): e1964:1–e1964:??, October 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rashid:2018:UAM

[RK18b] Nasir Rashid and Siffat Ullah Khan. Using agile methods for the development of green and sustainable software: Success factors for GSD vendors. Journal of Software: Evolution and Process, 30(8): e1927:1–e1927:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Razavian:2014:LMS

[RL14] Maryam Razavian and Patricia Lago. A lean and mean strategy: a data migration industrial study. Journal of Software: Evolution and Process, 26(2): 141–171, February 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Razavian:2015:SLR

[RL15] Maryam Razavian and Patricia Lago. A systematic literature review on SOA migration. Journal of Software: Evolution and Process, 27(5):337–372, May 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rajaram:2019:TBT

[RLMC19] Hamsini Ketheswarasarma Rajaram, John Loane, Silvana Togneri MacMahon, and Fergal Mc Caffery. Taxonomy-based testing and validation of a new defect classification for health software. Journal of Software: Evolution and Process, 31(1): e1985:1–e1985:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Reggio:2021:SOM

[RLR21] Gianna Reggio, Maurizio Leotta, and Filippo Ricca. A service-oriented method for domain and business process modelling. Journal of Software: Evolution and Process, 33(2): e2307:1–e2307:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Robillard:2014:TST

[RLTTC14] Pierre N. Robillard, Mathieu Lavallée, Yvan Ton-That, and François Chiocchio. Taxonomy for software teamwork measurement. Journal of Software: Evolution and Process, 26(10):910–922, October 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rucker:2022:FAL

[RM22] Nadine Rücker and Andreas Maier. FlexParser — the adaptive log file parser for continuous results in a changing world. Journal of Software: Evolution and Process, 34(3): e2426:1–e2426:??, March 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Raninen:2015:ASP

[RMAB15] Anu Raninen, Helena Merikoski, Jarmo J. Ahonen, and Sarah Beecham.

Applying software process modeling to improve customer support processes. Journal of Software: Evolution and Process, 27(4): 274–293, April 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ruano-Mayoral:2014:MFA

[RMCLGAM14] Marcos Ruano-Mayoral, Cristina Casado-Lumbreras, Helena Garbarino-Alberti, and Sanjay Misra. Methodological framework for the allocation of work packages in global software development. Journal of Software: Evolution and Process, 26(5):476–487, May 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ramirez-Mora:2020:GMT

[RMOP20] Sandra L. Ramírez-Mora, Hanna Oktaba, and Juana Patlán Pérez. Group maturity, team efficiency, and team effectiveness in software development: a case study in a CMMI-DEV Level 5 organization. Journal of Software: Evolution and Process, 32(4): e2232:1–e2232:??, April 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Regan:2020:QIM

[RMP^{+}20] Gilbert Regan, Fergal Mc-Caffery, Pangkaj Chandra Paul, Jan Reich, Eric Armengaud, Cem Kaypmaz, Marc Zeller, Joe Zhensheng Guo, Simone Longo, Eoin O’Carroll, and Ioannis Sorokos. Quality improvement mechanism for cyber physical systems — an evaluation. Journal of Software: Evolution and Process, 32(11):e2295:1– e2295:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Richardson:2012:ERD

[RPV12] Ita Richardson, Dan Paulish, and June Verner. Editorials: Research directions for global software engineering — where to next? Journal of Software: Evolution and Process, 24(2):115–117, March 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Raffo:2014:EIS

[RPZ14] David Raffo, Dietmar Pfahl, and Li Zhang. Editorial: ICSSP 2011 special issue: processes for tomorrow’s systems and software engineering: an evolving dynamic domain. Journal of Software: Evolution and Process, 26(4): 369–370, April 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ruiz-Rube:2013:UAS

[RRDPD^{+}13] Iván Ruiz-Rube, Juan Manuel Dodero, Manuel Palomo-Duarte, Mercedes Ruiz, and David Gawn. Uses and applications of Software & Systems Process Engineering Meta-Model process models. A systematic mapping study. Journal of Software: Evolution and Process, 25(9):999–1025, September 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rouhi:2022:FMT

[RRL22] Alireza Rouhi, Shekoufeh Kolahdouz Rahimi, and Kevin Lano. Formalizing model transformation patterns. Journal of Software: Evolution and Process, 34(2): e2406:1–e2406:??, February 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Razavian:2016:TMH

[RTCL16] Maryam Razavian, Antony Tang, Rafael Capilla, and Patricia Lago. In two minds: how reflections influence software design thinking. Journal of Software: Evolution and Process, 28(6):394–426, June 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ryan:2016:SPC

[Rya16] Kevin T. Ryan. Software processes for a changing world. Journal of Software: Evolution and Process, 28(4):236–240, April 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Rafi:2021:RMD

[RYA^{+}21] Saima Rafi, Wu Yu, Muhammad Azeem Akbar, Sajjad Mahmood, Ahmed Alsanad, and Abdu Gumaei. Readiness model for DevOps implementation in software organizations. Journal of Software: Evolution and Process, 33(4): e2323:1–e2323:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Raschke:2015:WDA

[RZL^{+}15] Wolfgang Raschke, Massimiliano Zilli, Johannes Loinig, Reinhold Weiss, Christian Steger, and Christian Kreiner. Where does all this waste come from? Journal of Software: Evolution and Process, 27(8):584–590, August 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Serebrenik:2017:FSS

[SA17] Alexander Serebrenik and Bram Adams. Foreword to the SANER 2015 special issue. Journal of Software: Evolution and Process, 29 (3):??, March 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sultana:2021:USM

[SAC21] Kazi Zakia Sultana, Vaibhav Anu, and Tai-Yin Chong. Using software metrics for predicting vulnerable classes and methods in Java projects: a machine learning approach. Journal of Software: Evolution and Process, 33(3): e2303:1–e2303:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Schoofs:2022:ATA

[SAD22] Ebert Schoofs, Mehrdad Abdi, and Serge Demeyer. AmPyfier: Test amplification in Python. Journal of Software: Evolution and Process, 34(11): e2490:1–e2490:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shawky:2016:CSD

[SAEH16] Doaa M. Shawky and Salwa K. Abd-El-Hafiz. Characterizing software development method using metrics. Journal of Software: Evolution and Process, 28(2):82–96, February 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sisman:2017:ESC

[SAK17] Bunyamin Sisman, Shayan A. Akbar, and Avinash C. Kak. Exploiting spatial code proximity and order for improved source code retrieval for bug localization. Journal of Software: Evolution and Process, 29 (1):??, January 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Seifzadeh:2013:SDS

[SAM13] Habib Seifzadeh, Hassan Abolhassani, and Mohsen Sadighi Moshkenani. A survey of dynamic software updating. Journal of Software: Evolution and Process, 25(5): 535–568, May 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Santos:2014:GCT

[San14] André L. Santos. GUI code tracing through direct program interaction. Journal of Software: Evolution and Process, 26(6): 552–568, June 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sas:2022:RBA

[SAPF22] Darius Sas, Paris Avgeriou, Ilaria Pigazzini, and Francesca Arcelli Fontana. On the relation between architectural smells and source code changes. Journal of Software: Evolution and Process, 34(1): e2398:1–e2398:??, January 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Steiner:2012:MTP

[SBPS12] Michael Steiner, Monique Blaschke, Michael Philipp, and Tomas Schweigert. Make test process assessment similar to software process assessment — the Test SPICE approach. Journal of Software: Evolution and Process, 24(5):471–480, August 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sajedi-Badashian:2020:GEB

[SBS20] Ali Sajedi-Badashian and Eleni Stroulia. Guidelines for evaluating bugassignment research. Journal of Software: Evolution and Process, 32(9): e2250:1–e2250:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Schmidt:2012:MSS

[Sch12] Rainer Schmidt. Metaservice support for the business process life cycle. Journal of Software: Evolution and Process, 24(3): 361–373, April 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Senapathi:2021:STA

[SDG21] Mali Senapathi and Meghann L. Drury-Grogan. Systems thinking approach to implementing Kanban: a case study. Journal of Software: Evolution and Process, 33(4):e2322:1– e2322:??, April 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Saeeda:2020:PFI

[SDWA20] Hina Saeeda, Junyu Dong, Yong Wang, and Muhammmad Adil Abid. A proposed framework for improved software requirements elicitation process in SCRUM: Implementation by a real-life Norwaybased IT project. Journal of Software: Evolution and Process, 32 (7):e2247:1–e2247:??, July 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Svahnberg:2017:MAR

[SG17] Mikael Svahnberg and Tony Gorschek. A model for assessing and reassessing the value of software reuse. Journal of Software: Evolution and Process, 29(4):??, April 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Singh:2020:MMI

[SG20] Ashima Singh and Sukhpal Singh Gill. Measuring the maturity of Indian small and medium enterprises for unofficial readiness for capability maturity model integrationbased software process improvement. Journal of Software: Evolution and Process, 32(9):e2261:1– e2261:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shah:2014:THT

[SGAP14] Syed Muhammad Ali Shah, Cigdem Gencel, Usman Sattar Alvi, and Kai Petersen. Towards a hybrid testing process unifying exploratory testing and scripted testing. Journal of Software: Evolution and Process, 26(2):220– 250, February 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sun:2013:AMN

[SGD^{+}13] Yu Sun, Jeff Gray, Romain Delamare, Benoit Baudry, and Jules White. Automating the maintenance of nonfunctional system properties using demonstration-based model transformation. Journal of Software: Evolution and Process, 25(12):1335– 1356, December 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sanchez-Gordon:2019:IAD

[SGSGYO19] Sandra Sanchez-Gordon, Mary Sánchez-Gordón, Murat Yilmaz, and Rory V. O’Connor. Integration of accessibility design patterns with the software implementation process of ISO/IEC 29110. Journal of Software: Evolution and Process, 31(1): e1987:1–e1987:??, January 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Szabo:2023:UCA

[SH23] Bálint Szabó and Károly Hercegfi. User-centered approaches in software development processes: Qualitative research into the practice of Hungarian companies. Journal of Software: Evolution and Process, 35(2): e2501:1–e2501:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shatnawi:2015:DMT

[Sha15] Raed Shatnawi. Deriving metrics thresholds using log transformation. Journal of Software: Evolution and Process, 27(2):95–113, February 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sirkia:2018:JKC

[Sir18] Teemu Sirkiä. Jsvee and Kelmu: Creating and tailoring program animations for computing education. Journal of Software: Evolution and Process, 30(2): e1924:1–e1924:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shang:2014:ESE

[SJA^{+}14] Weiyi Shang, Zhen Ming Jiang, Bram Adams, Ahmed E. Hassan, Michael W. Godfrey, Mohamed Nasser, and Parminder Flora. An exploratory study of the evolution of communicated information about the execution of large software systems. Journal of Software: Evolution and Process, 26(1):3–26, January 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Schrettner:2014:IAP

[SJG^{+}14] Lajos Schrettner, Judit Jász, Tamás Gergely, Árpád Beszédes, and Tibor Gyimóthy. Impact analysis in the presence of dependence clusters using Static Execute After in WebKit. Journal of Software: Evolution and Process, 26(6):569–588, June 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Salam:2018:CDG

[SK18] Muhammad Salam and Siffat Ullah Khan. Challenges in the development of green and sustainable software for software multisourcing vendors: Findings from a systematic literature review and industrial survey. Journal of Software: Evolution and Process, 30(8): e1939:1–e1939:??, August 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shameem:2018:PCA

[SKK^{+}18] Mohammad Shameem, Rakesh Ranjan Kumar, Chiranjeev Kumar, Bibhas Chandra, and Arif Ali Khan. Prioritizing challenges of agile process in distributed software development environment using analytic hierarchy process. Journal of Software: Evolution and Process, 30(11): e1979:1–e1979:??, November 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Svajlenko:2015:BDC

[SKR15] Jeffrey Svajlenko, Iman Keivanloo, and Chanchal K. Roy. Big data clone detection using classical detectors: an exploratory study. Journal of Software: Evolution and Process, 27(6):430–464, June 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Samalikova:2014:PMS

[SKTW14] J. Samalikova, R. J.

Kusters, J. J. M. Trienekens, and A. J. M. M. Weijters. Process mining support for Capability Maturity Model integration-based software process assessment, in principle and in practice. Journal of Software: Evolution and Process, 26(7):714–728, July 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sahin:2015:MTT

[SKWD15] Dilan Sahin, Marouane Kessentini, Manuel Wimmer, and Kalyanmoy Deb. Model transformation testing: a bi-level search-based software engineering approach. Journal of Software: Evolution and Process, 27(11):821– 837, November 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Saputri:2020:ASR

[SL20] Theresia Ratih Dewi Saputri and Seok-Won Lee. Addressing sustainability in the requirements engineering process: From elicitation to functional decomposition. Journal of Software: Evolution and Process, 32(8): e2254:1–e2254:??, August 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sae-Lim:2018:CBA

[SLHS18] Natthawute Sae-Lim, Shinpei Hayashi, and Motoshi Saeki. Context-based approach to prioritize code smells for prefactoring. Journal of Software: Evolution and Process, 30 (6):e1886:1–e1886:??, June 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shi:2021:MEU

[SLL^{+}21] Ke Shi, Yang Lu, Guangliang Liu, Zhenchun Wei, and Jingfei Chang. MPTembedding: an unsupervised representation learning of code for software defect prediction. Journal of Software: Evolution and Process, 33(4): e2330:1–e2330:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Singh:2021:FAR

[SLMW21] Neeraj Kumar Singh, Mark Lawford, Thomas S. E. Maibaum, and Alan Wassyng. A formal approach to rigorous development of critical systems. Journal of Software: Evolution and Process, 33(4): e2334:1–e2334:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Scalabrino:2018:CMC

[SLVOP18] Simone Scalabrino, Mario Linares-Vásquez, Rocco Oliveto, and Denys Poshyvanyk. A comprehensive model for code readability. Journal of Software: Evolution and Process, 30 (6):e1958:1–e1958:??, June 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Schweigert:2014:TIT

[SN14] Tomas Schweigert and Andreas Nehfort. Technical issues in test process assessment and their current handling in TestSPICE.

Journal of Software: Evolution and Process, 26(3): 350–356, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shailesh:2022:TSD

[SNP22] Tanuja Shailesh, Ashalatha Nayak, and Devi Prasad. Transformation of sequence diagram to timed Petri net using Atlas Transformation Language metamodel approach. Journal of Software: Evolution and Process, 34(1):e2412:1– e2412:??, January 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Santos:2022:BEK

[SP22a] José Amancio M. Santos and Gadiel Xavier Antunes Petronilo. Building empirical knowledge on the relationship between code smells and design patterns: an exploratory study. Journal of Software: Evolution and Process, 34(9): e2487:1–e2487:??, September 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sharma:2022:APL

[SP22b] Shachi Sharma and Parag C. Pendharkar. On the analysis of power law distribution in software component sizes. Journal of Software: Evolution and Process, 34(2):e2417:1– e2417:??, February 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Seo:2016:RFS

[SPC16] Jooyoung Seo, Jihyun Park, and Byoungju Choi. A runtime fault survival method for deployed software during production runs. Journal of Software: Evolution and Process, 28 (2):97–119, February 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Stehle:2019:PMC

[SR19] Tilmann Stehle and Matthias Riebisch. A porting method for coordinated multiplatform evolution. Journal of Software: Evolution and Process, 31(2): e2116:1–e2116:??, February 2019. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Saldana-Ramos:2014:SAW

[SRSEGA14] Javier Saldaña-Ramos, Ana Sanz-Esteban, Javier García, and Antonio Amescua. Skills and abilities for working in a global software development team: a competence model. Journal of Software: Evolution and Process, 26(3):329–338, March 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sharma:2018:FEE

[SS18] Pooja Sharma and Amrit Lal Sangal. Framework for empirical examination and modeling structural dependencies among inhibitors that impact SPI implementation initiatives in software SMEs. Journal of Software: Evolution and Process, 30(12): e1993:1–e1993:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shah:2015:LQI

[SSA15] Faisal Tehseen Shah, Shafay Shamail, and Niaz Ahmad Akhtar. Lean quality improvement model for quality practices in software industry in Pakistan. Journal of Software: Evolution and Process, 27(4):237–254, April 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Stolfa:2020:DEB

[SSB^{+}20] Jakub Stolfa, Svatopluk Stolfa, Christian Baio, Utimia Madaleno, Petr Dolejsi, Federico Brugnoli, and Richard Messnarz. DRIVES-EU blueprint project for the automotive sector — a literature review of drivers of change in automotive industry. Journal of Software: Evolution and Process, 32(3): e2222:1–e2222:??, March 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sajnani:2015:PEA

[SSL15] Hitesh Sajnani, Vaibhav Saini, and Cristina Lopes. A parallel and efficient approach to large scale clone detection. Journal of Software: Evolution and Process, 27(6):402–429, June 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sablis:2021:TEC

[SSM21] Aivars Sablis, Darja Smite, and Nils Moe. Teamexternal coordination in large-scale software development projects. Journal of Software: Evolution and Process, 33(3): e2297:1–e2297:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Stocco:2022:CDW

[ST22] Andrea Stocco and Paolo Tonella. Confidence-driven weighted retraining for predicting safety-critical failures in autonomous driving systems. Journal of Software: Evolution and Process, 34(10): e2386:1–e2386:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Saarikallio:2023:QCB

[ST23] Matti Saarikallio and Pasi Tyrväinen. Quality culture boosts agile transformation-Action research in a business-to-business software business. Journal of Software: Evolution and Process, 35(1): e2504:1–e2504:??, January 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sturm:2012:SBP

[Stu12] Arnon Sturm. Supporting business process analysis via data warehousing. Journal of Software: Evolution and Process, 24(3): 303–319, April 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sneed:2020:CDS

[SV20] Harry M. Sneed and Chris Verhoef. Cost-driven software migration: an experience report. Journal of Software: Evolution and Process, 32 (7):e2236:1–e2236:??, July 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shafagatova:2021:DTP

[SV21] Aygun Shafagatova and Amy Van Looy. Developing a tool for processoriented appraisals and rewards: Design science research. Journal of Software: Evolution and Process, 33(3):e2321:1– e2321:??, March 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Silva:2017:ICL

[SVB^{+}17] Leonardo Humberto Silva, Marco Tulio Valente, Alexandre Bergel, Nicolas Anquetil, and Anne Etien. Identifying classes in legacy JavaScript code. Journal of Software: Evolution and Process, 29(8): ??, August 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Schweigert:2014:AMM

[SVK^{+}14] Tomas Schweigert, Detlef Vohwinkel, Morten Korsaa, Risto Nevalainen, and Miklos Biro. Agile maturity model: analysing agile maturity characteristics from the SPICE perspective. Journal of Software: Evolution and Process, 26(5):513–520, May 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Saini:2020:IDI

[SVSC20] Munish Saini, Rohan Verma, Antarpuneet Singh, and Kuljit Kaur Chahal. Investigating diversity and impact of the popularity metrics for ranking software packages. Journal of Software: Evolution and Process, 32(9): e2265:1–e2265:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Smite:2012:LLT

[ŠW12] Darja Šmite and Claes Wohlin. Lessons learned from transferring software products to India. Journal of Software: Evolution and Process, 24(6):605– 623, October 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Sahin:2016:HDC

[SWT^{+}16] Cagri Sahin, Mian Wan, Philip Tornquist, Ryan McKenna, Zachary Pearson, William G. J. Halfond, and James Clause. How does code obfuscation impact energy usage? Journal of Software: Evolution and Process, 28(7): 565–588, July 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Shahzad:2012:PWP

[SZ12] Khurram Shahzad and Jelena Zdravkovic. Process warehouses in practice: a goal-driven method for business process analysis. Journal of Software: Evolution and Process, 24(3):321–339, April 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tufano:2017:ESD

[TBP^{+}17] Michele Tufano, Gabriele Bavota, Denys Poshyvanyk, Massimiliano Di Penta, Rocco Oliveto, and Andrea De Lucia. An empirical study on developerrelated factors characterizing fix-inducing commits. Journal of Software: Evolution and Process, 29(1): ??, January 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tan:2021:ETD

[TFAL21] Jie Tan, Daniel Feitosa, Paris Avgeriou, and Mircea Lungu. Evolution of technical debt remediation in Python: a case study on the Apache Software Ecosystem. Journal of Software: Evolution and Process, 33(4):e2319:1– e2319:??, April 2021. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Teyton:2014:SLM

[TFPB14] Cédric Teyton, Jean-Rémy Falleri, Marc Palyart, and Xavier Blanc. A study of library migrations in Java. Journal of Software: Evolution and Process, 26 (11):1030–1052, November 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tiwari:2020:UCS

[TG20] Saurabh Tiwari and Atul Gupta. Use case specifications: How complete are they? Journal of Software: Evolution and Process, 32 (1):e2218:1–e2218:??, January 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tanveer:2017:EEA

[TGE17] Binish Tanveer, Liliana Guzmán, and Ulf Martin Engel. Effort estimation in agile software development: Case study and improvement framework. Journal of Software: Evolution and Process, 29(11):??, November 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tunkelo:2013:IGD

[THP13] Teemu Tunkelo, Ari-Pekka Hameri, and Yves Pigneur. Improving globally distributed software development and support processes — a workflow view. Journal of Software: Evolution and Process, 25 (12):1305–1324, December 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tran:2016:IRT

[THP^{+}16] Hanh Nhi Tran, Mojtaba Hajmoosaei, Christian Percebois, Agnes Front, and Claudia Roncancio. Integrating runtime changes into system and software process enactment. Journal of Software: Evolution and Process, 28(9):762– 782, September 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tell:2021:TSC

[TKK^{+}21] Paolo Tell, Jil Klünder, Steffen Küpper, David Raffo, Stephen MacDonell, Jürgen Münch, Dietmar Pfahl, Oliver Linssen, and Marco Kuhrmann. Towards the statistical construction of hybrid development methods. Journal of Software: Evolution and Process, 33(1): e2315:1–e2315:??, January 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Takeuchi:2014:RAE

[TKS^{+}14] Motoko Takeuchi, Naohiko Kohtake, Seiko Shirasaka, Yumi Koishi, and Kazunori Shioya. Report on an assessment experience based on ISO/IEC 29110. Journal of Software: Evolution and Process, 26(3):306–312, March 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Trols:2021:ESC

[TMDE21] Michael Tröls, Atif Mashkoor, Andreas Demuth, and Alexander Egyed. Ensuring safe and consistent coengineering of cyberphysical production systems: a case study. Journal of Software: Evolution and Process, 33(9): e2308:1–e2308:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Teruel:2017:CFM

[TNLJ^{+}17] Miguel A. Teruel, Elena Navarro, Víctor López-Jaquero, Francisco Montero, and Pascual González. A comprehensive framework for modeling requirements of CSCW systems. Journal of Software: Evolution and Process, 29(5): ??, May 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tamburri:2021:ESF

[TP21] Damian A. Tamburri and Fabio Palomba. Evolving software forges: an experience report from Apache Allura. Journal of Software: Evolution and Process, 33(12):e2397:1– e2397:??, December 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tufano:2017:TBA

[TPB^{+}17] Michele Tufano, Fabio Palomba, Gabriele Bavota, Massimiliano Di Penta, Rocco Oliveto, Andrea De Lucia, and Denys Poshyvanyk. There and back again: Can you compile that snapshot? Journal of Software: Evolution and Process, 29(4):??, April 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Trektere:2017:MMA

[TRC^{+}17] Kitija Trektere, Gilbert Regan, Fergal Mc Caffery, Derek Flood, Marion Lepmets, and Grainne Barry. Mobile medical app development with a focus on traceability. Journal of Software: Evolution and Process, 29(11): ??, November 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Turetken:2017:AAL

[TST17] Oktay Turetken, Igor Stojanov, and Jos J. M. Trienekens. Assessing the adoption level of scaled agile development: a maturity model for Scaled Agile Framework. Journal of Software: Evolution and Process, 29(6): ??, June 2017. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tariq:2020:SBP

[TUF20] Sidra Tariq, Muhammad Usman, and Alvis C. M. Fong. Selecting best predictors from large software repositories for highly accurate software effort estimation. Journal of Software: Evolution and Process, 32(10):e2271:1– e2271:??, October 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tolfo:2018:APP

[TWFF18] Cristiano Tolfo, Raul Sidnei Wazlawick, Marcelo Gitirana Gomes Ferreira, and Fernando Antonio Forcellini. Agile practices and the promotion of entrepreneurial skills in software development. Journal of Software: Evolution and Process, 30(9): e1945:1–e1945:??, September 2018. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Tian:2021:ITS

[TWL^{+}21] Fangchao Tian, Tianlu Wang, Peng Liang, Chong Wang, Arif Ali Khan, and Muhammad Ali Babar. The impact of traceability on software maintenance and evolution: a mapping study. Journal of Software: Evolution and Process, 33 (10):e2374:1–e2374:??, October 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ukor:2012:SSH

[UC12] Roland Ukor and Andy Carpenter. Service selection and horizontal multi-sourcing in processoriented capability outsourcing. Journal of Software: Evolution and Process, 24(3):259–283, April 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Uzunov:2018:AIQ

[UFF18] Anton V. Uzunov, Eduardo B. Fernandez, and Katrina Falkner. Assessing and improving the quality of security methodologies for distributed systems. Journal of Software: Evolution and Process, 30(11): e1980:1–e1980:??, November 2018. CODEN ????

ISSN 2047-7473 (print), 2047-7481 (electronic).

Unterkalmsteiner:2014:CFS

[UGI^{+}14] Michael Unterkalmsteiner, Tony Gorschek, A. K. M. Moinul Islam, Chow Kian Cheng, Rahadian Bayu Permadi, and Robert Feldt. A conceptual framework for SPI evaluation. Journal of Software: Evolution and Process, 26(2): 251–279, February 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Usman:2020:AMB

[UIK20] Muhammad Usman, Muhammad Zohaib Iqbal, and Muhammad Uzair Khan. An automated modelbased approach for unitlevel performance test generation of mobile applications. Journal of Software: Evolution and Process, 32 (1):e2215:1–e2215:??, January 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Upadhyaya:2013:AMS

[UTZ13] Bipin Upadhyaya, Ran Tang, and Ying Zou. An approach for mining service composition patterns from execution logs. Journal of Software: Evolution and Process, 25(8):841– 870, August 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ulziit:2015:CFC

[UWGP15] Bayarbuyan Ulziit, Zeeshan Akhtar Warraich, Cigdem Gencel, and Kai Petersen. A conceptual framework of challenges and solutions for managing global software maintenance. Journal of Software: Evolution and Process, 27(10):763–792, October 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Unger-Windeler:2021:POC

[UWKRS21] Carolin Unger-Windeler, Jil Ann-Christin Klünder, Timothy Reuscher, and Kurt Schneider. Are Product Owners communicators? A multi-method research approach to provide a more comprehensive picture of Product Owners in practice. Journal of Software: Evolution and Process, 33(1): e2311:1–e2311:??, January 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ural:2013:RTS

[UY13] Hasan Ural and Hüsnü Yenigün. Regression test suite selection using dependence analysis. Journal of Software: Evolution and Process, 25(7):

681–709, July 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vavpotic:2022:SPE

[VBMH22] Damjan Vavpotic, Saimir Bala, Jan Mendling, and Tomaz Hovelja. Software process evaluation from user perceptions and log data. Journal of Software: Evolution and Process, 34(4):e2438:1– e2438:??, April 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

vandeGriend:2012:ITE

[vdGK12] P. A. van de Griend and R. J. Kusters. Integration test effort in SAP R/3 systems. Journal of Software: Evolution and Process, 24(4):421–435, June 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vince:2022:EHV

[VHBK22] Dániel Vince, Renáta Hodován, Daniella Bársony, and Ákos Kiss. The effect of hoisting on variants of Hierarchical Delta Debugging. Journal of Software: Evolution and Process, 34(11):e2483:1– e2483:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vogel:2021:MAS

[VKC^{+}21] Martin Vogel, Peter Knapik, Moritz Cohrs, Bernd Szyperrek, Winfried Pueschel, Haiko Etzel, Daniel Fiebig, Andreas Rausch, and Marco Kuhrmann. Metrics in automotive software development: a systematic literature review. Journal of Software: Evolution and Process, 33(2): e2296:1–e2296:??, February 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vanya:2013:SSA

[VKPvV13] Adam Vanya, Steven Klusener, Rahul Premraj, and Hans van Vliet. Supporting software architects to improve their software system’s decomposition — lessons learned. Journal of Software: Evolution and Process, 25(3): 219–232, March 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Varkoi:2020:VES

[VMCN20] Timo Varkoi, Timo Mäkinen, Frank Cameron, and Risto Nevalainen. Validating effectiveness of safety requirements’ compliance evaluation in process assessments. Journal of Software: Evolution and Process, 32(3):e2177:1– e2177:??, March 2020. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vaucher:2013:AEW

[VMSH13] Stéphane Vaucher, Antoine Moulart, Houari Sahraoui, and Naji Habra. Automated evaluation of website navigability: an empirical validation of multilevel quality models. Journal of Software: Evolution and Process, 25 (8):815–839, August 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Varkoi:2014:TNS

[VNM14] Timo Varkoi, Risto Nevalainen, and Timo Mäkinen. Toward nuclear SPICE — integrating IEC 61508, IEC 60880 and SPICE. Journal of Software: Evolution and Process, 26(3): 357–365, March 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vlietland:2014:III

[VvV14] Jan Vlietland and Hans van Vliet. Improving IT incident handling performance with information visibility. Journal of Software: Evolution and Process, 26(12):1106–1127, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Vlietland:2015:ISE

[VvV15] Jan Vlietland and Hans van Vliet. Information sharing for effective IT incident resolving in IT service provider networks: a financial service case study. Journal of Software: Evolution and Process, 27 (2):73–94, February 2015. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Walker:2018:CSC

[Wal18] Alastair Walker. Cybersecurity in safety-critical systems. Journal of Software: Evolution and Process, 30 (5):e1956:1–e1956:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2021:OSS

[Wan21]  Jinyong Wang.  Open

source software reliability model with nonlinear fault detection and fault introduction. Journal of Software: Evolution and Process, 33(12): e2385:1–e2385:??, December 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wen:2022:CSW

[WCDY22] Yingying Wen, Guanjie Cheng, Shuiguang Deng, and Jianwei Yin. Characterizing and synthesizing the workflow structure of microservices in ByteDance Cloud. Journal of Software: Evolution and Process, 34(8): e2467:1–e2467:??, August 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Walker:2014:EPC

[WCS14] Alastair Walker, Antonio Coletta, and Rama Sivaraman. An evaluation of the process capability implications of the requirements of ISO/IEC 20000-1. Journal of Software: Evolution and Process, 26(12):1316– 1326, December 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Weber:2018:ACD

[WFB^{+}18] Philip Weber, João Bosco Ferreira Filho, Behzad Bordbar, Mark Lee, Ian Litchfield, and Ruth Backman. Automated conflict detection between medical care pathways. Journal of Software: Evolution and Process, 30 (7):e1898:1–e1898:??, July 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wilde:2018:MTV

[WG18] Evan Wilde and Daniel German. Merge-Tree: Visualizing the integration of commits into Linux. Journal of Software: Evolution and Process, 30(2): e1936:1–e1936:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wnuk:2015:EFA

[WKB^{+}15] Krzysztof Wnuk, Jaap Kabbedijk, Sjaak Brinkkemper, Björn Regnell, and David Callele. Exploring factors affecting decision outcome and lead time in large-scale requirements engineering. Journal of Software: Evolution and Process, 27(9):647– 673, September 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2016:ACR

[WL16] Shaowei Wang and David Lo. AmaLgam+: Composing rich information sources for accurate bug localization. Journal of Software: Evolution and Process, 28(10):921–942, October 2016. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2023:TBC

[WLL^{+}23] Haifeng Wang, Hengyuan Liu, Zheng Li, Yong Liu, Fuxiang Sun, and Xiang Chen. A token-based compilation error categorization and its applications. Journal of Software: Evolution and Process, 35(2): e2512:1–e2512:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2014:ASM

[WPVS14] Xiaoran Wang, Lori Pollock, and K. Vijay-Shanker. Automatic segmentation of method code into meaningful blocks: Design and evaluation. Journal of Software: Evolution and Process, 26(1): 27–49, January 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2013:HDP

[WPXZ13] Jinshui Wang, Xin Peng, Zhenchang Xing, and Wenyun Zhao. How developers perform feature location tasks: a humancentric and process-oriented exploratory study. Journal of Software: Evolution and Process, 25(11):1193– 1224, November 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

White:2013:MSO

[WRC^{+}13] Laura J. White, Thomas Reichherzer, John Coffey, Norman Wilde, and Sharon Simmons. Maintenance of service oriented architecture composite applications: static and dynamic support. Journal of Software: Evolution and Process, 25(1):97– 109, January 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wilcox:2014:OBA

[WSB14] Chris Wilcox, Michelle Mills Strout, and James M. Bieman. An optimizationbased approach to lookup table program transformations. Journal of Software: Evolution and Process, 26(6):533–551, June 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wei:2021:CSS

[WSB^{+}21] Ying Wei, Xiaobing Sun, Lili Bo, Sicong Cao, Xin Xia, and Bin Li. A comprehensive study on security bug characteristics. Journal of Software: Evolution and Process, 33(10): e2376:1–e2376:??, October 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wolski:2018:SQM

[WWKC18] Marcin Wolski, Bartosz Walter, Szymon Kupiński, and Jakub Chojnacki. Software quality model for a research-driven organization — an experience report. Journal of Software: Evolution and Process, 30 (5):e1911:1–e1911:??, May 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2021:TMA

[WWQ^{+}21] Siwen Wang, Linhui Wang, Yang Qu, Rong Chen, and Shikai Guo. Toward more accurate developer recommendation via inference of development activities from interaction with bug repair process. Journal of Software: Evolution and Process, 33 (5):e2341:1–e2341:??, May 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wang:2013:ELK

[WYW^{+}13] Lei Wang, Pengzhi Yu, Zheng Wang, Chen Yang, and Qiang Ye. On the evolution of Linux kernels: a complex network perspective. Journal of Software: Evolution and Process, 25(5):439–458, May 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ward:2017:FSP

[WZ17] Martin Ward and Hussein Zedan. The formal semantics of program slicing for nonterminating computations. Journal of Software: Evolution and Process, 29 (1):??, January 2017. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Xu:2021:IAC

[XBS^{+}21] Yong Xu, Lili Bo, Xiaobing Sun, Bin Li, Jing Jiang, and Wei Zhou. image2emmet: Automatic code generation from web user interface image. Journal of Software: Evolution and Process, 33(8): e2369:1–e2369:??, August 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Xia:2015:DAR

[XLWZ15] Xin Xia, David Lo, Xinyu Wang, and Bo Zhou. Dual analysis for recommending developers to resolve bugs. Journal of Software: Evolution and Process, 27(3): 195–220, March 2015. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yaghoobi:2018:PKS

[Yag18] Tahere Yaghoobi. Prioritizing key success factors of software projects using fuzzy AHP. Journal of Software: Evolution and Process, 30(1):??, January 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yaseen:2022:SFA

[YAMM22] Muhammad Yaseen, Sikandar Ali, Aida Mustapha, and Nauman Mazhar. Success factors analysis for requirement elicitation in global software development paradigm: an empirical study. Journal of Software: Evolution and Process, 34(7):e2460:1– e2460:??, July 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chang:2012:DPR

[yCP12] Ray yaung Chang and Andy Podgurski. Discovering programming rules and violations by mining interprocedural dependences. Journal of Software: Evolution and Process, 24(1):51–66, January 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yaghoobirafi:2022:ASI

[YF22] Kamaleddin Yaghoobirafi and Ali Farahani. An approach for semantic interoperability in autonomic distributed intelligent systems. Journal of Software: Evolution and Process, 34 (10):e2436:1–e2436:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yasin:2022:UNQ

[YFL^{+}22] Affan Yasin, Rubia Fatima, Lin Liu, Javed Ali Khan, Raian Ali, and Jianmin Wang. On the utilization of non-quality assessed literature in software engineering research. Journal of Software: Evolution and Process, 34 (7):e2464:1–e2464:??, July 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yilmaz:2022:QEM

[YT22] Nebi Yilmaz and Ayça Kolukisa Tarhan. Quality evaluation models or frameworks for open source software: a systematic literature review. Journal of Software: Evolution and Process, 34 (6):e2458:1–e2458:??, June 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yang:2022:GSN

[YYF^{+}22] Kang Yang, Huiqun Yu, Guisheng Fan, Xingguang Yang, and Zijie Huang. A graph sequence neural architecture for code completion with semantic structure features. Journal of Software: Evolution and Process, 34(1): e2414:1–e2414:??, January 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Yang:2022:SAS

[YZT22] Xinyue Yang, Xiaofang Zhang, and Yao Tong. Simplified abstract syntax tree based semantic features learning for software change prediction. Journal of Software: Evolution and Process, 34(4): e2445:1–e2445:??, April 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zabardast:2022:FIS

[ZBGH22] Ehsan Zabardast, Kwabena Ebo Bennin, and Javier Gonzalez-Huerta. Further investigation of the survivability of code technical debt items. Journal of Software: Evolution and Process, 34(2): e2425:1–e2425:??, February 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zoppi:2018:LRE

[ZCL^{+}18] Tommaso Zoppi, Andrea Ceccarelli, Francesco Lo Piccolo, Paolo Lollini, Gabriele Giunta, Vito Morreale, and Andrea Bondavalli. Labelling relevant events to support the crisis management operator. Journal of Software: Evolution and Process, 30(3): e1874:1–e1874:??, March 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zarras:2022:AHS

[ZDV22] Apostolos V. Zarras, Ioannis Dinos, and Panos Vassiliadis. The athletic heart syndrome in web service evolution. Journal of Software: Evolution and Process, 34(10): e2418:1–e2418:??, October 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhu:2018:SCP

[ZHC^{+}18] Xiaoyan Zhu, Yueyang He, Long Cheng, Xiaolin Jia, and Lei Zhu. Software change-proneness prediction through combination of bagging and resampling methods. Journal of Software: Evolution and Process, 30(12):e2111:1– e2111:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zahid:2020:SRM

[ZIDM20] Maryam Zahid, Irum Inayat, Maya Daneva, and Zahid Mehmood. A security risk mitigation framework for cyber physical systems. Journal of Software: Evolution and Process, 32(2):e2219:1– e2219:??, February 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zahid:2021:SRC

[ZIDM21] Maryam Zahid, Irum Inayat, Maya Daneva, and Zahid Mehmood. Security risks in cyber physical systems — a systematic mapping study. Journal of Software: Evolution and Process, 33(9): e2346:1–e2346:??, September 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2012:TTS

[ZKJ12] He Zhang, Barbara Kitchenham, and Ross Jeffery. Toward trustworthy software process models: an exploratory study on transformable process modeling. Journal of Software: Evolution and Process, 24(7):741–763, November 2012. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2013:FSB

[ZKK13] Lingming Zhang, Miryung Kim, and Sarfraz Khurshid. FaultTracer: a spectrum-based approach to localizing failure-inducing program edits. Journal of Software: Evolution and Process, 25(12):1357–1383, December 2013. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhen:2021:CUT

[ZKN^{+}21] Ying Zhen, Abdullah Khan, Shah Nazir, Zhao Huiqi, Abdullah Alharbi, and Sulaiman Khan. Crowdsourcing usage, task assignment methods, and crowdsourcing platforms: a systematic literature review. Journal of Software: Evolution and Process, 33 (8):e2368:1–e2368:??, August 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2014:ESE

[ZKZH14] Feng Zhang, Foutse Khomh, Ying Zou, and Ahmed E. Hassan. An empirical study of the effect of file editing patterns on software quality. Journal of Software: Evolution and Process, 26(11):996–1029, November 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2021:IDL

[ZLM^{+}21] Zhuo Zhang, Yan Lei, Xiaoguang Mao, Meng Yan, Ling Xu, and Junhao Wen. Improving deep-learning-based fault localization with resampling. Journal of Software: Evolution and Process, 33(3): e2312:1–e2312:??, March 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhu:2022:SCP

[ZLW22] Xiaoyan Zhu, Nan Li, and Yong Wang. Software change-proneness prediction based on deep learning. Journal of Software: Evolution and Process, 34(4):e2434:1– e2434:??, April 2022. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zanoni:2014:PDC

[ZPFV14] Marco Zanoni, Fabrizio Perin, Francesca Arcelli Fontana, and Gianluigi Viscusi. Pattern detection for conceptual schema recovery in data-intensive systems. Journal of Software: Evolution and Process, 26(12):1172–1192, December 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zugal:2012:TEL

[ZPW12] Stefan Zugal, Jakob Pinggera, and Barbara Weber. Toward enhanced lifecycle support for declarative processes. Journal of Software: Evolution and Process, 24(3): 285–302, April 2012. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2014:SPS

[ZRB^{+}14] He Zhang, David Raffo, Thomas Birkhöltzer, Dan Houston, Raymond Madachy, Jürgen Münch, and Stanley M. Sutton, Jr. Software process simulation — at a crossroads? Journal of Software: Evolution and Process, 26(10):923– 928, October 2014. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhou:2016:CTM

[ZTGG16] Yu Zhou, Yanxiang Tong, Ruihang Gu, and Harald Gall. Combining text mining and data mining for bug report classification. Journal of Software: Evolution and Process, 28(3): 150–176, March 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2018:TRO

[ZWK18] Xuan Zhang, Xu Wang, and YanNi Kang. Trustworthiness requirementoriented software process modeling. Journal of Software: Evolution and Process, 30(12):e1991:1– e1991:??, December 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2020:GMT

[ZWW^{+}20] Yang Zhang, Huaimin Wang, Yiwen Wu, Dongyang Hu, and Tao Wang. GitHub’s milestone tool: a mixed-methods analysis on its use. Journal of Software: Evolution and Process, 32(4): e2229:1–e2229:??, April 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Zhang:2020:PBS

[ZWY^{+}20] Xuan Zhang, Xu Wang, Wei Yun, Chen Gao, Mengting Han, and Hui Liu. Pattern-based software process modeling for dependability. Journal of Software: Evolution and Process, 32(9): e2262:1–e2262:??, September 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).